### v1

# PlainQAFact: Automatic Factuality Evaluation Metric for Biomedical Plain Language Summaries Generation

Abstract
Hallucinated outputs from language models pose risks in the medical domain, especially for lay audiences making health-related decisions. Existing factuality evaluation methods, such as entailment- and question-answering-based (QA), struggle with plain language summary (PLS) generation due to elaborative explanation phenomenon, which introduces external content (e.g., definitions, background, examples) absent from the source document to enhance comprehension. To address this, we introduce PlainQAFact, a framework trained on a fine-grained, human-annotated dataset PlainFact, to evaluate the factuality of both source-simplified and elaboratively explained sentences. PlainQAFact first classifies factuality type and then assesses factuality using a retrieval-augmented QA-based scoring method. Our approach is lightweight and computationally efficient. Empirical results show that existing factuality metrics fail to effectively evaluate factuality in PLS, especially for elaborative explanations, whereas PlainQAFact achieves state-of-the-art performance. We further analyze its effectiveness across external knowledge sources, answer extraction strategies, overlap measures, and document granularity levels, refining its overall factuality assessment [^1] PlainQAFact and PlainFact are available at: https://github.com/zhiwenyou103/PlainQAFact.
PlainQAFact: Automatic Factuality Evaluation Metric for
Biomedical Plain Language Summaries Generation
Zhiwen You  Yue Guo University of Illinois Urbana-Champaign {zhiweny2, yueg}@illinois.edu
Figure 1: Illustration of our PlainQAFact framework. The inputs are PLSs and scientific abstracts. Our evaluation framework consists of a pre-trained classifier for factuality type categorization, retrieval of external domain knowledge for elaborative explanation sentences, and a local LLM-based question-answering system.
1 Introduction
Communicating biomedical scientific knowledge in plain language is essential for improving health information accessibility Stoll et al. (2022); Kuehne and Olden (2015). While large language models (LLMs) have made significant progress in generating plain language summaries (PLSs) Ondov et al. (2022); Guo et al. (2023); Goldsack et al. (2023); You et al. (2024b), the issue of hallucination remains unresolved Cao et al. (2022); Guo et al. (2024b).
Factuality evaluation for PLS is particularly challenging due to the presence of elaborative explanations, where external content such as definitions, background information, and examples is added to enhance comprehension but is absent from the original scientific abstract (source) Guo et al. (2021); Srikanth and Li (2021); Joseph et al. (2024). This creates two key issues. First, commonly used factuality metrics can effectively verify content supported by the source but fail to assess factual consistency of added information Guo et al. (2024a). Second, existing datasets for evaluating PLS factuality are limited in both data and annotation quality. FactPICO Joseph et al. (2024), for instance, focuses on LLM-generated summaries and annotates only newly added information, overlooking potential hallucinations from source simplifications. This limitation prevents a comprehensive evaluation of factual consistency across the entire PLS.
To address theses challenges, we hypothesize that improving factuality evaluation in elaborative explanations enhances overall PLS factuality assessment. To test this, we introduce the first granular factuality dataset for PLS, with expert annotations for factuality type (elaborative explanation vs. source simplification), functional role (e.g., definition, background, example), and sentence-alignment between human-authored PLS and corresponding scientific abstracts. These fine-grained annotations establish a benchmark for advancing PLS generation and factuality evaluation methods.
Building on this, we propose a dual-stage QA-based factuality evaluation framework PlainQAFact, that independently scores each PLS sentence (see Figure 1). First, a pre-trained classifier predicts whether a PLS sentence is a source simplification or an elaborative explanation. For elaborative explanations, dense retrieval via MedCPT Jin et al. (2023) retrieves relevant external information, for verification beyond the source abstract. The abstract and retrieved content are then concatenated as references for QA-based factuality evaluation. For source simplifications, the abstract alone serves as the reference for evaluation. Finally, sentence-level evaluation scores are averaged to produce a final factuality score for the PLS. Experiments on PlainFact and the CELLS dataset Guo et al. (2024b) demonstrate that PlainQAFact significantly enhances PLS factuality assessment, especially for elaborative explanations. By enabling more accurate factuality assessment in PLS, our approach enhances robustness and supports a foundation for enhancing open-ended PLS generation.
Our main contributions are as follows:
  • PlainFact: the first expert-annotated benchmark for granular factuality evaluation in PLS, covering factuality type, functional role, and sentence alignment (§3).
  • PlainQAFact: a factuality evaluation framework integrates sentence type classification with retrieval-augmented QA, leveraging expert-domain resources (§4).
  • PlainQAFact outperforms existing factuality metrics in PLS evaluation, particularly in handling elaborative explanations (§5, 6).
2 Related Work
Limitations of Existing Factuality Evaluation
The primary approach for evaluating PLS generation combines automated metrics with human evaluation (Jain et al., 2021; Ondov et al., 2022). While human assessment provides a thorough analysis Hardy et al. (2019), its high cost and time demands make it impractical for large-scale datasets. Evaluating factuality in biomedical PLS is particularly challenging, as it requires domain expertise. Entailment- Lee et al. (2022); Laban et al. (2022), similarity- Wan and Bansal (2022); Ye et al. (2024), model- Zha et al. (2023) and QA-based Deutsch et al. (2021); Fabbri et al. (2022b) metrics are commonly used for factuality assessment but rely heavily on high-quality reference summaries, which are often unavailable or difficult to obtain for PLS. Recent advancements in prompt-based evaluation show promise (Luo et al., 2023); however, their sensitivity to factuality perturbations in PLS remains limited Guo et al. (2024a). To address these limitations, we propose a reference-free solution for factuality evaluation of PLS that effectively assesses factual consistency with external information retrieval to augment the reference summary.
Retrieval-Augmented Generation
Retrieval-augmented methods enhance text generation by extracting relevant information from external sources to supplement input queries Lewis et al. (2020). These methods have been shown to be effective in open-domain QA Ren et al. (2023); Mao et al. (2020), knowledge-based QA Kim (2025), and multi-step reasoning Gao et al. (2023); Tang and Yang (2024). In the context of PLS, retrieval from structured knowledge bases (KBs) has been shown to improve factual accuracy compared to language models alone Guo et al. (2024b). However, retrieval-augmented approaches have not been extensively explored for factuality evaluation in PLS, despite their potential for addressing elaborative explanations. In this work, we investigate retrieval-augmented QA to enhance PLS factuality assessment while also examining its limitations.
3 PlainFact Benchmark
To develop a high-quality PLS factuality benchmark, we select a subset of the largest human-authored PLS dataset (§3.1) and hire domain experts to provide fine-grained annotations (§3.2).
3.1 Human-Authored PLS Dataset
Rather than relying on LLM-generated PLSs, we construct our benchmark using human-authored summaries. CELLS Guo et al. (2024b), the largest parallel corpus of scientific abstracts and their corresponding PLSs, is written by the original authors and sourced from 12 biomedical journals. We primarily select data from the Cochrane Database of Systematic Reviews (CDSR) [^2] http:///www.cochranelibrary.com within CELLS, as CDSR contains systematic reviews that support evidence-based medical decision-making across healthcare domains Murad et al. (2016). Since systematic reviews represent the highest level of scientific evidence, this selection enhances the factual rigor of our dataset. To ensure readability, we filter the 200 most readable PLSs based on average scores from three standard readability metrics: Flesch-Kincaid Grade Level (FKGL) Kincaid (1975), Dale-Chall Readability Score (DCRS) Dale and Chall (1948), and Coleman-Liau Index (CLI) Coleman (1975).
Explanation Simplification
# of Sentences 1,213 1,527
Average Length (token) 29 28
Vocabulary Size 4,230 4,046
Has Reference 417 1,527
# of Background 533 329
# of Definition 82 44
# of Method/Result 512 1,107
# of Example 10 3
# of Other 76 44
Chall $\downarrow$ 11.3 11.6
Coleman-Liau $\downarrow$ 13.5 13.9
Table 1: Overview of the PlainFact benchmark. Medical experts annotated 200 pairs of PLSs and their corresponding scientific abstracts from the CELLS dataset, categorizing each PLS sentence as either a source simplification or an elaborative explanation related to the abstract.
3.2 Expert Annotation
Since each PLS-abstract pair is authored by the same individual, we assume all information to be factual. The annotation aims to capture how PLS sentences relate to their scientific abstracts. Annotators analyzed each PLS sentence across three dimensions: (1) Factuality type: identify whether a sentence is a source simplification (derived from the abstract) or an elaborative explanation (introducing new content); (2) Functional role: categorize the sentence as background, definition, example, method/result, or other; and (3) Sentence alignment: map each PLS sentence to its corresponding sentence(s) in the scientific abstract. Details of annotation guidelines are in App. A.
Annotations were conducted by four independent annotators, each with at least a bachelor’s degree in biomedical sciences and prior fact-checking experience. Annotators were recruited via Upwork and compensated from \$15 to \$20 per hour. Each PLS-abstract pair was annotated by two independent annotators, with disagreements resolved by a third. Inter-rater agreement, measured by Cohen’s Kappa for factuality type, functional role, and sentence alignment are 0.43, 0.60, and 0.55 respectively, indicated moderate agreement for all tasks Artstein and Poesio (2008).
Table 1 summarizes the dataset characteristics. Notably, 44% of PLS sentences contain elaborative explanations, highlighting their role in enhancing PLS readability. This underscores the need for factuality evaluation methods that account for such phenomena. Moreover, elaborative explanations include more background, definitions, and examples than source simplifications. For annotation examples, see App. B.
4 PlainQAFact Framework
PlainQAFact conducts fine-grained factuality evaluation for PLS by assessing each sentence individually and retrieving tailored external knowledge within a retrieval-augmented QA framework. Figure 1 provides an overview of its three key components: sentence-level classification (§4.1), domain knowledge retrieval (§4.2), and dual-stage QA-based factuality evaluation (§4.3).
4.1 Learned Factuality Type Classifier
As elaborative explanations are prevalent in PLS generation and existing metrics struggle to capture added information Guo et al. (2024a), we fine-tune a classifier to categorize factuality types. Using factuality type annotations (source simplification vs. elaborative explanation) in PlainFact, we fine-tune a classifier with PubMedBERT-base model [^3] https://huggingface.co/microsoft/BiomedNLP-BiomedBERT-base-uncased-abstract-fulltext. The dataset is split 8:1:1 for training, validation, and testing. The classifier is fine-tuned with early stopping, stopping at the third epoch based on validation performance. Additionally, we compare the learned classifier with a zero-shot GPT-4o OpenAI (2024) classifier (details in App. C) to explore extending our factuality evaluation framework to domains and tasks lacking human-annotated data.
4.2 Domain Knowledge Retrieval
Retrieval-augmented methods have proven effective for explanation generation Guo et al. (2024b); Xu et al. (2024), making them a natural fit for evaluating elaborative explanations. Since our dataset is in the medical domain, we adopt MedCPT Jin et al. (2023), a retriever pretrained on large-scale PubMed search logs to generate biomedical text embeddings. For external resources, we incorporate StatPearls Jin et al. (2021) for clinical decision support and medical textbooks Xiong et al. (2024) for domain-specific knowledge.
4.3 QA Evaluation Components
QA-based metrics have outperformed other factuality metrics Guo et al. (2024a) in PLS and align more closely with human annotations Fabbri et al. (2022a); Deutsch et al. (2021). Therefore, we adopt a QA-based approach in our framework to verify factuality, incorporating answer extraction, question generation, question filtering, question answering, and answer overlap calculation.
Gold Answer Extraction
The first step in QA-based factuality evaluation is to extract answer entities (keyphrases) from the PLS as gold answer, then verify factuality by comparing them with answers generated by a QA model for the same questions. If the generated answer is correct or relevant, the summary is considered factual. To compare answer extraction strategies, we use noun phrase chunks (NP Chunks) as the baseline, and evaluate LLM-based extractors, including Llama 3.1 8B Instruct (Llama 3.1)Meta (2024), GPT-4o OpenAI (2024), and PromptRank Kong et al. (2023). Prompts for Llama and GPT-4o are in App. C. PromptRank is a keyphrase generation method based on the T5 model. The number of extracted entities is set to 50 and the T5-large is used for better performance.
Question Generation (QG)
Given an input document (PLS), the QG model generates questions based on the extracted answers and their context. Following previous studies Deutsch et al. (2021); Fabbri et al. (2022b) [^4] These studies use allennlp, an open-source NLP research library built on PyTorch, which is no longer actively maintained., we fine-tune BART-large model on standard QG datasets, including SQuAD Rajpurkar et al. (2016) and QA2D Demszky et al. (2018), for use as the QG model in our framework.
Question Filtering (QF)
To prevent unanswerable questions from impacting the final factuality score, we follow QAFactEval Fabbri et al. (2022a) and remove the unanswerable questions using a pre-trained Electra-large model Clark et al. . During QF, the QA model receives only the PLS and its corresponding questions. In the subsequent QA stage, answers are extracted from the source for answer overlap evaluation, while QF only determines whether the questions can be answered by the PLS.
Question Answering
The QA model extracts answers to filtered questions from the source document. To prevent hallucinated output, we use an extractive QA model, a pre-trained Electra-large, which was the best performing QA model in QAFactEval Fabbri et al. (2022a).
Answer Overlap Evaluation (AOE)
We evaluate the alignment between the gold and generated answers from the QA model using the LERC score Chen et al. (2020) and BERTScore Zhang et al. . The final step in PlainQAFact is weighting the scores for source simplification and elaborative explanation sentences. Specifically, the simplification score $s$ is computed using only abstracts as the source in QA, while the explanation score $e$ incorporates both abstracts and retrieved knowledge.
$\textsc{PlainQAFact}=\frac{s_{\text{Avg.}}\cdot n_{s}+e_{\text{Avg.}}\cdot n_{e}}{n_{s}+n_{e}},$ (1)
where $n_{s}$ is the number of simplification sentences, and $n_{e}$ is the number labeled as explanation.
5 Experiments
Since the classifier is trained on the PlainFact training set and its test set is limited to 20 scientific-PLS pairs, we conduct experiments on both PlainFact and the CELLS test set, excluding PlainFact to prevent data leakage.
5.1 Experiment Settings
Detailed model selections for each module are in §4. Prompts for GPT-4o OpenAI (2024) and Llama 3.1 8B Instruct Meta (2024) are in App. C. For the QG model, we fine-tune the BART-large with a batch size of 16, a learning rate of 3e-5, and two training epochs. In MedCPT retrieval, we set k (the number of retrieved snippets) to 3, ensuring the retrieved information remains within a short context window for the subsequent QA process (details in App. D).
5.2 Existing Factuality Metrics
To the best of our knowledge, no prior work has efficiently evaluated factuality metrics for detecting elaborative explanations in PLS. Moreover, most factuality metrics are designed for general-domain applications, largely due to limited quality and annotation of existing PLS datasets. To address this gap, we incorporate the following metrics in our experiment, informed by prior work Joseph et al. (2024): (1) Dependency-Arc Entailment (DAE) Goyal and Durrett (2021) is an entailment-based method that evaluates summary factuality by breaking it into smaller entailment tasks at the arc level. The model independently determines whether each arc’s relationship is supported by the input. (2) AlignScore Zha et al. (2023) is a model-based factuality metric using RoBERTa Liu (2019). It extracts claims from the summary and calculates the alignment scores with all the context chunks from the input document. The final score is an average of all highest alignment probabilities. (3) SummaC Laban et al. (2022) is an NLI-based inconsistency detection method designed for summarization tasks. It segments documents into sentences and aggregates scores between sentence pairs using a trained NLI model. (4) QAFactEval Fabbri et al. (2022a) is a QA-based factuality evaluation metric that assesses summary consistency by generating questions from the source document and comparing the model-generated answers with expected answers. (5) QuestEval Scialom et al. (2021a) is a reference-less evaluation metric employs a T5 model to generate questions from the source document and verifies whether the summary can correctly answer them.
6 Results and Analysis
We first analyze the impact of each component incorporated into PlainQAFact and conduct additional ablation studies to assess retrieval effectiveness on PlainFact6.1). Next, we evaluate PlainQAFact on two datasets, comparing it with five widely used factuality metrics (§6.2). Finally, we perform an error analysis to investigate the failure cases of PlainQAFact6.5).
Component Method Choice Benchmark
PlainQAFact 56.0 (0.02)
Classifier Learned classifier [^5] We acknowledge that applying the learned classifier to its training data may introduce data leakage. To further validate PlainQAFact, we evaluate its performance on the CELLS dataset, as shown in Figure 2. -
GPT-4o 56.1 (0.02)
No (retrieve for all) 58.4* (0.03)
Answer Extraction Llama3.1 8B Instruct -
GPT-4o 56.8 (0.05)
PromptRank 54.1*
NP Chunks 53.9*
Retrieval Source Abs + TB + SP -
TB + SP 48.5* (0.02)
Abs (no retrieval) 50.6* (0.03)
Answer Overlap BERTScore -
LERC 28.9* (0.03)
Granular Level Sentence -
Summary 55.8 (0.06)
Table 2: Ablation study on PlainFact, evaluating the impact of individual components in PlainQAFact. The first row represents our proposed combination. Ablations are conducted by replacing one component at a time while keeping all others fixed to match the best overall model. Abs: abstract; TB: Textbooks; SP: StatPearls. Using the first row as the baseline, * indicates a statistically significant difference based on a paired t-test ($p<0.05$). We run each setting for five times and report the standard deviation (std.). Note that only the settings using LLMs may produce fluctuated scores.
6.1 Ablation Study
Table 2 is the sentence-level ablation study of PlainQAFact where each component of our evaluation metric is modified individually. See App. E for the summary-level evaluation. Additional analysis of retrieval resources is presented in Table 3.
Learned Classifier
Evaluating all PLS sentences without classification and retrieving for all improves factuality assessment but significantly increase evaluation time. Retrieval for every sentence extends processing time from 2 hours and 40 minutes with our learned classifier to 6 hours and 30 minutes without it for 2,740 evaluation examples. This highlights the efficiency gains of incorporating a classifier in PLS factuality evaluation. Notably, the GPT-4o classifier achieves results comparable to our learned classifier trained on expert-annotated data, demonstrating the effectiveness of LLMs as alternative classifiers. This extends the applicability of PlainQAFact to factuality evaluation in other domains, as LLMs can serve as classifiers without requiring costly human annotations.
Llama for Gold Answer Extraction
The Llama-based answer extraction method outperforms NP Chunks by 2.1% by providing context-aware, flexible boundaries and robust answer selection Yu et al. (2024), making it more effective for factuality evaluation. Although GPT-4o achieves slightly better performance, the improvement is marginal, making Llama a more cost-effective alternative.
Sentence Type Source Reference Benchmark
Simplification Abs 63.3 (0.05)
Explanation Abs 30.5* (0.03)
TB 17.0* (0.02)
SP 18.4* (0.03)
TB + SP 25.2* (0.02)
Abs + TB 39.8* (0.08)
Abs + SP 40.1* (0.03)
Abs + TB + SP 44.6 (0.06)
Full PLS Abs 50.6* (0.03)
TB 45.4* (0.02)
SP 45.9* (0.03)
TB + SP 48.5* (0.02)
Abs + TB 54.1* (0.02)
Abs + SP 54.3* (0.01)
Abs + TB + SP 56.0 (0.02)
Table 3: Ablation study results on the PlainFact, evaluating how different retrieval sources affect various information types using PlainQAFact. The learned classifier categorizes sentences as source simplification (1,685) or elaborativeexplanation (1,055). Abs: abstracts; TB: Textbooks; SP: StatPearls. Using the Abs+TB+SP as the baseline, * indicates a statistically significant difference based on a paired t-test ($p<0.05$). We run each setting for five times and report the standard deviation (std.) in the brackets.
Combined Domain Resources for Retrieval
The best performance is achieved when abstracts, medical Textbooks, and StatPearls are combined as the source document. In contrast, excluding the source abstract for QA-based factuality evaluation achieves the worst performance, which is expected given that abstracts contain most of the key information. We further examine the impact of retrieval resources on different information types. As shown in Table 3, for elaborative explanation sentences, using abstracts alone leads to significantly lower scores compared to source simplification sentences, indicating that abstracts alone are insufficient for verifying external elaborative explanation content. Incorporating StatPearls alongside abstracts substantially improves performance for both elaborative explanation and full PLS, underscoring the importance of integrating high-quality domain-specific resources for factuality verification.
BERTScore for Answer Overlap
Although LERC performs well in QAFactEval, BERTScore achieves significantly better results for PLS factuality evaluation, improving performance by 27.1% on PlainFact. This disparity arises from their fundamental differences: LERC is a learned metric trained on human factual consistency scores, whereas BERTScore relies on embedding-based similarity. Given that PlainQAFact evaluates factuality by comparing generated answers with extracted gold answers from the source, BERTScore’s ability to measure fine-grained semantic similarity makes it more effective choice.
Sentence-level Evaluation
Sentence-level evaluation slightly outperforms summary-level evaluation, primarily due to differences in input context for the QG model. In sentence-level evaluation, the QG model processes one PLS sentence at a time, enabling more targeted and precise factual verification. In contrast, summary-level evaluation takes the entire PLS as input, introducing additional contextual dependencies that can lead to factual inconsistencies. This further confirms the effectiveness of finer granularity in PLS evaluation.
Figure 2: Scaled factuality accuracy on PlainFact and the CELLS test set. PlainQAFact significantly outperforms all other metrics on both datasets, except for AlignScore on PlainFact ($p<0.05$, paired t-test). We compute the std. by running the experiment five times for both our method and Llama 3.1.
6.2 PlainQAFact Achieves the Best Performance in PLS Factuality Evaluation
Figure 2 compares existing factuality metrics with PlainQAFact on the PlainFact and CELLS datasets. To ensure a fair comparison, we scale factuality accuracy to a range of 0 to 1, as QAFactEval scores range from 1 to 5. To further assess generalizability, we evaluate on 200 PLS-abstract pairs randomly sampled from the CELLS test set, sourced from journals outside PlainFact. Results indicate that AlignScore’s effectiveness depends on the proportion of elaborative versus simplification sentences. When fewer external elaborative sentences are present, its performance approaches that of PlainQAFact. However, as elaborative content increases, PlainQAFact consistently outperforms AlignScore, demonstrating its ability to better handle elaborative information absent from the source abstract.
Compared to QA-based metrics such as QAFactEval and QuestEval, PlainQAFact benefits from incorporating retrieval module, which enables a more comprehensive factuality assessment for elaborative explanations. DAE consistently performs the worst in PlainFact settings, highlighting the limitations of dependency-based factuality evaluation for PLS. Additionally, we assess the performance of Llama 3.1 as a judge for factuality evaluation. In summary-level evaluations, Llama 3.1 demonstrates superior performance across both the PlainFact and CELLS datasets, showing its robust capability for context-aware factuality assessment in plain language summarization tasks. However, its performance at the sentence level is inconsistent, while it slightly outperforms PlainQAFact on the CELLS dataset, it decreases by 16.5% on the PlainFact dataset (see App. C for the prompt used with Llama 3.1).
Figure 3: Scaled factuality accuracy on human-annotated elaborative explanation sentences from PlainFact. PlainQAFact significantly outperforms all other metrics on both evaluation levels ($p<0.05$, paired t-test). We compute the std. by running the experiment five times for both our method and Llama 3.1.
Misclassification Plain Language Sentence: This review looked at how well the methods worked to prevent pregnancy[extracted ${}_{\textit{[extracted answer]}}$, if they caused bleeding problems, if women used them as prescribed, and how safe they were. Lopez et al. (2013)
Scientific Abstract: To compare the contraceptive effectiveness, cycle control, compliance (adherence), and safety of the contraceptive patch or the vaginal ring versus combination oral contraceptives (COCs)… Lopez et al. (2013)
Generated Question: This review looked at how well the methods worked to prevent what?
Retrieved Knowledge: Appropriate treatment for the underlying etiology should start as soon as possible, and the patients and family members should receive appropriately targeted education…
QA Answer: cycle control
Retrieval Quality Plain Language Sentence: The patch is a small, thin, adhesive[extracted ${}_{\textit{[extracted answer]}}$ square that is applied to the skin. Lopez et al. (2013)
Scientific Abstract: Users of the norelgestromin-containing patch reported more breast discomfort, dysmenorrhea, nausea, and vomiting. In the levonorgestrel-containing patch trial, patch users reported less vomiting, headaches, and fatigue…Lopez et al. (2013)
Generated Question: The patch is a small, thin, what kind of square applied to the skin?
Retrieved Knowledge: Nonstick dressing Petrolatum-infused gauze strip or other material to form a bolster over the graft site. This may be sutured or taped securely in place to provide some pressure and to keep graft immobilized.
QA Answer: Nonstick dressing Petrolatum-infused gauze strip
Unanswerable Question Plain Language Sentence: Government[extracted ${}_{\textit{[extracted answer]}}$ officials, business people and health professionals implementing such measures should work together with researchers to find out more about their effects in the short and long term. von Philipsborn et al. (2019)
Scientific Abstract: To assess the effects of environmental interventions (excluding taxation) on the consumption of sugar-sweetened beverages and sugar-sweetened milk, diet-related anthropometric measures and health outcomes, and on any reported unintended consequences or adverse outcomes…von Philipsborn et al. (2019)
Generated Question: What officials, business people and health professionals implementing such measures should work together with researchers to find out more about their effects in the short and long term?
Retrieved Knowledge: Implementation should be accompanied by high-quality evaluations using appropriate study designs, with a particular focus on the long-term effects of approaches suitable for large-scale implementation.
QA Answer: long-term effects
Table 4: Error analysis of PlainQAFact with model intermediate outputs. Each plain language sentence is extracted from the PLS, the QG model generates questions based on the extracted answer and PLS sentence, and "Retrieved Knowledge" refers to the context used by the QA model for answer extraction.
6.3 Explanation-Only Evaluation
We evaluate PlainQAFact on human-annotated explanation-only sentences from PlainFact (see Figure 3), conducting both sentence- and summary-level evaluations. For summary-level setting, we aggregate all explanation sentences from each PLS to form a new summary, resulting in 196 summaries in total (4 PLS are labeled as all simplification sentences.). Results show PlainQAFact outperforms all other factuality metrics in assessing explanation information at both evaluation levels, surpassing the second-best metric by 12.2% at the sentence-level and 10.5% at the summary-level. These findings highlight PlainQAFact is effective in evaluating the factuality of elaborative explanations, particularly when external knowledge is required for verification beyond the scientific abstracts. While Llama 3.1 demonstrates superior performance at the summary-level evaluation, PlainQAFact emerges as the best performer in sentence-level setting. Compared to the results in Figure 2, AlignScore performs similarly to PlainQAFact overall on PlainFact but struggles with external sentences, where its performance drops significantly. Thus, PlainQAFact proves to be the most suitable method for the factuality evaluation of PLS.
Figure 4: Evaluation time comparison of PlainQAFact. We compare both sentence- and summary-level evaluation time over four settings: without information retrieval, PlainQAFact, and using GPT-4o as the classifier and answer extractor. We report the percentage reduction in time for summary-level evaluation given the sentence-level running time.
6.4 Evaluation Time
Figure 4 shows the evaluation times for PlainQAFact across both sentence-level and summary-level settings using PlainFact. The percentage scores indicate the reduction in evaluation time when transitioning from sentence-level to summary-level analysis. Our findings reveal that summary-level evaluation consistently reduces evaluation time across all cases. Moreover, employing PlainQAFact without MedRAG-based retrieval results in a 33.3% reduction in evaluation time for sentence-level analysis. Although integrating GPT-4o as both the classifier and answer extractor slightly improves the factuality evaluation performance (shown in Table 2), it also increases the evaluation time, particularly when GPT-4o is used as the classifier.
6.5 Error Analysis
To analyze cases where PlainQAFact fails on the PlainFact benchmark, we categorize three types of error in Table 4). For each example, we report the original PLS sentences, the corresponding scientific abstract with retrieved knowledge (from medical Textbooks and StatPearls), model-generated questions based on the extracted answers, and the QA model’s responses.
In the first case, the extracted answer is “pregnancy,” but the QA model returns “cycle control” from the abstract due to the absence of this term in the retrieved content. Since this sentence is labeled as “explanation” which requires additional knowledge for verification, the error suggests potential misclassification between simplification and explanation. The second example shows how noisy retrieved information can affect evaluation accuracy, as the QA model’s answer is irrelevant to the extracted answer. In the final case, an unanswerable question generated by the QG model (e.g., regarding “Government”) reveals limitations in QA-based evaluation, as the question remains unanswered even with retrieval. Additionally, we acknowledge that some PLS summaries or sentences may not generate any questions, resulting in empty question sets. Addressing these challenges requires further improvements in classification, retrieval, and question generation.
7 Discussion and Conclusion
Automated PLS generation has great potential for improving health information accessibility, but factuality evaluation remains challenging due to elaborative explanations. To address this, we introduce PlainQAFact, the first PLS factuality evaluation framework. PlainQAFact classifies sentence types, retrieves domain-specific knowledge, and utilizes a QA-based approach to assess the factual consistency of both source-simplified and elaborative sentences. By leveraging open-source models, PlainQAFact ensures both accessibility and computational efficiency.
To support granular factuality evaluation, we construct PlainFact, the first expert-annotated dataset for PLS, covering factuality type, functional role, and sentence alignment. Using PlainFact, our experiments demonstrate that PlainQAFact outperforms existing factuality metrics, particularly in handling external elaborative content. However, applying sentence-level evaluation to large datasets can be computationally expensive. For extensive data, we recommend using summary-level evaluation as a more efficient alternative while maintaining comparable performance.
Although our study focuses on biomedical PLS generation, the retrieval framework is adaptable to other domains by substituting domain-specific knowledge bases. Our findings highlight the effectiveness of retrieval-augmented QA for factuality evaluation, suggesting its broader applicability beyond PLS tasks. Future research should explore its integration into diverse summarization applications, particularly in LLM-generated text, where source coverage is often incomplete.
By introducing a fine-grained and explainable evaluation framework, our work establishes a foundation for advancing QA-based factuality metrics. We hope PlainQAFact will contribute to the development of more reliable and interpretable factuality assessment methods in NLP.
Limitations
We acknowledge that using an LLM as an answer extractor may yield non-deterministic outputs, even when setting the temperature to 0.01 for the Llama 3.1 8B Instruct model. For example, in our retrieval effectiveness evaluation (Table 3), we run PlainQAFact four times to examine the impact of LLM-based answer extraction on the final factuality prediction. The resulting scores were 50.60, 50.64, 50.57, and 50.63, with a standard deviation of 0.027. We consider this variability to be minimal, demonstrating that our factuality evaluation metric shows robust performance in the plain language summarization evaluation task.
Our study’s major limitations are as follows: 1) Although we provide substitute solutions for classifying input summaries or sentences, our pre-trained classifier is limited to the biomedical domain and specifically designed for plain language summarization tasks, fine-tuned through PlainFact. We acknowledge that the dataset used for model fine-tuning is insufficient, resulting in only moderate classification accuracy. Given the scarcity of high-quality human-annotated data for classifying source simplification and elaborative explanation information, future efforts may focus on developing domain- and task-specific classifiers You et al. (2024a) for factuality evaluation. 2) In the knowledge retrieval phase, we currently incorporate only Textbooks and StatPearls as external resources. This limited selection may not be sufficient for evaluating the factuality of plain language summaries. Further investigation into additional relevant knowledge bases, such as Wikipedia and PubMed, could enhance performance in explanation-needed factuality evaluation tasks. 3) For retrieval-augmented QA, we directly concatenate the scientific abstract with the retrieved information to form the reference for QA. However, we do not apply knowledge filtering to remove irrelevant information from external resources, which may introduce noise into the subsequent QA process for assessing answer-level factual correctness.
Acknowledgment
This work used Delta GPU at NCSA through allocation [CIS240504] from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S. National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
Appendix A Dataset Annotation Protocol
We developed a comprehensive annotation procedure for freelancers on Upwork to conduct fact-checking annotations.
Our annotation procedure involves two stages, starting with thorough training using detailed examples to ensure consistent understanding of the task between annotators. Annotators receive a spreadsheet where each row contains a pair of data: a sentence extracted from the PLS and its corresponding scientific abstract. For every sentence-abstract pair, the annotators are required to label three features: External, Category, and Relation, with the appropriate labels.
Step 1: Sentence Annotation: Compared to the scientific abstract, analyze each sentence of the plain language summary across three dimensions: external information, category, and relation.
1. External: Determine whether the sentence includes information does not present in the scientific abstract. 1) Yes: The sentence contains external information that is not explicitly mentioned, paraphrased, or implied in the scientific abstract. 2) No: The sentence contains information that is explicitly stated or closely paraphrased from the scientific abstract. 2. Category: Classify the sentence of the plain language summary into one of the following categories (you can only choose one category per annotation) 1) Definition: Provides a fundamental explanation of a term. 2) Background: Information that helps understand the term within the context of the abstract, such as relevance, significance, or motivation. 3) Example: Specific instances that illustrate the use of the term in the scientific abstract. 4) Method/result: Details about the methodology or results described in the scientific abstract. 5) Other: For sentences that do not fit into the categories above, please indicate the category 3. Relation: Identify the sentence(s) in the scientific abstract that the sentence of the plain language summary is related to. Use indices of sentences from the scientific abstract to link the sentence of the plain language summary. You can select one or more sentences from the scientific abstract. List the indices like s1_1,s2_3,s3_6. If no relation is found, mark it as “external.”
In Step 2, we provide explanations of the existing sentence-level indexes for plain language sentences and scientific abstracts.
Step 2: Using the Annotation Spreadsheet: You will work within a structured spreadsheet containing the segmented sentences from both the summaries and the scientific abstracts.
1. Target_Summary_ID: A unique identifier for each plain language summary. Sentences from the same plain language summary share the same ID.
2. Target_Sentence_Index: An identifier for each sentence within a summary, forming as tx_y, where ‘tx’ is the same as its ‘Target_Summary_ID’, and ‘y’ represents the index of each sentence in the plain language summary, starting from 1. e.g., t0_1 refers to the first sentence of the first summary)
3. Target_Sentence: The plain language sentence you are annotating.
4. Original_Abstract: The abstract corresponding to each summary, with each sentence indexed for easy reference.
Each annotator will annotate 40 PLS-abstract pairs to ensure each sentence of the PLS has two sets of annotations from different people. For each row of the spreadsheet, they need to annotate three columns: “External,” “Category,” and “Relation.”
$<$TO BE ANNOTATED$>$ External: Mark “Yes” or “No” to indicate if the sentence in the ‘Target_Sentence’ column contains external information.
$<$TO BE ANNOTATED$>$ Category: Choose the most fitting category of the sentence from the list (Definition, Background, Example, Method/Result, Other).
$<$TO BE ANNOTATED$>$ Relation: List the relevant sentence indices from the abstract in the ‘Original_Abstract’ column that relate to the plain language sentence (e.g., s10_1,s10_5). For example, filling in ‘s10_1,s10_5,s10_9’ if you think these three sentences from the abstract are relevant to sentence t1_3. Use commas to separate multiple indices.
To ensure annotators fully understand the context and task requirements, we provide comprehensive annotation training and a screening test prior to the annotation process. We select candidates from freelancers with a medical education background, and only those who pass the screening test are finalized as annotators.
Feature Annotation
Example 1 Plain Language Sentence Gout caused by crystal formation in the joints due to high uric acid levels in the blood.
Need External Information? yes
Category Background
Relation external
Corresponding Abstract None
Example 2 Plain Language Sentence Reducing blood pressure with drugs has been a strategy used in patients suffering from an acute event in the heart or in the brain, such as heart attack or stroke.
Need External Information? yes
Category Background
Relation s10_1,s10_2
Corresponding Abstract $<$s10_1$>$Acute cardiovascular events represent a therapeutic challenge. $<$s10_2$>$Blood pressure lowering drugs are commonly used and recommended in the early phase of these settings.
Example 3 Plain Language Sentence We looked at whether choice of antibiotic made a difference in the number of people who experienced failed treatment, and we determined the proportions who had resolution of fever at 48 hours.
Need External Information? no
Category Method/Result
Relation s15_16,s15_17,s15_20
Corresponding Abstract $<$s15_16$>$For treatment failure, the difference between doxycycline and tetracycline is uncertain (very low-certainty evidence). $<$s15_17$>$Doxycycline compared to tetracycline may make little or no difference in resolution of fever within 48 hours (risk ratio (RR) 1.14, 95% confidence interval (CI) 0.90 to 1.44, 55 participants; one trial; low-certainty evidence) and in time to defervescence (116 participants; one trial; low-certainty evidence). $<$s15_20$>$For most outcomes, including treatment failure, resolution of fever within 48 hours, time to defervescence, and serious adverse events, we are uncertain whether study results show a difference between doxycycline and macrolides (very low-certainty evidence).
Table 5: Examples of our curated dataset. Need External Information feature represents whether a plain language sentence is a simplification or an explanation. A label of “yes” indicates that the sentence is an explanation and requires additional elaborative information beyond the source abstract to verify its factuality. Conversely, a label of “no” shows that the sentence is a simplification that can be validated using only the source abstract.
Appendix B Dataset Examples
In Table 5, we presents three representative examples from PlainFact. Each example is annotated with five features: a plain language sentence, an indicator of whether the sentence is simplification or explanation, its category, its relation, and the corresponding abstract. All plain language sentences are factual, as they were written by the authors from the Cochrane database. “Need External Information?” feature specifies whether a sentence can be validated solely by the abstract. A “Yes” label indicates that the sentence includes information not explicitly mentioned in the abstract, and vice versa. The “Relation” feature identifies the sentence(s) in the scientific abstract most relevant to the PLS; if no corresponding content exists in the abstract, the relation is marked as “external.” We randomly sample three examples from the dataset to to illustrate the dataset’s structure. Additionally, indexes have been created for both the plain language sentences and the abstract sentences to facilitate annotation.
Appendix C LLM Prompts
We utilize two types of prompts to guide LLMs through two stages of factuality evaluation: classification and answer extraction. In accordance with the benchmark annotation protocol (Appendix A), we employ GPT-4o as a classifier to determine whether a given sentence or summary requires elaborative explanations for factual verification. For both stages, we set the max_tokens parameter to 512 and configure the temperature to 0 for GPT-4o and 0.01 for the Llama 3.1 8B Instruc model.
Developer
Annotate whether a sentence or summary includes information not present in the original abstract.
The sentence or summary contains external information that is not explicitly mentioned, paraphrased, or implied in the original abstract will be labeled as Yes’.
The sentence or summary contains information that is explicitly stated or closely paraphrased from the original abstract will be labeled as No’.
User
Sentence or summary: <input>
Original abstract: <abstract>
Example 1: Prompt of GPT-4o as the Classifier.
For the AE stage, we explore both GPT-4o and Llama 3.1 8B Instruct as backbone models to assess the factuality evaluation performance of PlainQAFact. Following the task description outlined in QAFactEval Fabbri et al. (2022a), we instruct both LLMs to extract potential answer entities from the input PLS.
Developer
QA-based metrics compare information units between the summary and source, so it is thus necessary to first extract such units, or answers, from the given summary. Please extract answers or information units from a plain language summary.
User
Extract a comma-separated list of the most important keywords from the following text: <input>
Example 2: Prompt of GPT-4o as the Answer Extractor.
System
QA-based metrics compare information units between the summary and source, so it is thus necessary to first extract such units, or answers, from the given summary. Please extract answers or information units from a plain language summary.
User
Extract a comma-separated list of the most important keywords from the following text: <input>
Example 3: Prompt of Llama 3.1 8B Instruct as the Answer Extractor.
Additionally, we report the performance of using Llama 3.1 8B Instruct as a judge to evaluate the factuality of a given PLS based on its source scientific abstract.
System
Rate the factuality of the given plain language sentence or summary compared with the scientific abstract. Output a numeric score from 0 to 100, with 100 meaning the sentence is completely factually consistent with the abstract and 0 meaning the sentence is completely non-factual with the abstract.
User
Sentence or summary: <input>
Original abstract: <abstract>
Factuality score (only output a numeric score): <score>
Example 4: Prompt of Llama 3.1 8B Instruct as a factuality judge.
Appendix D Detailed Experiment Settings
All experiments we conduct are under one NVIDIA A100 GPU with 40 GB GPU memory. We employ the Natural Language Toolkit (NLTK) [^6] https://www.nltk.org/ for PLS tokenization and sentence segmentation, utilizing the punkt package. For classifier fine-tuning, we set the random seed to 42 to split the dataset into training, validation, and test sets. We tune the PubMedBERT-base model for three epochs with early stopping under the validation set. The same seed (42) is also used to sample 200 PLS-abstract pairs from the CELLS test set for the comparison in Figure 2. The QA model used in our PlainQAFact is downloaded from QAFactEval Fabbri et al. (2022a); however, we set its maximum input length to 512 (from 364) tokens to incorporate more source context. Similarly, to ensure that the answers extracted during the AE stage are valid (i.e., within the 512-token limit to maintain consistency with the subsequent QA model), we configure the LLMs’ input lengths to 512 tokens.
Component Method Choice Benchmark
PlainQAFact 55.8 (0.12)
Classifier Learned classifier -
GPT-4o 59.1 (0.05)
No (retrieve for all) 60.1* (0.18)
Answer Extraction Llama3.1 8B Instruct -
GPT-4o 54.8 (0.15)
PromptRank 48.4*
NP Chunks 50.5*
Retrieval Source Abs + TB + SP -
TB + SP 47.9* (0.11)
Abs (no retrieval) 55.2 (0.07)
Answer Overlap BERTScore -
LERC 32.9* (0.18)
Table 6: Ablation study of PlainQAFact on PlainFact in the summary level, analyzing the impact of individual components in PlainQAFact. Abs: abstract; TB: Textbooks; SP: StatPearls. Using the first row as the baseline, * indicates a statistically significant difference based on a paired t-test ($p<0.05$). We run each setting for five times and report the standard deviation (std.). Note that only the settings using LLMs may produce fluctuated scores.
Appendix E Summary-Level Ablation Study of PlainQAFact
Similar to the sentence-level ablation studies in Table 2, we conduct experiments to evaluate the impact of various components in PlainQAFact at the summary level using PLSs as inputs. Although the overall evaluation score using GPT-4o as the classifier is 3.3% higher than that of PlainQAFact, the paired t-test (p<0.05) does not provide sufficient statistical evidence to conclude that GPT-4o significantly outperforms our learned classifier. This finding indicates that GPT-4o’s performance is comparable with our learned classifier, consistent with the sentence-level results.
Notably, in the summary-level evaluation, using only scientific abstracts as references for QA-based evaluation yields performance that is competitive with using retrieved external knowledge. We assume that summary-level retrieval may not consistently produce highly relevant knowledge snippets and might even introduce noise that undermines factuality evaluation accuracy.
Additionally, omitting the classification stage and retrieving information for all summaries produces the best performance among all settings. Although this approach increases the runtime to approximately 30 minutes compared to using the learned classifier for 200 summaries, the performance improvement is significant. Therefore, we recommend either employing more targeted knowledge bases for summary-level retrieval, compressing abstracts to include more relevant information, or retrieving for all input summaries when conducting summary-level factuality evaluation with PlainQAFact.
Figure 5: Evaluation performance across metrics for explanation sentences under PlainFact. We select only those summaries containing at least five explanation sentences, resulting in a total of 116 summaries. Note that this experiment is conducted at the summary level. A higher change percentage indicates a metric is more sensitive to the explanation information, implying that such metrics may not effectively verify the factuality of PLS. Ideally, we expect the score change percentage to increase slowly. Among the metrics evaluated, only AlignScore, SummaC, and our method show this expected pattern.
Appendix F Explanation Sentence Removal
To further assess the performance of existing factuality metrics on elaborative explanation sentences in PLSs, we conduct experiments in which these explanation sentences were iteratively removed from the summaries. Figure 5 illustrates the percentage change in scores relative to the baselines as more explanation sentences are removed. Notably, PlainQAFact can even predict 100% factual ratings for the given PLS when more explanation information is removed, whereas other metrics struggle to do so. Although our framework shows moderate performance in Figure 5, with a gradual increase in score as explanation sentences are removed—we believe that our retrieval-augmented, QA-based factuality evaluation framework more effectively assesses the factuality of plain language sentences that require elaborative explanations for verification.
Appendix G Pilot Study on FactPICO
In PLS generation tasks, we employ the recently introduced FactPICO dataset Joseph et al. (2024). This dataset comprises human-labeled plain language summaries of Randomized Controlled Trials (RCTs) that address several critical elements: Populations, Interventions, Comparators, Outcomes (PICO), and additional information. All summaries are generated by various LLMs based on scientific literature and include extensive explanations not present in the original abstracts. We hypothesize that existing factuality evaluation metrics for text summarization may struggle to accurately assess the factuality of this added information. To validate this assumption, we conduct pilot studies using four factuality evaluation metrics from FactPICO alongside one NLI-based metric: DAE Goyal and Durrett (2021), AlignScore Zha et al. (2023), SummaC Laban et al. (2022), QAFactEval Fabbri et al. (2022a), and QuestEval Scialom et al. (2021b).
G.1 Dataset Pre-processing
FactPICO provides span-level annotations for LLM-generated summaries, assessing whether the additional information is present and determining its factuality, labeled as either “yes” or “no.” We first remove all special identification tags from the abstracts, such as “ABSTRACT.BACKGROUND,” “ABSTRACT.RESULTS,” and “ABSTRACT.CONCLUSIONS.” We then deduplicate the generated summaries, while retaining duplicates within abstracts, as each abstract has three summaries generated by different LLMs, and each abstract may have varying numbers of generated summaries, from none to multiple. This pilot study focuses on two key research questions: evaluating the effectiveness of existing factuality metrics in assessing added information and non-factual added information.
G.2 Experiments and Results
We conduct experiments using the processed FactPICO dataset to address both research questions. It is important to note that there are some outliers in the PLSs which contain extraneous content and exceed the input length limitations of the DAE and QAFactEval metrics. Therefore, we are unable to include these two metrics in this pilot study.
Figure 6: Score change percentage from baselines over three metrics on the FactPICO dataset in removing factual added information (88 pairs). We expect each model get higher performance when more external information is removed. The evaluation dataset contains 88 valid summary-abstract pairs.
RQ1: Do existing metrics perform well when external information is added to PLS compared with no added information?
In this study, we evaluate the ability of existing metrics to detect added information in PLSs. We focus on PLSs where all added information is annotated as factual, resulting in a deduplicated dataset of 88 PLS-abstract pairs. To assess metric sensitivity to the added information, we iteratively remove sentences from each PLS that contain added spans. For example, if a span such as “of a medicine called haloperidol” is labeled as factual (“yes”), we remove the entire sentence in the original PLS containing that span through exact matching, continuing this process until no added information remains. This procedure allows us to determine how effectively existing metrics handle external information that is not present in the original abstracts. As shown in Figure 6, we report performance changes relative to baseline scores. For example, AlignScore increases by 3.33 (on a scale of 100) when six units of added information are removed, leading to a change of approximately 4.9%. These results indicate that added information has impacts on model-based factuality evaluation metrics such as AlignScore and SummaC, with both scores improving as more added information is removed. In contrast, QuestEval’s score decreases with the removal of added information.
Figure 7: Score change percentage from baselines over three metrics on the FactPICO dataset in removing non-factual added information (60 pairs). We expect each metric’s score increases when more non-factual information is removed.
RQ2: Can existing metrics distinguish between non-factual and factual added information?
In this research question, our goal is to evaluate the sensitivity of factuality evaluation metrics in detecting non-factual information within PLSs. The FactPICO dataset labels added information as either “yes” (factual) or “no” (non-factual). For this scenario, we sample only those PLSs that contain both “yes” and “no” labels for added spans. As with RQ1, we iteratively remove sentences containing non-factual added spans from each PLS until no non-factual sentences remain. Overall, as illustrated in Figure 7, most evaluation metrics show improved performance as more non-factual information is removed. Nevertheless, our findings suggest that existing factuality evaluation metrics have limited capacity to accurately distinguish between factual and non-factual added information in text summarization tasks.

### v2

@@
-PlainQAFact: Automatic Factuality Evaluation Metric for Biomedical Plain Language Summaries Generation
+PlainQAFact: Retrieval-augmented Factual Consistency Evaluation Metric for Biomedical Plain Language Summarization
 Abstract
-Hallucinated outputs from language models pose risks in the medical domain, especially for lay audiences making health-related decisions.
-Existing factuality evaluation methods, such as entailment- and question-answering-based (QA), struggle with plain language summary (PLS) generation due to elaborative explanation phenomenon, which introduces external content (e.g., definitions, background, examples) absent from the source document to enhance comprehension.
-To address this, we introduce PlainQAFact, a framework trained on a fine-grained, human-annotated dataset PlainFact, to evaluate the factuality of both source-simplified and elaboratively explained sentences.
-PlainQAFact first classifies factuality type and then assesses factuality using a retrieval-augmented QA-based scoring method.
-Our approach is lightweight and computationally efficient.
-Empirical results show that existing factuality metrics fail to effectively evaluate factuality in PLS, especially for elaborative explanations, whereas PlainQAFact achieves state-of-the-art performance.
-We further analyze its effectiveness across external knowledge sources, answer extraction strategies, overlap measures, and document granularity levels, refining its overall factuality assessment [^1] PlainQAFact and PlainFact are available at: https://github.com/zhiwenyou103/PlainQAFact.
-PlainQAFact: Automatic Factuality Evaluation Metric for
-Biomedical Plain Language Summaries Generation
-Zhiwen You  Yue Guo University of Illinois Urbana-Champaign {zhiweny2, yueg}@illinois.edu
-Figure 1: Illustration of our PlainQAFact framework.
-The inputs are PLSs and scientific abstracts.
-Our evaluation framework consists of a pre-trained classifier for factuality type categorization, retrieval of external domain knowledge for elaborative explanation sentences, and a local LLM-based question-answering system.
+Hallucinated outputs from large language models (LLMs) pose risks in the medical domain, especially for lay audiences making health-related decisions.
+Existing automatic factual consistency evaluation methods, such as entailment- and question-answering (QA) -based, struggle with plain language summarization (PLS) due to elaborative explanation phenomenon, which introduces external content (e.g., definitions, background, examples) absent from the scientific abstract to enhance comprehension.
+To address this, we introduce PlainQAFact, an automatic factual consistency evaluation metric trained on a fine-grained, human-annotated dataset PlainFact, for evaluating factual consistency of both source-simplified and elaborately explained sentences.
+PlainQAFact first classifies sentence type, then applies a retrieval-augmented QA scoring method.
+Empirical results show that existing evaluation metrics fail to evaluate the factual consistency in PLS, especially for elaborative explanations, whereas PlainQAFact consistently outperforms them across all evaluation settings.
+We further analyze PlainQAFact’s effectiveness across external knowledge sources, answer extraction strategies, answer overlap measures, and document granularity levels, refining its overall factual consistency assessment.
+Taken together, our work presents the first evaluation metric designed for PLS factual consistency evaluation, providing the community with both a robust benchmark and a practical tool to advance reliable and safe plain language communication in the medical domain.
+PlainQAFact and PlainFact are available at: https://github.com/zhiwenyou103/PlainQAFact.
+keywords:
+Plain language summarization , Factual consistency evaluation , Retrieval-augmented generation , Hallucination , Large language models
+[…]
+[label1]organization=School of Information Sciences,
+University of Illinois Urbana-Champaign,city=
+Champaign, postcode=61820, state=IL, country=United States
 1 Introduction
-Communicating biomedical scientific knowledge in plain language is essential for improving health information accessibility Stoll et al.
-(2022); Kuehne and Olden (2015).
-While large language models (LLMs) have made significant progress in generating plain language summaries (PLSs) Ondov et al.
-(2022); Guo et al.
-(2023); Goldsack et al.
-(2023); You et al.
-(2024b), the issue of hallucination remains unresolved Cao et al.
-(2022); Guo et al.
-(2024b).
-Factuality evaluation for PLS is particularly challenging due to the presence of elaborative explanations, where external content such as definitions, background information, and examples is added to enhance comprehension but is absent from the original scientific abstract (source) Guo et al.
-(2021); Srikanth and Li (2021); Joseph et al.
-(2024).
-This creates two key issues.
-First, commonly used factuality metrics can effectively verify content supported by the source but fail to assess factual consistency of added information Guo et al.
-(2024a).
-Second, existing datasets for evaluating PLS factuality are limited in both data and annotation quality.
-FactPICO Joseph et al.
-(2024), for instance, focuses on LLM-generated summaries and annotates only newly added information, overlooking potential hallucinations from source simplifications.
-This limitation prevents a comprehensive evaluation of factual consistency across the entire PLS.
-To address theses challenges, we hypothesize that improving factuality evaluation in elaborative explanations enhances overall PLS factuality assessment.
-To test this, we introduce the first granular factuality dataset for PLS, with expert annotations for factuality type (elaborative explanation vs.
-source simplification), functional role (e.g., definition, background, example), and sentence-alignment between human-authored PLS and corresponding scientific abstracts.
-These fine-grained annotations establish a benchmark for advancing PLS generation and factuality evaluation methods.
-Building on this, we propose a dual-stage QA-based factuality evaluation framework PlainQAFact, that independently scores each PLS sentence (see Figure 1).
-First, a pre-trained classifier predicts whether a PLS sentence is a source simplification or an elaborative explanation.
-For elaborative explanations, dense retrieval via MedCPT Jin et al.
-(2023) retrieves relevant external information, for verification beyond the source abstract.
-The abstract and retrieved content are then concatenated as references for QA-based factuality evaluation.
-For source simplifications, the abstract alone serves as the reference for evaluation.
-Finally, sentence-level evaluation scores are averaged to produce a final factuality score for the PLS.
-Experiments on PlainFact and the CELLS dataset Guo et al.
-(2024b) demonstrate that PlainQAFact significantly enhances PLS factuality assessment, especially for elaborative explanations.
-By enabling more accurate factuality assessment in PLS, our approach enhances robustness and supports a foundation for enhancing open-ended PLS generation.
-Our main contributions are as follows:
-  • PlainFact: the first expert-annotated benchmark for granular factuality evaluation in PLS, covering factuality type, functional role, and sentence alignment (§3).
-  • PlainQAFact: a factuality evaluation framework integrates sentence type classification with retrieval-augmented QA, leveraging expert-domain resources (§4).
-  • PlainQAFact outperforms existing factuality metrics in PLS evaluation, particularly in handling elaborative explanations (§5, 6).
+Communicating biomedical scientific knowledge in plain language is essential for improving health information accessibility and health literacy [1, 2].
+Recent advances in large language models (LLMs) have made significant progress in plain language summarization (PLS) of biomedical texts [3, 4, 5, 6].
+However, ensuring the factual consistency of these summaries remains a major challenge.
+A key source of inconsistency stems from elaborative explanations: content such as definitions, background information, and examples that enhance comprehension but are not explicitly present in the original scientific abstracts (i.e., source) [7, 8, 9].
+While such elaborations are critical for effective communication, they introduce external content that cannot be directly verified against the source, complicating automatic factual consistency evaluation.
+Factual consistency in PLS is typically assessed through a combination of human evaluation and automated metrics [3, 10].
+While human evaluation is reliable [11], it is costly and difficult to scale, particularly in biomedical contexts where domain expertise is required.
+Commonly used factuality metrics can effectively verify content supported by the source but fail to assess factual consistency of added information [12].
+However, these metrics depend heavily on high-quality reference summaries, which are often unavailable in plain language summaries.
+Recent prompt-based evaluation techniques show promise [13, 14], but their sensitivity to factual perturbations in elaborative content remains limited [12].
+The lack of suitable benchmark datasets further hinders progress.
+Many existing datasets are constructed from LLM-generated summaries or apply rule-based perturbations to simulate non-factual content.
+For example, FactPICO provides expert annotations for plain language summaries of randomized controlled trial abstracts, focusing on PICO elements and evidence inference [9].
+However, it includes factuality labels only for added content, leaving simplified sentences unannotated, which are generated by LLMs and potentially inaccurate.
+In contrast, APPLS perturbs human-written summaries using rule-based transformations [12], but cannot ensure that the resulting outputs remain coherent or factually plausible.
+These limitations underscore the need for high-quality, sentence-level annotations grounded in human-authored summaries.
+Figure 1: Overview of PlainQAFact.
+A fine-tuned classifier first identifies the sentence type, involving either source simplification or elaborative explanation.
+Then, a QA-based evaluation pipeline performs answer extraction, question generation, question answering, and answer overlap evaluation.
+For elaborative content not present in the scientific abstract, PlainQAFact retrieves external knowledge to verify factual consistency.
+The illustrated example shows an elaborative explanation involving a “bronchodilator” not mentioned in the source abstract but verifiable through external evidence.
+PlainQAFact assigns a high score, reflecting strong alignment between the extracted and gold answers.
+To address these challenges, we propose a targeted retrieval-based metric for factual consistency evaluation in PLS (Figure 1).
+We introduce a new expert-annotated dataset, PlainFact, consisting of human-written plain language summaries aligned with scientific abstracts.
+Each sentence is labeled with its type (elaborative explanation vs.
+source simplification), functional role, and alignment to the source (§3.1).
+Building on this, we present PlainQAFact, a dual-stage QA-based evaluation metric that selectively applies retrieval only to elaborative explanations (§3.2).
+This design ensures both efficiency and fidelity in evaluating factual consistency.
+Experiment results on several PLS datasets demonstrate the effectiveness of our approach, particularly in assessing complex, elaborative content (4, §5).
+Taken together, our work presents the first evaluation metric designed for PLS factual consistency evaluation, providing the community with both a robust benchmark and a practical tool to advance reliable and safe plain language communication in the medical domain.
 2 Related Work
 Limitations of Existing Factuality Evaluation
-The primary approach for evaluating PLS generation combines automated metrics with human evaluation (Jain et al., 2021; Ondov et al., 2022).
-While human assessment provides a thorough analysis Hardy et al.
-(2019), its high cost and time demands make it impractical for large-scale datasets.
-Evaluating factuality in biomedical PLS is particularly challenging, as it requires domain expertise.
-Entailment- Lee et al.
-(2022); Laban et al.
-(2022), similarity- Wan and Bansal (2022); Ye et al.
-(2024), model- Zha et al.
-(2023) and QA-based Deutsch et al.
-(2021); Fabbri et al.
-(2022b) metrics are commonly used for factuality assessment but rely heavily on high-quality reference summaries, which are often unavailable or difficult to obtain for PLS.
-Recent advancements in prompt-based evaluation show promise (Luo et al., 2023); however, their sensitivity to factuality perturbations in PLS remains limited Guo et al.
-(2024a).
-To address these limitations, we propose a reference-free solution for factuality evaluation of PLS that effectively assesses factual consistency with external information retrieval to augment the reference summary.
+The primary approach for evaluating plain language generation combines automated metrics with human evaluation [3, 10].
+While human assessment provides a thorough analysis [11], its high cost and time demands make it impractical for large-scale datasets.
+Evaluating factual consistency in biomedical PLS is particularly challenging, as it requires domain expertise.
+Entailment- [15, 16], similarity- [17, 18], model- [19] and QA-based [20, 21] metrics are commonly used for factual consistency assessment but rely heavily on high-quality reference summaries, which are often unavailable or difficult to obtain for PLS.
+Recent advancements in prompt-based evaluation show promise [13]; however, their sensitivity to factuality perturbations in PLS remains limited [12].
+To address these limitations, we propose a reference-free solution for factual consistency evaluation of PLS that effectively assesses factual consistency with external information retrieval to augment the reference summary.
 Retrieval-Augmented Generation
-Retrieval-augmented methods enhance text generation by extracting relevant information from external sources to supplement input queries Lewis et al.
-(2020).
-These methods have been shown to be effective in open-domain QA Ren et al.
-(2023); Mao et al.
-(2020), knowledge-based QA Kim (2025), and multi-step reasoning Gao et al.
-(2023); Tang and Yang (2024).
-In the context of PLS, retrieval from structured knowledge bases (KBs) has been shown to improve factual accuracy compared to language models alone Guo et al.
-(2024b).
-However, retrieval-augmented approaches have not been extensively explored for factuality evaluation in PLS, despite their potential for addressing elaborative explanations.
-In this work, we investigate retrieval-augmented QA to enhance PLS factuality assessment while also examining its limitations.
-3 PlainFact Benchmark
-To develop a high-quality PLS factuality benchmark, we select a subset of the largest human-authored PLS dataset (§3.1) and hire domain experts to provide fine-grained annotations (§3.2).
-3.1 Human-Authored PLS Dataset
-Rather than relying on LLM-generated PLSs, we construct our benchmark using human-authored summaries.
-CELLS Guo et al.
-(2024b), the largest parallel corpus of scientific abstracts and their corresponding PLSs, is written by the original authors and sourced from 12 biomedical journals.
-We primarily select data from the Cochrane Database of Systematic Reviews (CDSR) [^2] http:///www.cochranelibrary.com within CELLS, as CDSR contains systematic reviews that support evidence-based medical decision-making across healthcare domains Murad et al.
-(2016).
+Retrieval-augmented methods enhance text generation by extracting relevant information from external sources to supplement input queries [22].
+These methods have been shown to be effective in open-domain QA [23, 24], knowledge-based QA [25], and multi-step reasoning [26, 27].
+In the context of PLS, retrieval from structured knowledge bases (KBs) has been shown to improve factual accuracy compared to language models alone [28].
+However, retrieval-augmented approaches have not been extensively explored for factual consistency evaluation in PLS, despite their potential for addressing elaborative explanations.
+In this work, we investigate retrieval-augmented QA to enhance PLS factual consistency assessment while also examining its limitations.
+3 Methods
+We first introduce the protocol of PlainFact dataset curation (§3.1).
+Based on this dataset, we propose PlainQAFact, a two-stage retrieval-augmented QA framework for factual consistency evaluation in PLS tasks (§3.2).
+3.1 PlainFact Benchmark
+To develop a high-quality factual consistency evaluation benchmark in PLS tasks, we collect a subset from the largest human-authored CELLS [28] dataset (§3.1.1) and hire domain experts to provide fine-grained sentence-level annotations (§3.1.2).
+3.1.1 Human-Authored PLS Dataset
+Rather than relying on LLM-generated plain language summaries, we construct our benchmark using human-authored summaries.
+CELLS [28], the largest parallel corpus of scientific abstracts and their corresponding plain language summaries, is written by the original authors and sourced from 12 biomedical journals.
+We primarily select data from the Cochrane Database of Systematic Reviews (CDSR) [^1] http:///www.cochranelibrary.com within CELLS, as CDSR contains systematic reviews that support evidence-based medical decision-making across healthcare domains [29].
 Since systematic reviews represent the highest level of scientific evidence, this selection enhances the factual rigor of our dataset.
-To ensure readability, we filter the 200 most readable PLSs based on average scores from three standard readability metrics: Flesch-Kincaid Grade Level (FKGL) Kincaid (1975), Dale-Chall Readability Score (DCRS) Dale and Chall (1948), and Coleman-Liau Index (CLI) Coleman (1975).
-Explanation Simplification
+To ensure readability, we filter the 200 most readable plain language summaries based on average scores from three standard readability metrics: Flesch-Kincaid Grade Level (FKGL) [30], Dale-Chall Readability Score (DCRS) [31], and Coleman-Liau Index (CLI) [32].
+Additionally, given the collected plain language summaries from CELLS are all factual, we further conduct sentence-level perturbation to transform each plain language sentence into incorrect ones.
+Specifically, we use GPT-4o [^2] We use gpt-4o-2024-11-20 version for all experiments of GPT-4o in this paper.
+to perturb each plain language sentence based on prompts introduced in Guo et al.
+[12].
+Therefore, for PlainFact, we have 400 summary-abstract pairs in total, where 200 factual pairs and 200 non-factual pairs.
+Elaborative Explanation Source Simplification
 # of Sentences 1,213 1,527
@@
 # of Other 76 44
-Chall $\downarrow$ 11.3 11.6
-Coleman-Liau $\downarrow$ 13.5 13.9
+FKGL $\downarrow$ 12.5 12.4
+DCRS $\downarrow$ 11.3 11.6
+CLI $\downarrow$ 13.5 13.9
 Table 1: Overview of the PlainFact benchmark.
-Medical experts annotated 200 pairs of PLSs and their corresponding scientific abstracts from the CELLS dataset, categorizing each PLS sentence as either a source simplification or an elaborative explanation related to the abstract.
-3.2 Expert Annotation
-Since each PLS-abstract pair is authored by the same individual, we assume all information to be factual.
-The annotation aims to capture how PLS sentences relate to their scientific abstracts.
-Annotators analyzed each PLS sentence across three dimensions: (1) Factuality type: identify whether a sentence is a source simplification (derived from the abstract) or an elaborative explanation (introducing new content); (2) Functional role: categorize the sentence as background, definition, example, method/result, or other; and (3) Sentence alignment: map each PLS sentence to its corresponding sentence(s) in the scientific abstract.
-Details of annotation guidelines are in App.
-A.
-Annotations were conducted by four independent annotators, each with at least a bachelor’s degree in biomedical sciences and prior fact-checking experience.
-Annotators were recruited via Upwork and compensated from \$15 to \$20 per hour.
-Each PLS-abstract pair was annotated by two independent annotators, with disagreements resolved by a third.
-Inter-rater agreement, measured by Cohen’s Kappa for factuality type, functional role, and sentence alignment are 0.43, 0.60, and 0.55 respectively, indicated moderate agreement for all tasks Artstein and Poesio (2008).
+Medical experts annotated 200 pairs of plain language summaries and their corresponding scientific abstracts from the CELLS dataset [28], categorizing each plain language sentence as either a source simplification or an elaborative explanation related to the abstract.
+Lower scores of FKGL, DCRS, and CLI indicate better readability.
+3.1.2 Expert Annotation
+Since each summary-abstract pair is authored by the same individual, we assume all information to be factual.
+The annotation aims to capture how plain language sentences relate to their scientific abstracts.
+Annotators analyze each plain language sentence across three dimensions: (1) Factuality type: identify whether a sentence is a source simplification (derived from the abstract) or an elaborative explanation (introducing new content); (2) Functional role: categorize the sentence as background, definition, example, method/result, or other; and (3) Sentence alignment: map each plain language sentence to its corresponding sentence(s) in the scientific abstract.
+Details of annotation guidelines are provided in A.
+Annotations are conducted by four independent annotators, each with at least a bachelor’s degree in biomedical sciences and prior fact-checking experience.
+Annotators are recruited via Upwork and compensated from \$15 to \$20 per hour.
+Each summary-abstract pair is annotated by two independent annotators, with disagreements resolved by a third.
+Inter-rater agreement, measured by Cohen’s Kappa for factuality type, functional role, and sentence alignment are 0.43, 0.60, and 0.55 respectively, indicated moderate agreement for all tasks [33].
 Table 1 summarizes the dataset characteristics.
-Notably, 44% of PLS sentences contain elaborative explanations, highlighting their role in enhancing PLS readability.
-This underscores the need for factuality evaluation methods that account for such phenomena.
+Notably, 44% of plain language sentences are elaborative explanations, highlighting their role in enhancing the readability of plain language summaries.
+66% of these cannot be directly verified against the source abstract, which underscores the need for factuality evaluation methods that account for such phenomena.
 Moreover, elaborative explanations include more background, definitions, and examples than source simplifications.
-For annotation examples, see App.
-B.
-4 PlainQAFact Framework
-PlainQAFact conducts fine-grained factuality evaluation for PLS by assessing each sentence individually and retrieving tailored external knowledge within a retrieval-augmented QA framework.
-Figure 1 provides an overview of its three key components: sentence-level classification (§4.1), domain knowledge retrieval (§4.2), and dual-stage QA-based factuality evaluation (§4.3).
-4.1 Learned Factuality Type Classifier
-As elaborative explanations are prevalent in PLS generation and existing metrics struggle to capture added information Guo et al.
-(2024a), we fine-tune a classifier to categorize factuality types.
-Using factuality type annotations (source simplification vs.
-elaborative explanation) in PlainFact, we fine-tune a classifier with PubMedBERT-base model [^3] https://huggingface.co/microsoft/BiomedNLP-BiomedBERT-base-uncased-abstract-fulltext.
-The dataset is split 8:1:1 for training, validation, and testing.
-The classifier is fine-tuned with early stopping, stopping at the third epoch based on validation performance.
-Additionally, we compare the learned classifier with a zero-shot GPT-4o OpenAI (2024) classifier (details in App.
-C) to explore extending our factuality evaluation framework to domains and tasks lacking human-annotated data.
-4.2 Domain Knowledge Retrieval
-Retrieval-augmented methods have proven effective for explanation generation Guo et al.
-(2024b); Xu et al.
-(2024), making them a natural fit for evaluating elaborative explanations.
-Since our dataset is in the medical domain, we adopt MedCPT Jin et al.
-(2023), a retriever pretrained on large-scale PubMed search logs to generate biomedical text embeddings.
-For external resources, we incorporate StatPearls Jin et al.
-(2021) for clinical decision support and medical textbooks Xiong et al.
-(2024) for domain-specific knowledge.
-4.3 QA Evaluation Components
-QA-based metrics have outperformed other factuality metrics Guo et al.
-(2024a) in PLS and align more closely with human annotations Fabbri et al.
-(2022a); Deutsch et al.
-(2021).
-Therefore, we adopt a QA-based approach in our framework to verify factuality, incorporating answer extraction, question generation, question filtering, question answering, and answer overlap calculation.
+For annotation examples, see B.
+3.2 PlainQAFact Framework
+PlainQAFact conducts fine-grained factual consistency evaluation for plain language summaries by first segmenting each summary into sentences, classifying as either simplification or explanation, and retrieving tailored external knowledge within a retrieval-augmented QA framework.
+Figure 1 provides an overview of its three key components: sentence-level classification (§3.2.1), domain knowledge retrieval (§3.2.2), and dual-stage QA-based factual consistency evaluation (§3.2.3).
+3.2.1 Learned Factuality Type Classifier
+As elaborative explanations are prevalent in plain language generation and existing metrics struggle to capture added information [12], we first fine-tune a pre-trained language model to categorize factuality types of plain language sentences.
+Based on the factuality type annotations (source simplification vs.
+elaborative explanation) in PlainFact, we fine-tune the PubMedBERT-base model [^3] https://huggingface.co/microsoft/BiomedNLP-BiomedBERT-base-uncased-abstract-fulltext as a classifier.
+Our PlainFact is split 8:1:1 for training, validation, and testing.
+Additionally, we compare the pre-trained classifier with GPT-4o [34] as a zero-shot classifier (prompts are provided in C) to explore extending our factual consistency evaluation metric to domains and tasks lacking human-annotated data.
+3.2.2 Domain Knowledge Retrieval
+Retrieval-augmented methods have proven effective for explanation generation [28, 35], making them a natural fit for evaluating elaborative explanations.
+Since our dataset is in the medical domain, we employ MedCPT [36], a retriever pre-trained on large-scale PubMed search logs to generate biomedical text embeddings.
+For external resources, we incorporate StatPearls [37] for clinical decision support and medical textbooks [38] for domain-specific knowledge.
+3.2.3 QA Evaluation Components
+QA-based metrics have proved to be effective than other factual consistency evaluation metrics [12] in general summarization tasks and align more closely with human annotations [20, 39].
+In this study, we adopt a QA-based approach as a backbone of our evaluation metric to verify factual consistency, incorporating answer extraction, question generation, question filtering, question answering, and answer overlap calculation modules.
 Gold Answer Extraction
-The first step in QA-based factuality evaluation is to extract answer entities (keyphrases) from the PLS as gold answer, then verify factuality by comparing them with answers generated by a QA model for the same questions.
+The first step in QA-based factual consistency evaluation is to extract answer entities (keyphrases) from plain language summaries as gold answer, then verify factuality by comparing them with answers generated by a QA model for the same questions.
 If the generated answer is correct or relevant, the summary is considered factual.
-To compare answer extraction strategies, we use noun phrase chunks (NP Chunks) as the baseline, and evaluate LLM-based extractors, including Llama 3.1 8B Instruct (Llama 3.1)Meta (2024), GPT-4o OpenAI (2024), and PromptRank Kong et al.
-(2023).
-Prompts for Llama and GPT-4o are in App.
-C.
-PromptRank is a keyphrase generation method based on the T5 model.
-The number of extracted entities is set to 50 and the T5-large is used for better performance.
+The previous study used PromptRank [40], a keyphrase generation method based on the T5 model.
+To compare answer extraction strategies, we use PromptRank as the baseline, and evaluate LLM-based extractors, including an open-source Llama 3.1 8B Instruct model (Llama 3.1) [41] and close-source GPT-4o (GPT-4o) [34].
+Prompts for Llama 3.1 and GPT-4o are in C.
 Question Generation (QG)
-Given an input document (PLS), the QG model generates questions based on the extracted answers and their context.
-Following previous studies Deutsch et al.
-(2021); Fabbri et al.
-(2022b) [^4] These studies use allennlp, an open-source NLP research library built on PyTorch, which is no longer actively maintained., we fine-tune BART-large model on standard QG datasets, including SQuAD Rajpurkar et al.
-(2016) and QA2D Demszky et al.
-(2018), for use as the QG model in our framework.
+Given an input plain language summary, the QG model generates questions based on the extracted answers (§3.2.3) and the input plain language summary.
+Following previous studies [20, 21] [^4] These studies use allennlp, an open-source NLP research library built on PyTorch, which is no longer actively maintained., we fine-tune BART-large model on standard QG datasets, including SQuAD [42] and QA2D [43], for use as the QG model in our evaluation metric.
+The QG model will generate multiple relevant questions based on the input, which will increase the probability of verifying the factual consistency by answering these questions in the following stages.
 Question Filtering (QF)
-To prevent unanswerable questions from impacting the final factuality score, we follow QAFactEval Fabbri et al.
-(2022a) and remove the unanswerable questions using a pre-trained Electra-large model Clark et al.
-.
-During QF, the QA model receives only the PLS and its corresponding questions.
-In the subsequent QA stage, answers are extracted from the source for answer overlap evaluation, while QF only determines whether the questions can be answered by the PLS.
+Questions generated by the QG model (§3.2.3) may not always be answerable.
+For example, the QG model generates “<What can occur over several days?>,” which is vague and hard to predict the correct answer.
+To prevent these unanswerable questions from impacting the final evaluation performance, we follow QAFactEval [39] and remove the unanswerable questions using a pre-trained Electra-large model [44].
+During QF, the filtering model receives only the plain language summary and its corresponding questions.
+In the subsequent QA stage (§3.2.3), answers are extracted from the source for answer overlap evaluation, while QF only determines whether the questions can be answered by the plain language summary.
 Question Answering
-The QA model extracts answers to filtered questions from the source document.
-To prevent hallucinated output, we use an extractive QA model, a pre-trained Electra-large, which was the best performing QA model in QAFactEval Fabbri et al.
-(2022a).
+The QA model extracts answers to answer the filtered questions from the source document.
+To prevent hallucinated output, we use an extractive QA model, a pre-trained Electra-large, which was the best performing QA model in QAFactEval [39].
 Answer Overlap Evaluation (AOE)
-We evaluate the alignment between the gold and generated answers from the QA model using the LERC score Chen et al.
-(2020) and BERTScore Zhang et al.
-.
-The final step in PlainQAFact is weighting the scores for source simplification and elaborative explanation sentences.
-Specifically, the simplification score $s$ is computed using only abstracts as the source in QA, while the explanation score $e$ incorporates both abstracts and retrieved knowledge.
+We evaluate the alignment between the gold and generated answers from the QA model using the baseline Learned Evaluation metric for Reading Comprehension (LERC) score [45] and BERTScore [46].
+The final step in PlainQAFact is weighting the factual consistency scores for source simplification and elaborative explanation sentences.
+Specifically, the simplification score $s$ is computed using only abstracts as the source in QA, while the explanation score $e$ is calculated by incorporating both abstracts and retrieved knowledge as source for QA-based evaluation.
 $\textsc{PlainQAFact}=\frac{s_{\text{Avg.}}\cdot n_{s}+e_{\text{Avg.}}\cdot n_{e}}{n_{s}+n_{e}},$ (1)
 where $n_{s}$ is the number of simplification sentences, and $n_{e}$ is the number labeled as explanation.
+$s_{\text{Avg.}}$ and $e_{\text{Avg.}}$ denote the average PlainQAFact scores computed for instances classified as simplification and explanation sentences, respectively.
 4 Experiments
-Since the classifier is trained on the PlainFact training set and its test set is limited to 20 scientific-PLS pairs, we conduct experiments on both PlainFact and the CELLS test set, excluding PlainFact to prevent data leakage.
+We conduct main experiments on three publicly available datasets, including PlainFact, CELLS [28], and FareBio [47].
+To verify the effectiveness of our proposed PlainQAFact on PLS tasks, we compare with five widely used factual consistency evaluation metrics and two LLM-based evaluators.
+4.1 Datasets
+As introduced in Sec 3.1, we create PlainFact with fine-grained sentence-level annotations regarding the factuality types.
+Additionally, to test the generalizability of our pre-trained classifier, we also collect another subset from the test set of CELLS dataset [28], gathering 200 summary-abstract pairs.
+Similar to PlainFact3.1.1), we also conduct sentence-level perturbation for each plain language summary in our collected CELLS dataset.
+Therefore, we have 200 factual and 200 non-factual summary-abstract pairs for both PlainFact and CELLS datasets.
+We also compare PlainQAFact on the FareBio dataset, which contains 175 plain language summaries generated by seven different LLMs, together with the source articles [47].
+Specifically, FareBio includes 25 full scientific article examples, each source article is summarized by seven LLMs, resulting 175 plain language summaries in total.
+We sample plain language sentences that are not both faithful and factual as non-factual sentences, and the rest as factual sentences, resulting in 33 and 174 summaries respectively.
+We also extract sentences that are labeled as not faithful but factual in their “factual hallucination” label, which represents the plain language sentence that is not from the source article, but still conveying factual information, resulting in 53 valid summaries in total.
+These sentences are equal to our factual elaborative explanation sentences.
+Note in FareBio dataset, there are no non-factual simplification sentences.
+Therefore, we use the same set of non-factual sentences (33 sentences) for main experiment (Table 2) and explanation-only (Figure 2) evaluation.
 4.2 Experiment Settings
-Detailed model selections for each module are in §4.
-Prompts for GPT-4o OpenAI (2024) and Llama 3.1 8B Instruct Meta (2024) are in App.
-C.
-For the QG model, we fine-tune the BART-large with a batch size of 16, a learning rate of 3e-5, and two training epochs.
-In MedCPT retrieval, we set k (the number of retrieved snippets) to 3, ensuring the retrieved information remains within a short context window for the subsequent QA process (details in App.
-D).
+The classifier used in PlainQAFact is fine-tuned on PlainFact with a batch size of 32 for 10 epochs.
+We apply early stopping during training based on validation loss.
+The temperature is set to 0 for GPT-4o and 0.01 for Llama 3.1.
+The Llama 3.1 model used in AE also uses a temperature of 0.01.
+For the QG model, we fine-tune BART-large with a batch size of 16, a learning rate of 3e-5, and two training epochs.
+The maximum input length is set to 512 for all models in PlainQAFact.
+More details of prompts and model hyper-parameters for GPT-4o and Llama 3.1 are provided in C and  D.
 4.3 Existing Factuality Metrics
-To the best of our knowledge, no prior work has efficiently evaluated factuality metrics for detecting elaborative explanations in PLS.
+To the best of our knowledge, no prior work has efficiently evaluated factuality metrics for detecting elaborative explanations in PLS tasks.
 Moreover, most factuality metrics are designed for general-domain applications, largely due to limited quality and annotation of existing PLS datasets.
-To address this gap, we incorporate the following metrics in our experiment, informed by prior work Joseph et al.
-(2024): (1) Dependency-Arc Entailment (DAE) Goyal and Durrett (2021) is an entailment-based method that evaluates summary factuality by breaking it into smaller entailment tasks at the arc level.
+To address this gap, we incorporate the following metrics in our experiment, informed by prior work [9]: (1) Dependency-Arc Entailment (DAE) [48] is an entailment-based method that evaluates summary factuality by breaking it into smaller entailment tasks at the arc level.
 The model independently determines whether each arc’s relationship is supported by the input.
-(2) AlignScore Zha et al.
-(2023) is a model-based factuality metric using RoBERTa Liu (2019).
+(2) AlignScore [19] is a model-based factuality metric using RoBERTa [49].
 It extracts claims from the summary and calculates the alignment scores with all the context chunks from the input document.
 The final score is an average of all highest alignment probabilities.
-(3) SummaC Laban et al.
-(2022) is an NLI-based inconsistency detection method designed for summarization tasks.
+(3) SummaC [16] is an NLI-based inconsistency detection method designed for summarization tasks.
 It segments documents into sentences and aggregates scores between sentence pairs using a trained NLI model.
-(4) QAFactEval Fabbri et al.
-(2022a) is a QA-based factuality evaluation metric that assesses summary consistency by generating questions from the source document and comparing the model-generated answers with expected answers.
-(5) QuestEval Scialom et al.
-(2021a) is a reference-less evaluation metric employs a T5 model to generate questions from the source document and verifies whether the summary can correctly answer them.
+(4) QAFactEval [39] is a QA-based factuality evaluation metric that assesses summary consistency by generating questions from the source document and comparing the model-generated answers with expected answers.
+(5) QuestEval [50] is a reference-less evaluation metric employs a T5 model to generate questions from the source document and verifies whether the summary can correctly answer them.
 5 Results and Analysis
-We first analyze the impact of each component incorporated into PlainQAFact and conduct additional ablation studies to assess retrieval effectiveness on PlainFact6.1).
-Next, we evaluate PlainQAFact on two datasets, comparing it with five widely used factuality metrics (§6.2).
-Finally, we perform an error analysis to investigate the failure cases of PlainQAFact6.5).
-Component Method Choice Benchmark
-PlainQAFact 56.0 (0.02)
-Classifier Learned classifier [^5] We acknowledge that applying the learned classifier to its training data may introduce data leakage.
-To further validate PlainQAFact, we evaluate its performance on the CELLS dataset, as shown in Figure 2.
--
-GPT-4o 56.1 (0.02)
-No (retrieve for all) 58.4* (0.03)
-Answer Extraction Llama3.1 8B Instruct -
-GPT-4o 56.8 (0.05)
-PromptRank 54.1*
-NP Chunks 53.9*
-Retrieval Source Abs + TB + SP -
-TB + SP 48.5* (0.02)
-Abs (no retrieval) 50.6* (0.03)
-Answer Overlap BERTScore -
-LERC 28.9* (0.03)
-Granular Level Sentence -
-Summary 55.8 (0.06)
-Table 2: Ablation study on PlainFact, evaluating the impact of individual components in PlainQAFact.
-The first row represents our proposed combination.
-Ablations are conducted by replacing one component at a time while keeping all others fixed to match the best overall model.
+We first conduct a pilot study on the FactPICO dataset [9] to investigate the performance of existing factual consistency evaluation metrics on PLS tasks.
+Our study (G) reveals two major limitations.
+First, many LLM-generated plain language summaries in FactPICO contain incomplete or incoherent sentences, highlighting the need for a benchmark based on high-quality, human-authored plain language summaries.
+Second, existing automatic factual consistency metrics are not sensitive to evaluate the factuality of “added information”, underscoring the need for a more domain-related factual consistency evaluation metric, especially for those summaries with additional explanations that are not mentioned by the original source texts.
+To address these challenges, we introduce a new expert-annotated dataset, PlainFact, and propose PlainQAFact, a QA-based, retrieval-augmented metric for evaluating factual consistency in plain language summaries.
+In this section, we first benchmark PlainQAFact against five widely used automatic factual consistency evaluation metrics on three PLS datasets (§5.1).
+We then evaluate its performance in an explanation-only setting using PlainFact5.2), followed by an ablation study to assess the contribution of each component (§5.3).
+Next, we conduct an error analysis to examine failure cases and highlight directions for further improvement (§5.4).
+Datasets Metrics Kendall’s $\tau$ Pearson AUC-ROC
+CELLS Llama 3.1 57.5 42.1 83.4
+GPT-4o 70.0 75.4 99.3
+QAFactEval 61.6 70.6 93.5
+QuestEval 23.2 28.3 66.4*
+SummaC 24.8 29.8 67.5*
+AlignScore 56.0 66.2 89.6
+DAE 14.4 14.4 55.0*
+PlainQAFact 62.1 75.2 93.8
+FareBio Llama 3.1 40.5 38.9 83.0
+GPT-4o 27.3 35.7 75.7
+QAFactEval 32.8 43.9 81.6
+QuestEval 37.8 45.2 86.4
+SummaC 33.8 32.1 82.5
+AlignScore 32.2 47.2 81.0
+DAE 8.8 8.8 53.4
+PlainQAFact 16.8 35.6 66.1
+PlainFact Llama 3.1 60.7 52.6 85.3
+GPT-4o 69.8 80.3 99.2
+QAFactEval 65.8 79.3 96.4
+QuestEval 28.4 34.7 70.1*
+SummaC 34.1 42.8 74.1*
+AlignScore 60.2 72.6 92.5
+DAE 5.3 5.3 51.8*
+PlainQAFact 65.7 81.3 96.4
+Table 2: Evaluation results of automatic metrics on the CELLS [28], PlainFact, and FareBio [47] datasets.
+Results are evaluated using Kendall’s $\tau$, Pearson correlation, and AUC-ROC, with eight metric scores compared against human-labeled factuality.
+The standard deviations (std.) over five runs are: 0.2 (PlainQAFact), 0.5 (Llama 3.1), and 6.0 (GPT-4o) on PlainFact; 0.1 (PlainQAFact), 0.2 (Llama 3.1), and 6.3 (GPT-4o) on CELLS; and 0.1 (PlainQAFact), 2.9 (Llama 3.1), and 17.0 (GPT-4o) on FareBio.
+* indicates an improvement of our metric over prior work with 95% confidence interval (details in  E).
+Note the plain language summaries in FareBio are generated by seven LLMs, while CELLS and PlainFact are written by the authors of original articles.
+5.1 Main Results
+We report rank correlation (Kendall’s $\tau$), linear correlation (Pearson), and discrimination (AUC-ROC [51]) for three datasets (PlainFact, CELLS [28], and FareBio [47]) following previous studies [16, 39, 49] in Table 2.
+Overall, GPT-4o performs the best on CELLS and PlainFact, but its performance drops on FareBio.
+DAE is consistently the weakest across all datasets and metrics, suggesting that a dependency-only metric is insufficient for capturing the factual consistency required in PLS tasks.
+On CELLS, PlainQAFact outperforms all other automatic metrics on all three criteria (Kendall’s $\tau$=62.1, Pearson=75.2, AUC-ROC=93.8).
+This shows that our approach, first detecting elaborative content and then verifying it with retrieval, works better than existing QA-based metrics, which ask all questions from the source only, and also better than NLI-based evaluation metrics.
+Similarly, on PlainFact, PlainQAFact achieves the highest Pearson (81.3), surpassing GPT-4o by 1.0 point.
+It ties AUC-ROC (96.4) score with QAFactEval and has a competitive Kendall’s $\tau$ (65.7).
+These results on both datasets show that PlainQAFact is reliable for evaluating both factual and non-factual plain language summaries.
+On FareBio, no single method performs the best on all criteria.
+QuestEval achieves the best AUC-ROC (86.4), AlignScore has the best Pearson (47.2), while Llama 3.1 reaches the best Kendall’s $\tau$ (40.5).
+However, FareBio differs from the other two datasets in several ways: (1) the original labels in FareBio are defined at the sentence level, and a summary is marked as non-factual if any sentence is non-factual; (2) their summaries are generated from full scientific articles rather than abstracts, making it harder for metrics to locate matching source text; (3) the dataset is imbalanced (33 non-factual vs.
+174 factual); and (4) All plain language summaries are generated by LLMs and no elaborative explanation annotations are provided for each sentence or summary.
+These factors likely increase evaluation difficulty on FareBio and impact the performance of multiple metrics, including PlainQAFact and GPT-4o.
+In summary, PlainQAFact is a more consistent and effective metric for factual consistency compared to existing automatic approaches.
+It clearly improves over prior QA-based metrics on CELLS and matches or exceeds them on PlainFact.
+Although GPT-4o performs better in most of the settings, our PlainQAFact is built entirely on the open-source LLM (i.e., Llama 3.1), ensuring transparency, reproducibility, and accessibility.
+Its two-step evaluation process, detecting explanations and verifying them with retrieval, helps align metric scores more closely with human judgments in PLS evaluation tasks.
+Figure 2: Overall performance on human-annotated elaborative explanation summaries from PlainFact (392 summaries).
+The std.
+of PlainQAFact, Llama 3.1, and GPT-4o are 0.1, 1.0, and 7.7, respectively based on five runs for each metric.
+* indicates an improvement of our metric over prior work with 95% confidence interval (details in  E).
+PlainQAFact significantly outperforms most of the automatic factual consistency evaluation metrics in AUC-ROC.
+Note that the CELLS dataset does not contain annotations for elaborative explanations.
+Results of explanation-only evaluation on FactPICO and FareBio are reported in F.
+5.2 Explanation-Only Evaluation
+To test whether existing evaluation metrics are limited in assessing elaborative explanations in PLS, we evaluate PlainQAFact on human-annotated explanation-only summaries from PlainFact.
+We select only the sentences labeled as “elaborative explanation” and group them into plain language summaries, resulting in 392 summary–abstract pairs for PlainFact.
+Note that CELLS does not provide sentence-level annotations for elaborative explanations, so we evaluate it only as a general PLS dataset (§5.1).
+As shown in Figure 2, PlainQAFact outperforms all existing automatic metrics across all three criteria.
+It is 5.7 points higher in Kendall’s $\tau$, 7.5 points higher in Pearson, and 4.0 points higher in AUC-ROC compared with the second-best metric, QAFactEval.
+These results indicate that PlainQAFact captures the factual consistency of elaborative explanations more effectively than prior automatic metrics.
+While GPT-4o achieves higher overall scores (Kendall’s $\tau$=54.9, Pearson=60.7, AUC-ROC=88.7), it shows much higher variance (std.
+7.7) compared to PlainQAFact (std.
+0.1), making it less stable for factual consistency evaluation.
+Compared to the results in Table 2, AlignScore is competitive on general PLS tasks but drops on evaluating explanation-only sentences.
+In contrast, PlainQAFact remains consistently strong when we isolate explanation-only content.
+PlainQAFact detects added elaborative explanations and checks them with domain knowledge retrieval, which helps when factual consistency depends on facts beyond simple restatement from the source.
+Since 44% of sentences in our curated benchmark are labeled as elaborative by human annotators, we believe that PlainQAFact is the more suitable and robust factual consistency evaluation metric in this scenario.
+PlainQAFact also outperforms Llama 3.1 on all three criteria (27.9/28.3/67.0), showing that instruction-tuned LLM judges alone are not enough for reliable evaluation of elaborative content.
+Component Method Choice Kendall’s $\tau$ Pearson AUC std.
+PlainQAFact 65.7 81.3 96.4 0.2
+Classifier Fine-tuned classifier - - - -
+GPT-4o 62.8 77.1 94.3* 0.1
+No (retrieve for all) 61.2 74.2 93.2* 0.7
+Answer Extraction Llama 3.1 - - - -
+GPT-4o 66.9 81.9 97.2 0.6
+PromptRank 66.3 80.2 96.0 -
+Retrieval Source Abs + TB + SP - - - -
+TB + SP 63.4 76.8 94.8 0.1
+Abs (no retrieval) 62.8 77.1 94.3* 0.1
+Answer Overlap BERTScore - - - -
+LERC 64.5 75.9 95.5 0.1
+Granular Level Sentence - - - -
+Summary 49.3 59.6 84.7* 0.5
+Table 3: Ablation study of PlainQAFact on PlainFact, analyzing the impact of individual components in PlainQAFact.
 Abs: abstract; TB: Textbooks; SP: StatPearls.
-Using the first row as the baseline, * indicates a statistically significant difference based on a paired t-test ($p<0.05$).
+AUC represents AUC-ROC.
+The first row of each component setting represents our best combination.
+* indicates an improvement of our metric over prior work with 95% confidence interval.
 We run each setting for five times and report the standard deviation (std.).
@@
 5.3 Ablation Study
-Table 2 is the sentence-level ablation study of PlainQAFact where each component of our evaluation metric is modified individually.
-See App.
-E for the summary-level evaluation.
-Additional analysis of retrieval resources is presented in Table 3.
-Learned Classifier
-Evaluating all PLS sentences without classification and retrieving for all improves factuality assessment but significantly increase evaluation time.
-Retrieval for every sentence extends processing time from 2 hours and 40 minutes with our learned classifier to 6 hours and 30 minutes without it for 2,740 evaluation examples.
-This highlights the efficiency gains of incorporating a classifier in PLS factuality evaluation.
-Notably, the GPT-4o classifier achieves results comparable to our learned classifier trained on expert-annotated data, demonstrating the effectiveness of LLMs as alternative classifiers.
-This extends the applicability of PlainQAFact to factuality evaluation in other domains, as LLMs can serve as classifiers without requiring costly human annotations.
-Llama for Gold Answer Extraction
-The Llama-based answer extraction method outperforms NP Chunks by 2.1% by providing context-aware, flexible boundaries and robust answer selection Yu et al.
-(2024), making it more effective for factuality evaluation.
-Although GPT-4o achieves slightly better performance, the improvement is marginal, making Llama a more cost-effective alternative.
-Sentence Type Source Reference Benchmark
-Simplification Abs 63.3 (0.05)
-Explanation Abs 30.5* (0.03)
-TB 17.0* (0.02)
-SP 18.4* (0.03)
-TB + SP 25.2* (0.02)
-Abs + TB 39.8* (0.08)
-Abs + SP 40.1* (0.03)
-Abs + TB + SP 44.6 (0.06)
-Full PLS Abs 50.6* (0.03)
-TB 45.4* (0.02)
-SP 45.9* (0.03)
-TB + SP 48.5* (0.02)
-Abs + TB 54.1* (0.02)
-Abs + SP 54.3* (0.01)
-Abs + TB + SP 56.0 (0.02)
+PlainQAFact consists of several modules, including a fine-tuned classifier, QA modules, a retrieval function, and an answer overlap evaluation process.
+To understand the contribution of each component, we conduct an ablation study on PlainFact5.3).
+Table 3 summarizes the ablation results, where each component of PlainQAFact is individually modified to measure its impact on overall performance.
+Fine-tuned Classifier Table 3 shows that removing the classifier and retrieving for every sentence does not improve factual consistency assessment.
+Performance drops across all criteria (Kendall’s $\tau$ from 65.7 to 61.2; Pearson from 81.3 to 74.2; AUC-ROC from 96.4 to 93.2), and the std.
+increases to 0.7.
+This approach also requires notably more computation.
+Using GPT-4o as a simple classifier provides reasonable results but still underperforms compared to our fine-tuned classifier.
+These findings support the use of a lightweight fine-tuned classifier to trigger retrieval only when necessary.
+Answer Extraction For extracting gold answers from plain language summaries, the LLM-based extractors are better than PromptRank [40].
+GPT-4o achieves the best overall scores (Kendall’s $\tau$=66.9, Pearson=81.9, AUC-ROC=97.2), with small gains over our Llama-based method.
+PromptRank performs better on Kendall’s $\tau$ (increases 0.6), but fails on other two criteria compared to our best combination.
+Although GPT-4o achieves slightly higher performance, the marginal gain does not justify the increased cost, making Llama 3.1 a more practical and cost-effective choice for large-scale factual consistency evaluation.
+Component Method Choice Kendall’s $\tau$ Pearson AUC std.
+Source Simplification Abs 63.5 79.0 94.8 0.2
+Elaborative Explanation Abs 32.8 38.2 72.8* 0.2
+TB 31.2 37.2 70.6* 0.1
+SP 33.3 37.8 72.5 1.7
+TB + SP 36.7 42.0 75.4 0.1
+Abs + TB 40.3 47.0 78.2 0.2
+Abs + SP 40.5 47.4 78.3 0.2
+Abs + TB + SP 44.2 51.3 81.0 0.1
+Full Dataset Abs 62.8 77.1 94.3* 0.1
+TB 60.8 73.8 93.0* 0.1
+SP 61.3 74.2 93.3* 0.1
+TB + SP 63.4 76.8 94.8 0.1
+Abs + TB 64.3 79.7 95.4 0.2
+Abs + SP 65.0 80.3 95.9 0.2
+Abs + TB + SP 65.7 81.3 96.4 0.2
 Table 4: Ablation study results on the PlainFact, evaluating how different retrieval sources affect various information types using PlainQAFact.
-The learned classifier categorizes sentences as source simplification (1,685) or elaborativeexplanation (1,055).
+The fine-tuned classifier categorizes input sentences as source simplification or elaborative explanation.
+Overall, the numbers of factual summaries that only include source simplification and elaborative explanation sentences are 198 and 191 respectively, and 198 and 194 for non-factual summaries.
 Abs: abstracts; TB: Textbooks; SP: StatPearls.
-Using the Abs+TB+SP as the baseline, * indicates a statistically significant difference based on a paired t-test ($p<0.05$).
+AUC represents AUC-ROC.
+* indicates an improvement of Abs + TB + SP over other settings with 95% confidence interval.
 We run each setting for five times and report the standard deviation (std.) in the brackets.
-Combined Domain Resources for Retrieval
-The best performance is achieved when abstracts, medical Textbooks, and StatPearls are combined as the source document.
-In contrast, excluding the source abstract for QA-based factuality evaluation achieves the worst performance, which is expected given that abstracts contain most of the key information.
-We further examine the impact of retrieval resources on different information types.
-As shown in Table 3, for elaborative explanation sentences, using abstracts alone leads to significantly lower scores compared to source simplification sentences, indicating that abstracts alone are insufficient for verifying external elaborative explanation content.
-Incorporating StatPearls alongside abstracts substantially improves performance for both elaborative explanation and full PLS, underscoring the importance of integrating high-quality domain-specific resources for factuality verification.
-BERTScore for Answer Overlap
-Although LERC performs well in QAFactEval, BERTScore achieves significantly better results for PLS factuality evaluation, improving performance by 27.1% on PlainFact.
-This disparity arises from their fundamental differences: LERC is a learned metric trained on human factual consistency scores, whereas BERTScore relies on embedding-based similarity.
-Given that PlainQAFact evaluates factuality by comparing generated answers with extracted gold answers from the source, BERTScore’s ability to measure fine-grained semantic similarity makes it more effective choice.
-Sentence-level Evaluation
-Sentence-level evaluation slightly outperforms summary-level evaluation, primarily due to differences in input context for the QG model.
-In sentence-level evaluation, the QG model processes one PLS sentence at a time, enabling more targeted and precise factual verification.
-In contrast, summary-level evaluation takes the entire PLS as input, introducing additional contextual dependencies that can lead to factual inconsistencies.
-This further confirms the effectiveness of finer granularity in PLS evaluation.
-Figure 2: Scaled factuality accuracy on PlainFact and the CELLS test set.
-PlainQAFact significantly outperforms all other metrics on both datasets, except for AlignScore on PlainFact ($p<0.05$, paired t-test).
-We compute the std.
-by running the experiment five times for both our method and Llama 3.1.
-6.2 PlainQAFact Achieves the Best Performance in PLS Factuality Evaluation
-Figure 2 compares existing factuality metrics with PlainQAFact on the PlainFact and CELLS datasets.
-To ensure a fair comparison, we scale factuality accuracy to a range of 0 to 1, as QAFactEval scores range from 1 to 5.
-To further assess generalizability, we evaluate on 200 PLS-abstract pairs randomly sampled from the CELLS test set, sourced from journals outside PlainFact.
-Results indicate that AlignScore’s effectiveness depends on the proportion of elaborative versus simplification sentences.
-When fewer external elaborative sentences are present, its performance approaches that of PlainQAFact.
-However, as elaborative content increases, PlainQAFact consistently outperforms AlignScore, demonstrating its ability to better handle elaborative information absent from the source abstract.
-Compared to QA-based metrics such as QAFactEval and QuestEval, PlainQAFact benefits from incorporating retrieval module, which enables a more comprehensive factuality assessment for elaborative explanations.
-DAE consistently performs the worst in PlainFact settings, highlighting the limitations of dependency-based factuality evaluation for PLS.
-Additionally, we assess the performance of Llama 3.1 as a judge for factuality evaluation.
-In summary-level evaluations, Llama 3.1 demonstrates superior performance across both the PlainFact and CELLS datasets, showing its robust capability for context-aware factuality assessment in plain language summarization tasks.
-However, its performance at the sentence level is inconsistent, while it slightly outperforms PlainQAFact on the CELLS dataset, it decreases by 16.5% on the PlainFact dataset (see App.
-C for the prompt used with Llama 3.1).
-Figure 3: Scaled factuality accuracy on human-annotated elaborative explanation sentences from PlainFact.
-PlainQAFact significantly outperforms all other metrics on both evaluation levels ($p<0.05$, paired t-test).
-We compute the std.
-by running the experiment five times for both our method and Llama 3.1.
-Misclassification Plain Language Sentence: This review looked at how well the methods worked to prevent pregnancy[extracted ${}_{\textit{[extracted answer]}}$, if they caused bleeding problems, if women used them as prescribed, and how safe they were.
-Lopez et al.
-(2013)
-Scientific Abstract: To compare the contraceptive effectiveness, cycle control, compliance (adherence), and safety of the contraceptive patch or the vaginal ring versus combination oral contraceptives (COCs)… Lopez et al.
-(2013)
+Combined Domain Resources for Retrieval The highest performance is achieved when combining all three sources (Abs + TB + SP).
+This suggests that comprehensive retrieval, which includes the original source abstracts, improves overall factual consistency assessment.
+Compared to using only external sources (TB + SP), performance drops by 2.3/4.5/1.6 across the three criteria.
+Using only abstracts (i.e., no elaborative retrieval) also lowers performance (2.9/4.2/2.1).
+To further examine the effectiveness of retrieval for source simplification and elaborative explanation, Table 4 breaks down the results as follows:
+(1) Source simplification: These are sentences in plain language summaries classified as “source simplification.” We evaluate summaries containing only simplification sentences, using abstracts as the source.
+Using abstracts as sole source texts achieves strong performance (63.5/79.0/94.8) across all three criteria.
+(2) Elaborative explanation: For summaries containing only explanation sentences, abstracts alone are not sufficient (32.8/38.2/72.8), performing much worse than in simplification cases.
+Adding abstracts with external sources improves performance, with the best results obtained by combining all three sources (Abs + TB + SP).
+(3)Full dataset: Retrieval from StatPearls provides better results than retrieval from Textbooks, highlighting the importance of using high-quality, domain-specific knowledge bases for PLS evaluation.
+Overall, the best combination across all settings is Abs + TB + SP (65.7/81.3/96.4).
+These findings show that effective retrieval for factual consistency should include both the source abstract and reliable external medical references, especially when summaries contain elaborative explanations.
+Answer Overlap Evaluation BERTScore outperforms the LERC approach introduced in QAFactEval [21], with gains of 1.2/5.4/0.9 across the three evaluation criteria.
+This performance gap can be attributed to the fundamental differences between the two metrics.
+LERC is a learned model trained to predict human factual consistency scores, which may reduce its generalizability to new domains or sentence types.
+In contrast, BERTScore computes embedding-based semantic similarity, allowing it to capture fine-grained semantic overlap between generated and gold answers.
+Since PlainQAFact evaluates factual consistency by comparing model-generated answers to those extracted from the source, BERTScore’s sensitivity to semantic alignment makes it a more effective and robust choice for this step of the evaluation.
+Sentence-level vs.
+Summary-level Evaluation To further examine the effect of input granularity in factual consistency evaluation, we deactivate the sentence-splitting function and instead pair each plain language summary with an abstract to evaluate factual consistency at the summary level.
+Our best sentence-level evaluation achieves an AUC-ROC of 96.4, while the summary-level setting performs worse across all three criteria (Kendall’s $\tau$=49.3, Pearson=59.6, AUC-ROC=84.7; std.
+0.5).
+Evaluating at the summary level introduces broader contextual dependencies, which can omit sentence-specific factual errors or create inaccurate entailments.
+These results suggest that processing one sentence at a time makes QA-based question generation and answering more focused and reduces noise from unrelated context.
+Correct Case Plain Language Sentence: Limits to the availability of SSB in schools (e.g. replacing SSBs with water in school cafeterias). [52]
+Scientific Abstract: Frequent consumption of excess amounts of sugar-sweetened beverages (SSB) is a risk factor for obesity, type 2 diabetes, cardiovascular disease and dental caries… [52]
+Generated Question: Limits to the availability of what in schools?
+Retrieved Knowledge: Even though additional data is required to determine the impact of limiting the availability of nutrient-poor or high-sugar goods in schools on obesity, some study results have shown a net-positive result…
+Final Answer: nutrient-poor or high-sugar goods
+Misclassification Plain Language Sentence: This review looked at how well the methods worked to prevent pregnancy, if they caused bleeding problems, if women used them as prescribed, and how safe they were.
+[53]
+Scientific Abstract: To compare the contraceptive effectiveness, cycle control, compliance (adherence), and safety of the contraceptive patch or the vaginal ring versus combination oral contraceptives (COCs)… [53]
 Generated Question: This review looked at how well the methods worked to prevent what?
 Retrieved Knowledge: Appropriate treatment for the underlying etiology should start as soon as possible, and the patients and family members should receive appropriately targeted education…
-QA Answer: cycle control
-Retrieval Quality Plain Language Sentence: The patch is a small, thin, adhesive[extracted ${}_{\textit{[extracted answer]}}$ square that is applied to the skin. Lopez et al. (2013)
+Final Answer: cycle control
+Retrieval Quality Plain Language Sentence: The patch is a small, thin, adhesive square that is applied to the skin. [53]
 Scientific Abstract: Users of the norelgestromin-containing patch reported more breast discomfort, dysmenorrhea, nausea, and vomiting.
-In the levonorgestrel-containing patch trial, patch users reported less vomiting, headaches, and fatigue…Lopez et al.
-(2013)
+In the levonorgestrel-containing patch trial, patch users reported less vomiting, headaches, and fatigue…[53]
 Generated Question: The patch is a small, thin, what kind of square applied to the skin?
@@
 This may be sutured or taped securely in place to provide some pressure and to keep graft immobilized.
-QA Answer: Nonstick dressing Petrolatum-infused gauze strip
-Unanswerable Question Plain Language Sentence: Government[extracted ${}_{\textit{[extracted answer]}}$ officials, business people and health professionals implementing such measures should work together with researchers to find out more about their effects in the short and long term.
-von Philipsborn et al.
-(2019)
-Scientific Abstract: To assess the effects of environmental interventions (excluding taxation) on the consumption of sugar-sweetened beverages and sugar-sweetened milk, diet-related anthropometric measures and health outcomes, and on any reported unintended consequences or adverse outcomes…von Philipsborn et al.
-(2019)
+Final Answer: Nonstick dressing Petrolatum-infused gauze strip
+Unanswerable Question Plain Language Sentence: Government officials, business people and health professionals implementing such measures should work together with researchers to find out more about their effects in the short and long term.
+[52]
+Scientific Abstract: To assess the effects of environmental interventions (excluding taxation) on the consumption of sugar-sweetened beverages and sugar-sweetened milk, diet-related anthropometric measures and health outcomes, and on any reported unintended consequences or adverse outcomes…[52]
 Generated Question: What officials, business people and health professionals implementing such measures should work together with researchers to find out more about their effects in the short and long term?
 Retrieved Knowledge: Implementation should be accompanied by high-quality evaluations using appropriate study designs, with a particular focus on the long-term effects of approaches suitable for large-scale implementation.
-QA Answer: long-term effects
-Table 4: Error analysis of PlainQAFact with model intermediate outputs.
-Each plain language sentence is extracted from the PLS, the QG model generates questions based on the extracted answer and PLS sentence, and "Retrieved Knowledge" refers to the context used by the QA model for answer extraction.
-6.3 Explanation-Only Evaluation
-We evaluate PlainQAFact on human-annotated explanation-only sentences from PlainFact (see Figure 3), conducting both sentence- and summary-level evaluations.
-For summary-level setting, we aggregate all explanation sentences from each PLS to form a new summary, resulting in 196 summaries in total (4 PLS are labeled as all simplification sentences.).
-Results show PlainQAFact outperforms all other factuality metrics in assessing explanation information at both evaluation levels, surpassing the second-best metric by 12.2% at the sentence-level and 10.5% at the summary-level.
-These findings highlight PlainQAFact is effective in evaluating the factuality of elaborative explanations, particularly when external knowledge is required for verification beyond the scientific abstracts.
-While Llama 3.1 demonstrates superior performance at the summary-level evaluation, PlainQAFact emerges as the best performer in sentence-level setting.
-Compared to the results in Figure 2, AlignScore performs similarly to PlainQAFact overall on PlainFact but struggles with external sentences, where its performance drops significantly.
-Thus, PlainQAFact proves to be the most suitable method for the factuality evaluation of PLS.
-Figure 4: Evaluation time comparison of PlainQAFact.
-We compare both sentence- and summary-level evaluation time over four settings: without information retrieval, PlainQAFact, and using GPT-4o as the classifier and answer extractor.
-We report the percentage reduction in time for summary-level evaluation given the sentence-level running time.
-6.4 Evaluation Time
-Figure 4 shows the evaluation times for PlainQAFact across both sentence-level and summary-level settings using PlainFact.
-The percentage scores indicate the reduction in evaluation time when transitioning from sentence-level to summary-level analysis.
-Our findings reveal that summary-level evaluation consistently reduces evaluation time across all cases.
-Moreover, employing PlainQAFact without MedRAG-based retrieval results in a 33.3% reduction in evaluation time for sentence-level analysis.
-Although integrating GPT-4o as both the classifier and answer extractor slightly improves the factuality evaluation performance (shown in Table 2), it also increases the evaluation time, particularly when GPT-4o is used as the classifier.
+Final Answer: long-term effects
+Table 5: Error analysis of PlainQAFact with intermediate metric outputs.
+“Retrieved Knowledge” refers to the source texts retrieved for each plain language sentence.
+We present correct and failure cases sampled only from elaborative explanation examples (i.e., sentences classified as explanations during evaluation).
+The correct case has a PlainQAFact score above 0.6, while the failure cases have scores below 0.5.
+Color legend: extracted answer, answer origin, correct answer, incorrect answer.
 5.4 Error Analysis
-To analyze cases where PlainQAFact fails on the PlainFact benchmark, we categorize three types of error in Table 4).
-For each example, we report the original PLS sentences, the corresponding scientific abstract with retrieved knowledge (from medical Textbooks and StatPearls), model-generated questions based on the extracted answers, and the QA model’s responses.
-In the first case, the extracted answer is “pregnancy,” but the QA model returns “cycle control” from the abstract due to the absence of this term in the retrieved content.
-Since this sentence is labeled as “explanation” which requires additional knowledge for verification, the error suggests potential misclassification between simplification and explanation.
-The second example shows how noisy retrieved information can affect evaluation accuracy, as the QA model’s answer is irrelevant to the extracted answer.
-In the final case, an unanswerable question generated by the QG model (e.g., regarding “Government”) reveals limitations in QA-based evaluation, as the question remains unanswered even with retrieval.
-Additionally, we acknowledge that some PLS summaries or sentences may not generate any questions, resulting in empty question sets.
-Addressing these challenges requires further improvements in classification, retrieval, and question generation.
+To analyze cases where PlainQAFact fails on the PlainFact benchmark, we categorize one correct example and three types of errors in Table 5.
+For each case, we present the original plain language sentence, the corresponding scientific abstract with retrieved knowledge (from medical textbooks and StatPearls), the model-generated questions based on extracted answers, and the QA model’s final responses.
+Each question is generated based on the extracted answer and its corresponding plain language sentence.
+In the correct case, the QG model generates a question from the plain language sentence and the extracted answer “SSB.” The QA model then provides the correct answer using the retrieved knowledge rather than the original abstract, demonstrating PlainQAFact ’s effectiveness in evaluating explanation sentences.
+In the first failure case, the extracted answer is “pregnancy,” but the QA model returns “cycle control” from the abstract because the correct term is missing in the retrieved content.
+Since this sentence is classified as an “elaborative explanation,” which requires external knowledge for verification, this error points to a potential misclassification between simplification and explanation by the fine-tuned classifier.
+The second failure illustrates how noisy retrieved knowledge can impair evaluation.
+The QA model provides an irrelevant answer from the retrieved content that does not match the extracted gold answer.
+The third case shows an unanswerable question generated by the QG model (e.g., regarding “Government”), highlighting a limitation of QA-based factual consistency evaluation.
+Some questions remain unanswered even with retrieval.
+Additionally, we acknowledge that some plain language sentences may not generate any questions, resulting in empty question sets.
+Addressing these challenges requires further improvements in classification, retrieval, and domain-specific question generation.
 6 Discussion and Conclusion
-Automated PLS generation has great potential for improving health information accessibility, but factuality evaluation remains challenging due to elaborative explanations.
-To address this, we introduce PlainQAFact, the first PLS factuality evaluation framework.
-PlainQAFact classifies sentence types, retrieves domain-specific knowledge, and utilizes a QA-based approach to assess the factual consistency of both source-simplified and elaborative sentences.
-By leveraging open-source models, PlainQAFact ensures both accessibility and computational efficiency.
-To support granular factuality evaluation, we construct PlainFact, the first expert-annotated dataset for PLS, covering factuality type, functional role, and sentence alignment.
-Using PlainFact, our experiments demonstrate that PlainQAFact outperforms existing factuality metrics, particularly in handling external elaborative content.
-However, applying sentence-level evaluation to large datasets can be computationally expensive.
-For extensive data, we recommend using summary-level evaluation as a more efficient alternative while maintaining comparable performance.
-Although our study focuses on biomedical PLS generation, the retrieval framework is adaptable to other domains by substituting domain-specific knowledge bases.
-Our findings highlight the effectiveness of retrieval-augmented QA for factuality evaluation, suggesting its broader applicability beyond PLS tasks.
-Future research should explore its integration into diverse summarization applications, particularly in LLM-generated text, where source coverage is often incomplete.
-By introducing a fine-grained and explainable evaluation framework, our work establishes a foundation for advancing QA-based factuality metrics.
-We hope PlainQAFact will contribute to the development of more reliable and interpretable factuality assessment methods in NLP.
-Limitations
-We acknowledge that using an LLM as an answer extractor may yield non-deterministic outputs, even when setting the temperature to 0.01 for the Llama 3.1 8B Instruct model.
-For example, in our retrieval effectiveness evaluation (Table 3), we run PlainQAFact four times to examine the impact of LLM-based answer extraction on the final factuality prediction.
-The resulting scores were 50.60, 50.64, 50.57, and 50.63, with a standard deviation of 0.027.
-We consider this variability to be minimal, demonstrating that our factuality evaluation metric shows robust performance in the plain language summarization evaluation task.
-Our study’s major limitations are as follows: 1) Although we provide substitute solutions for classifying input summaries or sentences, our pre-trained classifier is limited to the biomedical domain and specifically designed for plain language summarization tasks, fine-tuned through PlainFact.
+Our study advances the assessment of hallucinations in the field of PLS by supplying both the first domain expert-annotated biomedical PLS benchmark PlainFact and a novel retrieval-augmented QA-based factual consistency evaluation metric PlainQAFact.
+Unlike existing biomedical domain corpora that either lack fine-grained labels for elaborative content [54] or focus solely on text simplification (e.g., FactPICO[9]), PlainFact captures sentence-level distinctions, including simplification versus explanation, functional roles, and explicit alignment to source sentences.
+This granularity not only enables precise error analysis of hallucinations introduced for clarity, but also provides a reusable framework for researchers in other disciplines (e.g., legal, technical) to replicate our annotation protocol and build high-quality factual consistency datasets tailored to their specialized texts.
+By offering these data and detailed annotation protocol (A), we anticipate community-driven extensions, such as multilingual adaptations or integration with domain-specific ontologies that will democratize development of trustworthy summarization models.
+Building on this resource, PlainQAFact differentiates from prior QA-based metrics (e.g., QAFactEval[21], QuestEval[50]) by first classifying whether a sentence requires external verification given elaborative explanations are often ignored by metrics relying solely on source text, and improve the evaluation efficiency without retrieving for every plain language summary instance.
+Then, retrieving targeted domain knowledge for each elaborative explanation before posing and answering fact-checking questions.
+The ablations demonstrate that selective retrieval yields substantial gains in Kendall’s $\tau$, Pearson correlation, and AUC-ROC over both end-to-end LLM judges[41] and alignment-based scorers[19], highlighting that judicious incorporation of external evidence is essential for robust factual consistency assessment.
+Beyond its empirical strengths, PlainQAFact serves as a plug-and-play evaluation metric for future summarization systems.
+Developers can leverage our classifier to classify sentences, apply retrieval only where necessary, thus containing computational costs and obtain interpretable question–answer pairs that manifest non-factual hallucinations.
+More importantly, PlainQAFact is developed with the open-source backbone model (i.e., Llama 3.1), ensuring broad use with transparency, reproducibility, and fewer computational costs.
+Looking forward, extending our proposed evaluation method to interactive human–AI workflows, where clinicians or subject-matter experts verify QA outputs on-the-fly, could further enhance trust and adoption in high-stakes domains, including medicine, healthcare, or law.
+Despite these promising results, we consider the following limitations of this study.
+(1) Our error analysis reveals that misclassifications, particularly between source simplifications and elaborative explanations, can lead to retrieval mismatches and unanswerable QA prompts.
+These issues underscore the need for more nuanced classification and question generation modules.
+Although we provide substitute solutions for classifying input summaries or sentences, our fine-tuned classifier is limited to the biomedical domain and specifically designed for PLS tasks, fine-tuned through PlainFact.
 We acknowledge that the dataset used for model fine-tuning is insufficient, resulting in only moderate classification accuracy.
-Given the scarcity of high-quality human-annotated data for classifying source simplification and elaborative explanation information, future efforts may focus on developing domain- and task-specific classifiers You et al.
-(2024a) for factuality evaluation.
-2) In the knowledge retrieval phase, we currently incorporate only Textbooks and StatPearls as external resources.
-This limited selection may not be sufficient for evaluating the factuality of plain language summaries.
-Further investigation into additional relevant knowledge bases, such as Wikipedia and PubMed, could enhance performance in explanation-needed factuality evaluation tasks.
-3) For retrieval-augmented QA, we directly concatenate the scientific abstract with the retrieved information to form the reference for QA.
-However, we do not apply knowledge filtering to remove irrelevant information from external resources, which may introduce noise into the subsequent QA process for assessing answer-level factual correctness.
+Given the scarcity of high-quality human-annotated data for classifying source simplification and elaborative explanation information, future efforts are needed on developing domain-specific (e.g., law, healthcare, social science, etc.) classifiers [55] for factual consistency evaluation.
+(2) While our experiments indicate that our retrieval-augmented evaluation metric can improve factual consistency assessment in most of the settings, especially in elaborative explanation evaluation, the computational time increases compared to NLI-based evaluation metrics and LLM-based evaluators.
+The trade-off between evaluation precision and efficiency suggests that further optimization will be beneficial.
+We suggest to balance wisely based on the trade-offs we report in this study regarding the evaluation time and accuracy.
 Acknowledgment
@@
 Our annotation procedure involves two stages, starting with thorough training using detailed examples to ensure consistent understanding of the task between annotators.
-Annotators receive a spreadsheet where each row contains a pair of data: a sentence extracted from the PLS and its corresponding scientific abstract.
+Annotators receive a spreadsheet where each row contains a pair of data: a sentence extracted from the plain language summary and its corresponding scientific abstract.
 For every sentence-abstract pair, the annotators are required to label three features: External, Category, and Relation, with the appropriate labels.
@@
 4. Original_Abstract: The abstract corresponding to each summary, with each sentence indexed for easy reference.
-Each annotator will annotate 40 PLS-abstract pairs to ensure each sentence of the PLS has two sets of annotations from different people.
+Each annotator will annotate 40 summary-abstract pairs to ensure each sentence of the plain language summary has two sets of annotations from different people.
 For each row of the spreadsheet, they need to annotate three columns: “External,” “Category,” and “Relation.”
-$<$TO BE ANNOTATED$>$ External: Mark “Yes” or “No” to indicate if the sentence in the ‘Target_Sentence’ column contains external information.
-$<$TO BE ANNOTATED$>$ Category: Choose the most fitting category of the sentence from the list (Definition, Background, Example, Method/Result, Other).
-$<$TO BE ANNOTATED$>$ Relation: List the relevant sentence indices from the abstract in the ‘Original_Abstract’ column that relate to the plain language sentence (e.g., s10_1,s10_5).
+<<TO BE ANNOTATED>> External: Mark “Yes” or “No” to indicate if the sentence in the ‘Target_Sentence’ column contains external information.
+<<TO BE ANNOTATED>> Category: Choose the most fitting category of the sentence from the list (Definition, Background, Example, Method/Result, Other).
+<<TO BE ANNOTATED>> Relation: List the relevant sentence indices from the abstract in the ‘Original_Abstract’ column that relate to the plain language sentence (e.g., s10_1,s10_5).
 For example, filling in ‘s10_1,s10_5,s10_9’ if you think these three sentences from the abstract are relevant to sentence t1_3.
@@
 Relation s10_1,s10_2
-Corresponding Abstract $<$s10_1$>$Acute cardiovascular events represent a therapeutic challenge.
-$<$s10_2$>$Blood pressure lowering drugs are commonly used and recommended in the early phase of these settings.
+Corresponding Abstract <<s10_1>>Acute cardiovascular events represent a therapeutic challenge.
+<<s10_2>>Blood pressure lowering drugs are commonly used and recommended in the early phase of these settings.
 Example 3 Plain Language Sentence We looked at whether choice of antibiotic made a difference in the number of people who experienced failed treatment, and we determined the proportions who had resolution of fever at 48 hours.
@@
 Relation s15_16,s15_17,s15_20
-Corresponding Abstract $<$s15_16$>$For treatment failure, the difference between doxycycline and tetracycline is uncertain (very low-certainty evidence).
-$<$s15_17$>$Doxycycline compared to tetracycline may make little or no difference in resolution of fever within 48 hours (risk ratio (RR) 1.14, 95% confidence interval (CI) 0.90 to 1.44, 55 participants; one trial; low-certainty evidence) and in time to defervescence (116 participants; one trial; low-certainty evidence).
-$<$s15_20$>$For most outcomes, including treatment failure, resolution of fever within 48 hours, time to defervescence, and serious adverse events, we are uncertain whether study results show a difference between doxycycline and macrolides (very low-certainty evidence).
+Corresponding Abstract <<s15_16>>For treatment failure, the difference between doxycycline and tetracycline is uncertain (very low-certainty evidence).
+<<s15_17>>Doxycycline compared to tetracycline may make little or no difference in resolution of fever within 48 hours (risk ratio (RR) 1.14, 95% confidence interval (CI) 0.90 to 1.44, 55 participants; one trial; low-certainty evidence) and in time to defervescence (116 participants; one trial; low-certainty evidence).
+<<s15_20>>For most outcomes, including treatment failure, resolution of fever within 48 hours, time to defervescence, and serious adverse events, we are uncertain whether study results show a difference between doxycycline and macrolides (very low-certainty evidence).
 Table 6: Examples of our curated dataset.
 Need External Information feature represents whether a plain language sentence is a simplification or an explanation.
-A label of “yes” indicates that the sentence is an explanation and requires additional elaborative information beyond the source abstract to verify its factuality.
+A label of “yes” indicates that the sentence is an explanation and requires additional elaborative information beyond the source abstract to verify its factual consistency.
 Conversely, a label of “no” shows that the sentence is a simplification that can be validated using only the source abstract.
@@
 A “Yes” label indicates that the sentence includes information not explicitly mentioned in the abstract, and vice versa.
-The “Relation” feature identifies the sentence(s) in the scientific abstract most relevant to the PLS; if no corresponding content exists in the abstract, the relation is marked as “external.” We randomly sample three examples from the dataset to to illustrate the dataset’s structure.
+The “Relation” feature identifies the sentence(s) in the scientific abstract most relevant to the plain language summaries; if no corresponding content exists in the abstract, the relation is marked as “external.” We randomly sample three examples from the dataset to to illustrate the dataset’s structure.
 Additionally, indexes have been created for both the plain language sentences and the abstract sentences to facilitate annotation.
 Appendix C LLM Prompts
-We utilize two types of prompts to guide LLMs through two stages of factuality evaluation: classification and answer extraction.
-In accordance with the benchmark annotation protocol (Appendix A), we employ GPT-4o as a classifier to determine whether a given sentence or summary requires elaborative explanations for factual verification.
+We utilize two types of prompts to guide LLMs through two stages of FC evaluation: classification and answer extraction.
+In accordance with the benchmark annotation protocol (A), we employ GPT-4o as a classifier to determine whether a given sentence or summary requires elaborative explanations for factual verification.
 For both stages, we set the max_tokens parameter to 512 and configure the temperature to 0 for GPT-4o and 0.01 for the Llama 3.1 8B Instruc model.
@@
 Example 1: Prompt of GPT-4o as the Classifier.
-For the AE stage, we explore both GPT-4o and Llama 3.1 8B Instruct as backbone models to assess the factuality evaluation performance of PlainQAFact.
-Following the task description outlined in QAFactEval Fabbri et al.
-(2022a), we instruct both LLMs to extract potential answer entities from the input PLS.
+For the AE stage, we explore both GPT-4o and Llama 3.1 8B Instruct as backbone models to assess the FC evaluation performance of PlainQAFact.
+Following the task description outlined in QAFactEval [39], we instruct both LLMs to extract potential answer entities from the input PLS.
 Developer
@@
 Example 3: Prompt of Llama 3.1 8B Instruct as the Answer Extractor.
-Additionally, we report the performance of using Llama 3.1 8B Instruct as a judge to evaluate the factuality of a given PLS based on its source scientific abstract.
+Additionally, we report the performance of using Llama 3.1 8B Instruct as a judge to evaluate the FC of a given PLS based on its source scientific abstract.
 System
@@
 Factuality score (only output a numeric score): <score>
-Example 4: Prompt of Llama 3.1 8B Instruct as a factuality judge.
+Example 4: Prompt of Llama 3.1 8B Instruct as a FC judge.
+For the LLM-based perturbation of PlainFact and CELLS datasets, we follow the protocol of APPLS [12] on faithfulness criteria.
+System
+You are a data transformation assistant.
+You will receive a sentence from a biomedical literature.
+You will generate a new version of the given sentence based on the following rules for faithfulness perturbations:
+1. Number Swap
+Locate any numeric value(s) in the sentence and swap them with different numeric value(s).
+Example: "infected more than 59 million people" -> "infected more than 64 million people"
+2. Entity Swap
+Locate a key entity (e.g., virus name, drug name, organization) in the sentence and swap it with a different entity.
+Example: "coronavirus 2 (SARS-CoV-2)" -> "canine adenovirus (CAV-2)"
+3. Synonym Verb Swap
+Identify a key verb in the sentence and replace it with a near-synonym or related verb that changes the nuance or meaning slightly.
+Example: "killed more than one of them" -> "stamped out more than one of them"
+4. Hypernym/Antonym Swap
+Select a word and replace it with either a hypernym (a more general term) or an antonym (opposite meaning), as appropriate.
+Example (antonym): "killed more than one of them" -> "saved more than one of them"
+Example (hypernym): "dog" -> "animal" (if relevant)
+5. Negation
+Negate a key part of the sentence to flip its meaning.
+Example: "has infected more than 59 million people" -> "hasnt infected more than 59 million people" or "has not infected more than 59 million people"
+Your task:
+Read each sentence, generate a new sentence based on one of the five types of perturbation stratigies (Number Swap, Entity Swap, Synonym Verb Swap, Hypernym/Antonym Swap, Negation) above.
+Return the perturbation sentence.
+Do not change the rest parts of the sentence except for the perturbation content.
+For example, the original sentence is "The skin patch and the vaginal (birth canal) ring are two methods of birth control." The perturbation sentence should be "The skin patch and the vaginal (birth canal) ring are five methods of birth control."
+User
+Sentence: <input>
+Perturbation sentence: <sentence>
+Example 5: Prompt of GPT-4o for faithfulness perturbation
 Appendix D Detailed Experiment Settings
 All experiments we conduct are under one NVIDIA A100 GPU with 40 GB GPU memory.
-We employ the Natural Language Toolkit (NLTK) [^6] https://www.nltk.org/ for PLS tokenization and sentence segmentation, utilizing the punkt package.
-For classifier fine-tuning, we set the random seed to 42 to split the dataset into training, validation, and test sets.
+We employ the Natural Language Toolkit (NLTK) [^5] https://www.nltk.org/ to split the plain language summaries into sentences using the punkt package.
+For the classifier fine-tuning, we set the random seed to 42 to split PlainFact into training, validation, and test sets.
 We tune the PubMedBERT-base model for three epochs with early stopping under the validation set.
-The same seed (42) is also used to sample 200 PLS-abstract pairs from the CELLS test set for the comparison in Figure 2.
-The QA model used in our PlainQAFact is downloaded from QAFactEval Fabbri et al.
-(2022a); however, we set its maximum input length to 512 (from 364) tokens to incorporate more source context.
+The final fine-tuning accuracy of the classifier on the test set is 0.77.
+The same seed (42) is also used to sample 200 summary-abstract pairs from the CELLS test set for the comparison in Table 2.
+The QA model used in our PlainQAFact is downloaded from QAFactEval [39].
+However, we set its maximum input length to 512 (from 364) tokens to incorporate more source context.
 Similarly, to ensure that the answers extracted during the AE stage are valid (i.e., within the 512-token limit to maintain consistency with the subsequent QA model), we configure the LLMs’ input lengths to 512 tokens.
-Component Method Choice Benchmark
-PlainQAFact 55.8 (0.12)
-Classifier Learned classifier -
-GPT-4o 59.1 (0.05)
-No (retrieve for all) 60.1* (0.18)
-Answer Extraction Llama3.1 8B Instruct -
-GPT-4o 54.8 (0.15)
-PromptRank 48.4*
-NP Chunks 50.5*
-Retrieval Source Abs + TB + SP -
-TB + SP 47.9* (0.11)
-Abs (no retrieval) 55.2 (0.07)
-Answer Overlap BERTScore -
-LERC 32.9* (0.18)
-Table 6: Ablation study of PlainQAFact on PlainFact in the summary level, analyzing the impact of individual components in PlainQAFact.
-Abs: abstract; TB: Textbooks; SP: StatPearls.
-Using the first row as the baseline, * indicates a statistically significant difference based on a paired t-test ($p<0.05$).
-We run each setting for five times and report the standard deviation (std.).
-Note that only the settings using LLMs may produce fluctuated scores.
-Appendix E Summary-Level Ablation Study of PlainQAFact
-Similar to the sentence-level ablation studies in Table 2, we conduct experiments to evaluate the impact of various components in PlainQAFact at the summary level using PLSs as inputs.
-Although the overall evaluation score using GPT-4o as the classifier is 3.3% higher than that of PlainQAFact, the paired t-test (p<0.05) does not provide sufficient statistical evidence to conclude that GPT-4o significantly outperforms our learned classifier.
-This finding indicates that GPT-4o’s performance is comparable with our learned classifier, consistent with the sentence-level results.
-Notably, in the summary-level evaluation, using only scientific abstracts as references for QA-based evaluation yields performance that is competitive with using retrieved external knowledge.
-We assume that summary-level retrieval may not consistently produce highly relevant knowledge snippets and might even introduce noise that undermines factuality evaluation accuracy.
-Additionally, omitting the classification stage and retrieving information for all summaries produces the best performance among all settings.
-Although this approach increases the runtime to approximately 30 minutes compared to using the learned classifier for 200 summaries, the performance improvement is significant.
-Therefore, we recommend either employing more targeted knowledge bases for summary-level retrieval, compressing abstracts to include more relevant information, or retrieving for all input summaries when conducting summary-level factuality evaluation with PlainQAFact.
-Figure 5: Evaluation performance across metrics for explanation sentences under PlainFact.
-We select only those summaries containing at least five explanation sentences, resulting in a total of 116 summaries.
-Note that this experiment is conducted at the summary level.
-A higher change percentage indicates a metric is more sensitive to the explanation information, implying that such metrics may not effectively verify the factuality of PLS.
-Ideally, we expect the score change percentage to increase slowly.
-Among the metrics evaluated, only AlignScore, SummaC, and our method show this expected pattern.
-Appendix F Explanation Sentence Removal
-To further assess the performance of existing factuality metrics on elaborative explanation sentences in PLSs, we conduct experiments in which these explanation sentences were iteratively removed from the summaries.
-Figure 5 illustrates the percentage change in scores relative to the baselines as more explanation sentences are removed.
-Notably, PlainQAFact can even predict 100% factual ratings for the given PLS when more explanation information is removed, whereas other metrics struggle to do so.
-Although our framework shows moderate performance in Figure 5, with a gradual increase in score as explanation sentences are removed—we believe that our retrieval-augmented, QA-based factuality evaluation framework more effectively assesses the factuality of plain language sentences that require elaborative explanations for verification.
+In MedCPT retrieval, we set k (the number of retrieved snippets) to 3, ensuring the retrieved information remains within a short context window for the subsequent QA process (D).
+We apply the default setting (i.e., RoBERTa-base) for AlignScore [19] and SummaC-Conv for SummaC [16].
+For GPT-4o experiments as an factual consistency evaluator, we set the temperature as 1.0 in this paper.
+Appendix E Statistical Testing
+Following SummaC [16] and QAFactEval [39], we test whether our proposed metric achieves statistically significant improvements over other methods.
+Given Kendall’s $\tau$ and Pearson correlations do not show the discriminative feature of continuous scores in various evaluation metric, we run a systematic evaluation on four datasets (CELLS, FareBio, FactPICO, and PlainFact) through bootstrap resampling [56] on the AUC-ROC results.
+We compare our best metric to other methods using the confidence intervals as significance level of 0.05 and apply Bonferroni correction [57] following SummaC [16].
+Our results show that PlainQAFact achieves statistically significant improvements over QuestEval [50], SummaC [16], and DAE [48] on the CELLS and PlainFact datasets.
+As shown in Figure 2, PlainQAFact also outperforms other automatic metrics in evaluating elaborative explanation sentences, except for QAFactEval [39].
+This highlights the effectiveness of PlainQAFact for biomedical PLS tasks that include external explanations.
+On the other hand, when evaluating elaborative content in the FactPICO and FareBIO datasets (Figures 3 and 4), we do not find statistically significant differences across metrics.
+While some existing metrics perform similarly in certain settings, our metric consistently shows better performance across all three evaluation criteria.
+Figure 3: Overall performance on summaries containing added information (i.e., elaborative explanations) from FactPICO [9].
+The std.
+of PlainQAFact, Llama 3.1, and GPT-4o are 0.2, 0.2, and 3.3, respectively, based on five runs of each metric (details in  E).
+Figure 4: Overall performance on summaries with elaborative explanations from FareBio [47].
+The std.
+for PlainQAFact, Llama 3.1, and GPT-4o are 0.03, 0.3, and 5.7, respectively, computed over five runs (details in  E.)
+Appendix F Summary-Level Explanation-Only Evaluation
+Similar to our sentence-level explanation annotations in PlainFact, FactPICO [9] and FareBio [47] datasets also provide human-annotated explanation information in plain language summaries.
+According to the results shown in Section 5.2, we also assess the performance of five factual consistency evaluation metrics under FactPICO and FareBio datasets.
+In the FactPICO, the elaborative explanation is defined as “added information.” We first select summaries that include added information, and then we label a summary as “non-factual” if it contains any non-factual added information as determined by annotators.
+We generate a summary-level explanation-only FactPICO dataset consisting of 231 summary-abstract pairs.
+As shown in Figure 3, AlignScore achieves the best performance across all three evaluation criteria.
+However, since the summaries in the FactPICO dataset are generated using LLMs, the factual consistency between the generated summaries and the original source abstracts is not guaranteed.
+For instance, we treat a summary as “factual” indicates that all the added information is factual, but it does not necessarily reflect the factual consistency of other content.
+Therefore, the results presented in Figure 3 may become biased.
+For FareBio [47], we collect explanation sentences based on both “faithfulness” and “factual hallucination” labels in the original dataset.
+As shown in Figure 4, PlainQAFact achieves the best performance compared with all other metrics, showing a clear advantage over GPT-4o.
+Since the FareBio dataset provides sentence-level annotations of factual consistency for each sentence in the plain language summaries, the results are more reliable than those from FactPICO.
 Appendix G Pilot Study on FactPICO
-In PLS generation tasks, we employ the recently introduced FactPICO dataset Joseph et al.
-(2024).
+To investigate the performance of existing automatic factual consistency evaluation metrics on plain language generation tasks, we employ the recently introduced FactPICO dataset [9].
 This dataset comprises human-labeled plain language summaries of Randomized Controlled Trials (RCTs) that address several critical elements: Populations, Interventions, Comparators, Outcomes (PICO), and additional information.
-All summaries are generated by various LLMs based on scientific literature and include extensive explanations not present in the original abstracts.
+All summaries are generated by various LLMs based on medical literature and include added information (i.e., extensive explanations) not present in the original abstracts.
 We hypothesize that existing factuality evaluation metrics for text summarization may struggle to accurately assess the factuality of this added information.
-To validate this assumption, we conduct pilot studies using four factuality evaluation metrics from FactPICO alongside one NLI-based metric: DAE Goyal and Durrett (2021), AlignScore Zha et al.
-(2023), SummaC Laban et al.
-(2022), QAFactEval Fabbri et al.
-(2022a), and QuestEval Scialom et al.
-(2021b).
+To validate this assumption, we conduct pilot studies using four factual consistency evaluation metrics from FactPICO alongside one NLI-based metric: DAE [48], AlignScore [19], SummaC [16], QAFactEval [39], and QuestEval [50].
 G.1 Dataset Pre-processing
@@
 G.2 Experiments and Results
-We conduct experiments using the processed FactPICO dataset to address both research questions.
-It is important to note that there are some outliers in the PLSs which contain extraneous content and exceed the input length limitations of the DAE and QAFactEval metrics.
-Therefore, we are unable to include these two metrics in this pilot study.
-Figure 6: Score change percentage from baselines over three metrics on the FactPICO dataset in removing factual added information (88 pairs).
-We expect each model get higher performance when more external information is removed.
+We conduct experiments using the processed FactPICO dataset to conduct pilot studies that assess the performance of existing factual consistency evaluation metrics in detecting added information in plain language summarization tasks.
+It is important to note that some outliers in the plain language summaries of the FactPICO dataset contain extraneous content, which causes the input lengths to exceed the limitations of both the DAE and QAFactEval metrics.
+Furthermore, the factuality of the LLM-generated plain language summaries is not guaranteed.
+Figure 5: Score change percentage from baselines over five metrics on the FactPICO dataset in removing factual added information.
+We expect each metric stays unchanged even when more added factual information is removed.
 The evaluation dataset contains 88 valid summary-abstract pairs.
-RQ1: Do existing metrics perform well when external information is added to PLS compared with no added information?
-In this study, we evaluate the ability of existing metrics to detect added information in PLSs.
-We focus on PLSs where all added information is annotated as factual, resulting in a deduplicated dataset of 88 PLS-abstract pairs.
-To assess metric sensitivity to the added information, we iteratively remove sentences from each PLS that contain added spans.
-For example, if a span such as “of a medicine called haloperidol” is labeled as factual (“yes”), we remove the entire sentence in the original PLS containing that span through exact matching, continuing this process until no added information remains.
-This procedure allows us to determine how effectively existing metrics handle external information that is not present in the original abstracts.
+Figure 6: Score change percentage from baselines over five metrics on the FactPICO dataset in removing non-factual added information (60 pairs).
+We expect the change percentage from baseline increases when more added non-factual information is removed.
+RQ1: Do existing metrics perform well when external information is added to plain language summaries compared with no added information?
+In this study, we evaluate the ability of existing factual consistency evaluation metrics to detect added information in plain language summaries.
+We focus on summaries where all added information is annotated as factual, resulting in a dataset of 88 summary-abstract pairs.
+To assess metric sensitivity to the added information, we iteratively remove sentences from each plain language summary that contain added spans.
+For example, if a span such as “of a medicine called haloperidol” is labeled as factual (“yes”), we remove the entire sentence in the original plain language summaries containing that span through exact matching, continuing this process until no added information remains.
+This procedure enables us to determine how effectively current metrics handle added information that is absent from the original abstracts.
 As shown in Figure 5, we report performance changes relative to baseline scores.
-For example, AlignScore increases by 3.33 (on a scale of 100) when six units of added information are removed, leading to a change of approximately 4.9%.
-These results indicate that added information has impacts on model-based factuality evaluation metrics such as AlignScore and SummaC, with both scores improving as more added information is removed.
-In contrast, QuestEval’s score decreases with the removal of added information.
-Figure 7: Score change percentage from baselines over three metrics on the FactPICO dataset in removing non-factual added information (60 pairs).
-We expect each metric’s score increases when more non-factual information is removed.
+For example, the AlignScore evaluation increases by 3.3% (on a 100-point scale) when six spans of added information are removed, resulting in an overall change of approximately 4.9% compared to the setting in which no added information is removed.
+These findings indicate that added information affects model-based factuality evaluation metrics such as QAFactEval (6.0%), AlignScore (4.9%), and SummaC (1.9%), with scores improving as more added information is removed.
+In contrast, both QuestEval and DAE scores decline with the removal of added information, and notably, DAE exhibits the most significant performance drop.
+Overall, these findings suggest that all the five metrics have difficulty accurately assessing added factual information, as evidenced by both increases and decreases in their performance.
 RQ2: Can existing metrics distinguish between non-factual and factual added information?
-In this research question, our goal is to evaluate the sensitivity of factuality evaluation metrics in detecting non-factual information within PLSs.
+In this research question, our goal is to evaluate the sensitivity of factuality evaluation metrics in detecting non-factual information within plain language summaries.
 The FactPICO dataset labels added information as either “yes” (factual) or “no” (non-factual).
-For this scenario, we sample only those PLSs that contain both “yes” and “no” labels for added spans.
-As with RQ1, we iteratively remove sentences containing non-factual added spans from each PLS until no non-factual sentences remain.
-Overall, as illustrated in Figure 7, most evaluation metrics show improved performance as more non-factual information is removed.
-Nevertheless, our findings suggest that existing factuality evaluation metrics have limited capacity to accurately distinguish between factual and non-factual added information in text summarization tasks.
+In this scenario, we sample only those plain language summaries that contain both “yes” and “no” labels for added spans.
+As with RQ1, we iteratively remove sentences containing non-factual added spans from each plain language summary until no non-factual sentences remain.
+Overall, as illustrated in Figure 6, only AlignScore, SummaC, and QAFactEval show improved performance as more non-factual information is removed, indicating that these metrics are sensitive in added non-factual information.
+Nevertheless, based on the results of RQ 1 and 2, our findings suggest that existing factual consistency evaluation metrics have limited capacity to accurately distinguish between factual and non-factual added information in plain language summarization tasks.

### v3

@@
 We further analyze PlainQAFact’s effectiveness across external knowledge sources, answer extraction strategies, answer overlap measures, and document granularity levels, refining its overall factual consistency assessment.
-Taken together, our work presents the first evaluation metric designed for PLS factual consistency evaluation, providing the community with both a robust benchmark and a practical tool to advance reliable and safe plain language communication in the medical domain.
-PlainQAFact and PlainFact are available at: https://github.com/zhiwenyou103/PlainQAFact.
+Taken together, our work presents a sentence-aware, retrieval-augmented metric targeted at elaborative explanations in biomedical PLS tasks, providing the community with both a robust benchmark and a practical tool to advance reliable and safe plain language communication in the medical domain.
+PlainQAFact and PlainFact are available at: [Uncaptioned image] https://github.com/zhiwenyou103/PlainQAFact.
 keywords:
@@
 source simplification), functional role, and alignment to the source (§3.1).
-Building on this, we present PlainQAFact, a dual-stage QA-based evaluation metric that selectively applies retrieval only to elaborative explanations (§3.2).
-This design ensures both efficiency and fidelity in evaluating factual consistency.
-Experiment results on several PLS datasets demonstrate the effectiveness of our approach, particularly in assessing complex, elaborative content (4, §5).
-Taken together, our work presents the first evaluation metric designed for PLS factual consistency evaluation, providing the community with both a robust benchmark and a practical tool to advance reliable and safe plain language communication in the medical domain.
+Building on this and prior studies for the QA-based evaluation metrics in the general domain [20, 21], we present PlainQAFact, a dual-stage QA-based evaluation metric that selectively applies retrieval only to elaborative explanations (§3.2).
+Experiment results on several PLS datasets demonstrate the effectiveness of our approach, particularly in assessing complex, elaborative content (§4, §5).
+Taken together, our work presents the first metric specifically designed to evaluate the factual consistency in elaborative explanations in PLS, providing the community with both a robust benchmark and a practical tool to advance reliable and safe plain language communication in the medical domain.
 2 Related Work
@@
 Details of annotation guidelines are provided in A.
+Note that “functional role” and “sentence alignment” are not components of developing PlainQAFact.
+We use these two fields in PlainFact as metadata to conduct error analysis and document our annotation process in A and  B.
 Annotations are conducted by four independent annotators, each with at least a bachelor’s degree in biomedical sciences and prior fact-checking experience.
@@
 Each summary-abstract pair is annotated by two independent annotators, with disagreements resolved by a third.
-Inter-rater agreement, measured by Cohen’s Kappa for factuality type, functional role, and sentence alignment are 0.43, 0.60, and 0.55 respectively, indicated moderate agreement for all tasks [33].
+Inter-rater agreement (IRA), measured by Cohen’s Kappa for factuality type, functional role, and sentence alignment are 0.43, 0.60, and 0.55 respectively computed from 2,740 annotated plain language sentences, indicated fair agreement for all tasks [33].
 Table 1 summarizes the dataset characteristics.
@@
 For annotation examples, see B.
+Additionally, we provide qualitative analysis of PlainFact in  C given the fair IRA of factuality type.
 3.2 PlainQAFact Framework
@@
 Our PlainFact is split 8:1:1 for training, validation, and testing.
-Additionally, we compare the pre-trained classifier with GPT-4o [34] as a zero-shot classifier (prompts are provided in C) to explore extending our factual consistency evaluation metric to domains and tasks lacking human-annotated data.
+Additionally, we compare the pre-trained classifier with GPT-4o [34] as a zero-shot classifier to explore extending our factual consistency evaluation metric to domains and tasks lacking human-annotated data.
 3.2.2 Domain Knowledge Retrieval
@@
 Since our dataset is in the medical domain, we employ MedCPT [36], a retriever pre-trained on large-scale PubMed search logs to generate biomedical text embeddings.
-For external resources, we incorporate StatPearls [37] for clinical decision support and medical textbooks [38] for domain-specific knowledge.
+For external courpus, we incorporate StatPearls [^4] https://huggingface.co/datasets/MedRAG/statpearls [37] for clinical decision support and medical Textbooks [^5] https://huggingface.co/datasets/MedRAG/textbooks [38] for domain-specific medical knowledge (see details in  D.
 3.2.3 QA Evaluation Components
@@
 In this study, we adopt a QA-based approach as a backbone of our evaluation metric to verify factual consistency, incorporating answer extraction, question generation, question filtering, question answering, and answer overlap calculation modules.
-Gold Answer Extraction
+Gold Answer Extraction (AE)
 The first step in QA-based factual consistency evaluation is to extract answer entities (keyphrases) from plain language summaries as gold answer, then verify factuality by comparing them with answers generated by a QA model for the same questions.
@@
 To compare answer extraction strategies, we use PromptRank as the baseline, and evaluate LLM-based extractors, including an open-source Llama 3.1 8B Instruct model (Llama 3.1) [41] and close-source GPT-4o (GPT-4o) [34].
-Prompts for Llama 3.1 and GPT-4o are in C.
+Prompts for Llama 3.1 and GPT-4o are in E.4.
 Question Generation (QG)
@@
 Specifically, the simplification score $s$ is computed using only abstracts as the source in QA, while the explanation score $e$ is calculated by incorporating both abstracts and retrieved knowledge as source for QA-based evaluation.
+$n$ is the number of sentences in a plain language summary, after we split the summary using Natural Language Toolkit (NLTK) [^7] https://www.nltk.org/ and and classify each sentence with our fine-tuned classifier.
 $\textsc{PlainQAFact}=\frac{s_{\text{Avg.}}\cdot n_{s}+e_{\text{Avg.}}\cdot n_{e}}{n_{s}+n_{e}},$ (1)
@@
 We also compare PlainQAFact on the FareBio dataset, which contains 175 plain language summaries generated by seven different LLMs, together with the source articles [47].
-Specifically, FareBio includes 25 full scientific article examples, each source article is summarized by seven LLMs, resulting 175 plain language summaries in total.
-We sample plain language sentences that are not both faithful and factual as non-factual sentences, and the rest as factual sentences, resulting in 33 and 174 summaries respectively.
-We also extract sentences that are labeled as not faithful but factual in their “factual hallucination” label, which represents the plain language sentence that is not from the source article, but still conveying factual information, resulting in 53 valid summaries in total.
-These sentences are equal to our factual elaborative explanation sentences.
-Note in FareBio dataset, there are no non-factual simplification sentences.
-Therefore, we use the same set of non-factual sentences (33 sentences) for main experiment (Table 2) and explanation-only (Figure 2) evaluation.
+Specifically, FareBio includes 25 full scientific articles, each summarized by seven LLMs, resulting 175 plain language summaries.
+Because FareBio does not include non-factual simplification sentences, we treat all simplification sentences as “factual.” For each original summary, we form new plain language summaries by aggregating all sentences annotated as factual and all sentences annotated as non factual separately.
+This produces 33 non factual and 174 factual summaries.
+We also identify explanation sentences that receive the “factual hallucination” label (i.e., they are not verifiable from the source but are factually correct based on external knowledge) and build 53 factual explanation only summaries from them.
+We use the same set of 33 non factual summaries for both the main experiment (Table 2) and the explanation only evaluation (Figure 6).
 4.2 Experiment Settings
@@
 The maximum input length is set to 512 for all models in PlainQAFact.
-More details of prompts and model hyper-parameters for GPT-4o and Llama 3.1 are provided in C and  D.
+More details of prompts and model hyper-parameters for GPT-4o and Llama 3.1 are provided in E.
 4.3 Existing Factuality Metrics
@@
 (4) QAFactEval [39] is a QA-based factuality evaluation metric that assesses summary consistency by generating questions from the source document and comparing the model-generated answers with expected answers.
-(5) QuestEval [50] is a reference-less evaluation metric employs a T5 model to generate questions from the source document and verifies whether the summary can correctly answer them.
+(5) QuestEval [50] is a reference‐less evaluation metric employs a T5 model to generate questions from the source document and verifies whether the summary can correctly answer them.
 5 Results and Analysis
 We first conduct a pilot study on the FactPICO dataset [9] to investigate the performance of existing factual consistency evaluation metrics on PLS tasks.
-Our study (G) reveals two major limitations.
+Our pilot study (F) reveals two major limitations.
 First, many LLM-generated plain language summaries in FactPICO contain incomplete or incoherent sentences, highlighting the need for a benchmark based on high-quality, human-authored plain language summaries.
-Second, existing automatic factual consistency metrics are not sensitive to evaluate the factuality of “added information”, underscoring the need for a more domain-related factual consistency evaluation metric, especially for those summaries with additional explanations that are not mentioned by the original source texts.
+Second, existing automatic factual consistency metrics are not sensitive to evaluate the factuality of “added information,” underscoring the need for a more domain-related factual consistency evaluation metric, especially for those summaries with additional explanations that are not mentioned by the original source texts.
 To address these challenges, we introduce a new expert-annotated dataset, PlainFact, and propose PlainQAFact, a QA-based, retrieval-augmented metric for evaluating factual consistency in plain language summaries.
@@
 Next, we conduct an error analysis to examine failure cases and highlight directions for further improvement (§5.4).
-Datasets Metrics Kendall’s $\tau$ Pearson AUC-ROC
-CELLS Llama 3.1 57.5 42.1 83.4
-GPT-4o 70.0 75.4 99.3
-QAFactEval 61.6 70.6 93.5
-QuestEval 23.2 28.3 66.4*
-SummaC 24.8 29.8 67.5*
-AlignScore 56.0 66.2 89.6
-DAE 14.4 14.4 55.0*
-PlainQAFact 62.1 75.2 93.8
-FareBio Llama 3.1 40.5 38.9 83.0
-GPT-4o 27.3 35.7 75.7
-QAFactEval 32.8 43.9 81.6
-QuestEval 37.8 45.2 86.4
-SummaC 33.8 32.1 82.5
-AlignScore 32.2 47.2 81.0
-DAE 8.8 8.8 53.4
-PlainQAFact 16.8 35.6 66.1
-PlainFact Llama 3.1 60.7 52.6 85.3
-GPT-4o 69.8 80.3 99.2
-QAFactEval 65.8 79.3 96.4
-QuestEval 28.4 34.7 70.1*
-SummaC 34.1 42.8 74.1*
-AlignScore 60.2 72.6 92.5
-DAE 5.3 5.3 51.8*
-PlainQAFact 65.7 81.3 96.4
+Datasets Metrics Kendall’s $\tau$ Pearson AUC (99% CI)
+CELLS Llama 3.1 57.5 42.1 83.4* [78.4, 88.0]
+GPT-4o 70.0 75.4 99.3 [98.6, 99.8]
+QAFactEval 61.6 70.6 93.5 [90.2, 96.3]
+QuestEval 23.2 28.3 66.4* [59.1, 73.3]
+SummaC 24.8 29.8 67.6* [60.3, 74.3]
+AlignScore 56.0 66.2 89.6 [85.3, 93.3]
+DAE 14.4 14.4 55.0* [50.7, 59.6]
+PlainQAFact 62.1 75.2 93.8 [90.3, 96.7]
+FareBio Llama 3.1 40.5 38.9 83.0 [69.0, 94.0]
+GPT-4o 27.3 35.7 75.7 [59.7, 89.8]
+QAFactEval 32.8 43.9 81.6 [69.2, 91.7]
+QuestEval 37.8 45.2 86.4 [78.0, 93.3]
+SummaC 33.8 32.1 82.5 [69.7, 92.9]
+AlignScore 32.2 47.2 81.0 [65.9, 93.0]
+DAE 8.8 8.8 53.4 [47.6, 57.6]
+PlainQAFact 16.8 35.6 66.1 [48.8, 82.1]
+PlainFact Llama 3.1 60.7 52.6 85.3* [80.6, 89.7]
+GPT-4o 69.8 80.3 99.2 [98.4, 99.8]
+QAFactEval 65.8 79.3 96.4 [94.0, 98.4]
+QuestEval 28.4 34.7 70.1* [63.1, 76.5]
+SummaC 34.1 42.8 74.1* [67.6, 80.4]
+AlignScore 60.2 72.6 92.5 [93.8, 98.4]
+DAE 5.3 5.3 51.7* [47.5, 55.9]
+PlainQAFact 65.7 81.3 96.4 [93.6, 98.5]
 Table 2: Evaluation results of automatic metrics on the CELLS [28], PlainFact, and FareBio [47] datasets.
-Results are evaluated using Kendall’s $\tau$, Pearson correlation, and AUC-ROC, with eight metric scores compared against human-labeled factuality.
+Results are evaluated using Kendall’s $\tau$, Pearson correlation, and AUC-ROC (AUC), with eight metric scores compared against human-labeled factual consistency.
 The standard deviations (std.) over five runs are: 0.2 (PlainQAFact), 0.5 (Llama 3.1), and 6.0 (GPT-4o) on PlainFact; 0.1 (PlainQAFact), 0.2 (Llama 3.1), and 6.3 (GPT-4o) on CELLS; and 0.1 (PlainQAFact), 2.9 (Llama 3.1), and 17.0 (GPT-4o) on FareBio.
-* indicates an improvement of our metric over prior work with 95% confidence interval (details in  E).
+* indicates a statistically significant difference compared to PlainQAFact ($p<0.01$).
 Note the plain language summaries in FareBio are generated by seven LLMs, while CELLS and PlainFact are written by the authors of original articles.
@@
 DAE is consistently the weakest across all datasets and metrics, suggesting that a dependency-only metric is insufficient for capturing the factual consistency required in PLS tasks.
-On CELLS, PlainQAFact outperforms all other automatic metrics on all three criteria (Kendall’s $\tau$=62.1, Pearson=75.2, AUC-ROC=93.8).
-This shows that our approach, first detecting elaborative content and then verifying it with retrieval, works better than existing QA-based metrics, which ask all questions from the source only, and also better than NLI-based evaluation metrics.
-Similarly, on PlainFact, PlainQAFact achieves the highest Pearson (81.3), surpassing GPT-4o by 1.0 point.
+On CELLS, PlainQAFact outperforms all other automatic metrics on all three criteria.
+Specifically, PlainQAFact achieves the best Kendall’s $\tau$ ($62.1$) and Pearson ($75.2$) among all the automatic factual consistency evaluation metrics.
+As a comparative effect size, the pairwise difference in AUC-ROC between PlainQAFact and QAFactEval is small ($\Delta\mathrm{AUC-ROC}=0.3$, where $\Delta$ is PlainQAFact minus QAFactEval).
+After Holm–Bonferroni correction [52] on AUC-ROC, PlainQAFact shows significantly higher AUC-ROC than Llama 3.1, QuestEval, SummaC, DAE, and PromptRank ($p<0.01$), whereas the differences with QAFactEval, AlignScore, and GPT-4o are not statistically significant (adjusted $p\geq 0.91$).
+This shows that our approach, first detecting elaborative content and then verifying it with retrieval, performs better than most of the existing reference-based metrics, which ask all questions from the source only.
+Similarly, on PlainFact, PlainQAFact achieves the highest linear correlation with human ratings among automatic metrics (Pearson= $81.3$, 99% CI [76.3, 85.4]), surpassing GPT-4o.
 It ties AUC-ROC (96.4) score with QAFactEval and has a competitive Kendall’s $\tau$ (65.7).
-These results on both datasets show that PlainQAFact is reliable for evaluating both factual and non-factual plain language summaries.
-On FareBio, no single method performs the best on all criteria.
-QuestEval achieves the best AUC-ROC (86.4), AlignScore has the best Pearson (47.2), while Llama 3.1 reaches the best Kendall’s $\tau$ (40.5).
-However, FareBio differs from the other two datasets in several ways: (1) the original labels in FareBio are defined at the sentence level, and a summary is marked as non-factual if any sentence is non-factual; (2) their summaries are generated from full scientific articles rather than abstracts, making it harder for metrics to locate matching source text; (3) the dataset is imbalanced (33 non-factual vs.
-174 factual); and (4) All plain language summaries are generated by LLMs and no elaborative explanation annotations are provided for each sentence or summary.
-These factors likely increase evaluation difficulty on FareBio and impact the performance of multiple metrics, including PlainQAFact and GPT-4o.
+As a comparative effect size, the pairwise difference in AUC-ROC between PlainQAFact and QAFactEval is negligible ($\Delta\mathrm{AUC-ROC}=-0.1$).
+These results on both datasets show that PlainQAFact is reliable for evaluating the factual consistency of plain language summaries.
+See  G for more details of statistical tests.
+On FareBio, no single method is best across all criteria.
+QuestEval achieves the best Kendall’s $\tau$ ($37.8$) and AUC-ROC ($86.4$), while AlignScore achieves the best Pearson correlation ($47.2$).
+However, FareBio differs from the other two datasets in several important ways: (1) the original sentence labels in FareBio are incomplete.
+Any sentence that can be traced back to the source is labeled as factual, so the dataset does not include sentences labeled as non-factual simplifications.
+This design leads to unreliable final summary-level labels.
+For example, even if a summary is labeled as factual and all of its sentences are labeled as factual, some sentences may still contain factual errors; (2) the processed dataset is highly imbalanced (33 non-factual vs.
+174 factual summaries); (3) FareBio reports a fair Cohen’s Kappa of $0.48$ computed from 34 plain language sentences without adjudication, which indicates notable annotation quality issues in their curated dataset; and (4) all plain language summaries are generated by LLMs, and the plain language sentences are not guaranteed to be correct, since the annotation only checks whether the sentence is a source simplification without post-processed non-factual perturbation.
+These factors likely increase the difficulty of evaluation on FareBio and affect the performance of factual consistency evaluation metrics, including PlainQAFact and GPT-4o.
+We also provide a detailed error analysis of FareBio that explains the main limitations of both PlainQAFact and the dataset (§5.5).
+In addition, because the non-factual plain language summaries of PlainFact and CELLS are generated by GPT-4o3.1.1), they may contain detectable artifacts that are not representative of real PLS errors.
+To further test the robustness of our proposed metric, we conduct additional experiments using the heuristic entity swap approach introduced in Guo et al.
+[12] to create non-factual plain language summaries.
+Specifically, following the same experiment setting as APPLS [12], we replace entities in each plain language sentence using the KBIN method [53], which replaces specific entity spans (such as names of viruses or diseases) with related but different concepts derived from the Unified Medical Language System (UMLS) [^8] https://www.nlm.nih.gov/research/umls/ while maximizing NLI contradiction and minimizing LM perplexity.
+This approach ensures that each new perturbed sentence is not only different from the original sentence, but also logically contradicts it while remaining grammatically fluent.
+Overall, on the CELLS dataset, for both the LLM-perturbed setting and the heuristic entity swap method, our metric achieves the best performance across all three evaluation criteria compared with other automatic metrics, showing the robustness of our proposed metric.
+Detailed experiment results are reported in  H.
 In summary, PlainQAFact is a more consistent and effective metric for factual consistency compared to existing automatic approaches.
@@
 of PlainQAFact, Llama 3.1, and GPT-4o are 0.1, 1.0, and 7.7, respectively based on five runs for each metric.
-* indicates an improvement of our metric over prior work with 95% confidence interval (details in  E).
+* indicates a statistically significant difference compared to PlainQAFact ($p<0.01$).
 PlainQAFact significantly outperforms most of the automatic factual consistency evaluation metrics in AUC-ROC.
 Note that the CELLS dataset does not contain annotations for elaborative explanations.
-Results of explanation-only evaluation on FactPICO and FareBio are reported in F.
+Results of explanation-only evaluation on FactPICO and FareBio are reported in I.
 5.2 Explanation-Only Evaluation
@@
 Note that CELLS does not provide sentence-level annotations for elaborative explanations, so we evaluate it only as a general PLS dataset (§5.1).
-As shown in Figure 2, PlainQAFact outperforms all existing automatic metrics across all three criteria.
-It is 5.7 points higher in Kendall’s $\tau$, 7.5 points higher in Pearson, and 4.0 points higher in AUC-ROC compared with the second-best metric, QAFactEval.
-These results indicate that PlainQAFact captures the factual consistency of elaborative explanations more effectively than prior automatic metrics.
-While GPT-4o achieves higher overall scores (Kendall’s $\tau$=54.9, Pearson=60.7, AUC-ROC=88.7), it shows much higher variance (std.
-7.7) compared to PlainQAFact (std.
-0.1), making it less stable for factual consistency evaluation.
+[…]
 Compared to the results in Table 2, AlignScore is competitive on general PLS tasks but drops on evaluating explanation-only sentences.
@@
 Since 44% of sentences in our curated benchmark are labeled as elaborative by human annotators, we believe that PlainQAFact is the more suitable and robust factual consistency evaluation metric in this scenario.
-PlainQAFact also outperforms Llama 3.1 on all three criteria (27.9/28.3/67.0), showing that instruction-tuned LLM judges alone are not enough for reliable evaluation of elaborative content.
-Component Method Choice Kendall’s $\tau$ Pearson AUC std.
-PlainQAFact 65.7 81.3 96.4 0.2
+Component Method Choice $\tau$ Pearson AUC (99% CI) std.
+PlainQAFact 65.7 81.3 96.4 [93.6, 98.5] 0.2
 Classifier Fine-tuned classifier - - - -
-GPT-4o 62.8 77.1 94.3* 0.1
-No (retrieve for all) 61.2 74.2 93.2* 0.7
-Answer Extraction Llama 3.1 - - - -
-GPT-4o 66.9 81.9 97.2 0.6
-PromptRank 66.3 80.2 96.0 -
+GPT-4o 62.8 77.1 94.3* [90.8, 97.1] 0.1
+No (retrieve for all) 61.2 74.2 93.2* [89.6, 96.2] 0.7
+AE Llama 3.1 - - - -
+GPT-4o 66.9 81.9 97.2 [95.2, 98.9] 0.6
+PromptRank 66.3 80.2 96.0 [94.2, 98.4] -
 Retrieval Source Abs + TB + SP - - - -
-TB + SP 63.4 76.8 94.8 0.1
-Abs (no retrieval) 62.8 77.1 94.3* 0.1
+TB + SP 63.4 76.8 94.8* [91.7, 97.3] 0.1
+Abs (no retrieval) 62.8 77.1 94.3* [90.8, 97.1] 0.1
 Answer Overlap BERTScore - - - -
-LERC 64.5 75.9 95.5 0.1
+LERC 64.5 75.9 95.6 [92.7, 97.9] 0.1
 Granular Level Sentence - - - -
-Summary 49.3 59.6 84.7* 0.5
+Summary 49.3 59.6 84.7* [79.6, 89.5] 0.5
 Table 3: Ablation study of PlainQAFact on PlainFact, analyzing the impact of individual components in PlainQAFact.
-Abs: abstract; TB: Textbooks; SP: StatPearls.
-AUC represents AUC-ROC.
+$\tau$: Kendall’s $\tau$; AE: answer extraction; Abs: abstract; TB: Textbooks; SP: StatPearls; AUC: AUC-ROC.
 The first row of each component setting represents our best combination.
-* indicates an improvement of our metric over prior work with 95% confidence interval.
+* indicates a statistically significant difference compared to PlainQAFact ($p<0.01$).
 We run each setting for five times and report the standard deviation (std.).
@@
 PlainQAFact consists of several modules, including a fine-tuned classifier, QA modules, a retrieval function, and an answer overlap evaluation process.
-To understand the contribution of each component, we conduct an ablation study on PlainFact5.3).
+To understand the contribution of each component, we conduct an ablation study on PlainFact.
 Table 3 summarizes the ablation results, where each component of PlainQAFact is individually modified to measure its impact on overall performance.
 Fine-tuned Classifier Table 3 shows that removing the classifier and retrieving for every sentence does not improve factual consistency assessment.
+PlainQAFact statistically significantly outperforms the settings using GPT-4o as the classifier and the setting without a classifier ($p<0.01$).
 Performance drops across all criteria (Kendall’s $\tau$ from 65.7 to 61.2; Pearson from 81.3 to 74.2; AUC-ROC from 96.4 to 93.2), and the std.
@@
 These findings support the use of a lightweight fine-tuned classifier to trigger retrieval only when necessary.
-Answer Extraction For extracting gold answers from plain language summaries, the LLM-based extractors are better than PromptRank [40].
-GPT-4o achieves the best overall scores (Kendall’s $\tau$=66.9, Pearson=81.9, AUC-ROC=97.2), with small gains over our Llama-based method.
-PromptRank performs better on Kendall’s $\tau$ (increases 0.6), but fails on other two criteria compared to our best combination.
-Although GPT-4o achieves slightly higher performance, the marginal gain does not justify the increased cost, making Llama 3.1 a more practical and cost-effective choice for large-scale factual consistency evaluation.
-Component Method Choice Kendall’s $\tau$ Pearson AUC std.
-Source Simplification Abs 63.5 79.0 94.8 0.2
-Elaborative Explanation Abs 32.8 38.2 72.8* 0.2
-TB 31.2 37.2 70.6* 0.1
-SP 33.3 37.8 72.5 1.7
-TB + SP 36.7 42.0 75.4 0.1
-Abs + TB 40.3 47.0 78.2 0.2
-Abs + SP 40.5 47.4 78.3 0.2
-Abs + TB + SP 44.2 51.3 81.0 0.1
-Full Dataset Abs 62.8 77.1 94.3* 0.1
-TB 60.8 73.8 93.0* 0.1
-SP 61.3 74.2 93.3* 0.1
-TB + SP 63.4 76.8 94.8 0.1
-Abs + TB 64.3 79.7 95.4 0.2
-Abs + SP 65.0 80.3 95.9 0.2
-Abs + TB + SP 65.7 81.3 96.4 0.2
+Answer Extraction For extracting gold answers from plain language summaries, the LLM based extractors perform better than PromptRank [40].
+GPT-4o achieves the best overall scores (Kendall’s $\tau$=66.9, Pearson=81.9, AUC-ROC=97.2), with only small gains over our Llama-based method.
+In PlainFact, PromptRank obtains a slightly lower Pearson correlation than PlainQAFact.
+However, in the ablation study on the CELLS dataset in Table 10, PlainQAFact significantly outperforms PromptRank in both perturbation settings (see details in  H).
+Overall, we do not find a statistically significant difference between PlainQAFact and the other answer extraction methods.
+PlainQAFact is statistically tied with PromptRank (AUC-ROC=$96.0$, $p=0.34$) and with GPT-4o AE (AUC-ROC=$97.2$, $p=1.0$).
+Although GPT-4o achieves slightly higher performance overall, the marginal gain does not justify the increased cost, making Llama 3.1 a more practical and cost-effective choice for large-scale factual consistency evaluation.
+Component Method Choice $\tau$ Pearson AUC (99% CI) std.
+Simplification Abs 63.5 79.0 94.8 [91.6, 97.4] 0.2
+Explanation Abs 32.8 38.2 72.7* [65.8, 79.2] 0.2
+TB 31.2 37.2 74.6* [68.0, 80.9] 0.1
+SP 33.3 37.8 72.5* [65.9, 79.1] 1.7
+TB + SP 36.7 42.0 75.4* [68.8, 81.7] 0.1
+Abs + TB 40.3 47.0 78.1* [71.8, 84.1] 0.2
+Abs + SP 40.5 47.4 78.3* [71.7, 84.2] 0.2
+Abs + TB + SP 44.2 51.3 81.0 [74.9, 86.7] 0.1
+Full Dataset Abs 62.8 77.1 94.3* [90.8, 97.1] 0.1
+TB 60.8 73.8 94.0* [90.7, 96.7] 0.1
+SP 61.3 74.2 93.3* [89.6, 96.4] 0.1
+TB + SP 63.4 76.8 94.8* [91.7, 97.3] 0.1
+Abs + TB 64.3 79.7 95.4* [92.2, 97.9] 0.2
+Abs + SP 65.0 80.3 95.9 [92.9, 98.2] 0.2
+Abs + TB + SP 65.7 81.3 96.4 [93.6, 98.5] 0.2
 Table 4: Ablation study results on the PlainFact, evaluating how different retrieval sources affect various information types using PlainQAFact.
@@
 Overall, the numbers of factual summaries that only include source simplification and elaborative explanation sentences are 198 and 191 respectively, and 198 and 194 for non-factual summaries.
-Abs: abstracts; TB: Textbooks; SP: StatPearls.
-AUC represents AUC-ROC.
-* indicates an improvement of Abs + TB + SP over other settings with 95% confidence interval.
+Simplification: Source Simplification; Explanation: Elaborative Explanation; $\tau$: Kendall’s $\tau$; Abs: abstracts; TB: Textbooks; SP: StatPearls; AUC: AUC-ROC.
+* indicates a statistically significant difference compared to Abs+TB+SP ($p<0.01$).
 We run each setting for five times and report the standard deviation (std.) in the brackets.
-Combined Domain Resources for Retrieval The highest performance is achieved when combining all three sources (Abs + TB + SP).
+Combined Domain Resources for Retrieval As shown in Table 3, combining all three evidence sources (Abs+TB+SP) significantly outperforms the other two settings ($p<0.01$).
 This suggests that comprehensive retrieval, which includes the original source abstracts, improves overall factual consistency assessment.
-Compared to using only external sources (TB + SP), performance drops by 2.3/4.5/1.6 across the three criteria.
-Using only abstracts (i.e., no elaborative retrieval) also lowers performance (2.9/4.2/2.1).
+Table 4 shows the retrieval ablation study results in detail.
+In the full dataset setting, Abs+TB+SP achieves AUC-ROC=$96.4$ and Pearson=$81.3$, outperforming all retrieval variants that use only a subset of the retrieval corpus, including TB+SP (AUC-ROC=$94.8$), TB only ($94.0$), SP only ($93.3$), and Abs+TB ($95.4$).
 To further examine the effectiveness of retrieval for source simplification and elaborative explanation, Table 4 breaks down the results as follows:
@@
 Using abstracts as sole source texts achieves strong performance (63.5/79.0/94.8) across all three criteria.
-(2) Elaborative explanation: For summaries containing only explanation sentences, abstracts alone are not sufficient (32.8/38.2/72.8), performing much worse than in simplification cases.
-Adding abstracts with external sources improves performance, with the best results obtained by combining all three sources (Abs + TB + SP).
+(2) Elaborative explanation: For summaries containing only explanation sentences, abstracts alone are not sufficient, performing much worse than in simplification cases.
+Adding abstracts with external sources improves performance, with the best results obtained by combining all three sources.
+Abs+TB+SP outperforms the abstract-only baseline Abs (AUC-ROC=$72.7$) and all partial-view variants (TB+SP, TB, SP, Abs+TB, and Abs+SP).
+The gain over Abs is substantial, with $\Delta\mathrm{AUC-ROC}=8.3$ on (99% CI [3.7, 13.0]), and similarly large improvements in correlation.
 (3)Full dataset: Retrieval from StatPearls provides better results than retrieval from Textbooks, highlighting the importance of using high-quality, domain-specific knowledge bases for PLS evaluation.
 Overall, the best combination across all settings is Abs+TB+SP (65.7/81.3/96.4).
+Relative to the abstract-only baseline, Abs (AUC-ROC=$94.3$), the gain in discrimination is $\Delta\mathrm{AUC-ROC}=2.1$ (99% CI [0.5, 4.2], where $\Delta$ is Abs+TB+SP minus Abs).
 These findings show that effective retrieval for factual consistency should include both the source abstract and reliable external medical references, especially when summaries contain elaborative explanations.
-Answer Overlap Evaluation BERTScore outperforms the LERC approach introduced in QAFactEval [21], with gains of 1.2/5.4/0.9 across the three evaluation criteria.
-This performance gap can be attributed to the fundamental differences between the two metrics.
+Answer Overlap Evaluation BERTScore ties with the LERC approach introduced in QAFactEval [21], showing no significant difference in AUC-ROC ($p=0.29$).
+BERTScore achieves slightly higher nominal values in both correlation metrics.
+There are fundamental differences between the two metrics.
 LERC is a learned model trained to predict human factual consistency scores, which may reduce its generalizability to new domains or sentence types.
@@
 Sentence-level vs.
-Summary-level Evaluation To further examine the effect of input granularity in factual consistency evaluation, we deactivate the sentence-splitting function and instead pair each plain language summary with an abstract to evaluate factual consistency at the summary level.
-Our best sentence-level evaluation achieves an AUC-ROC of 96.4, while the summary-level setting performs worse across all three criteria (Kendall’s $\tau$=49.3, Pearson=59.6, AUC-ROC=84.7; std.
+Summary-level Evaluation To further examine how input granularity affects factual consistency evaluation, we deactivate the sentence-splitting function and instead pair each plain language summary with its abstract to evaluate factual consistency at the summary level.
+Our best sentence level evaluation achieves an AUC-ROC of $96.4$, which is significantly higher than the summary level setting ($p<0.01$).
+The summary level setting performs worse across all three criteria (Kendall’s $\tau$=$49.3$, Pearson=$59.6$, AUC-ROC=$84.7$; std.
 $0.5$).
 Evaluating at the summary level introduces broader contextual dependencies, which can omit sentence-specific factual errors or create inaccurate entailments.
-These results suggest that processing one sentence at a time makes QA-based question generation and answering more focused and reduces noise from unrelated context.
+These results suggest that processing one sentence at a time makes QA-based factual consistency evaluation more focused and reduces noise from unrelated context.
 Correct Case Plain Language Sentence: Limits to the availability of SSB in schools (e.g. replacing SSBs with water in school cafeterias). [54]
-Scientific Abstract: Frequent consumption of excess amounts of sugar-sweetened beverages (SSB) is a risk factor for obesity, type 2 diabetes, cardiovascular disease and dental caries… [52]
+Scientific Abstract: Frequent consumption of excess amounts of sugar‐sweetened beverages (SSB) is a risk factor for obesity, type 2 diabetes, cardiovascular disease and dental caries… [54]
 Generated Question: Limits to the availability of what in schools?
@@
 Retrieval Quality Plain Language Sentence: The patch is a small, thin, adhesive square that is applied to the skin. [55]
-Scientific Abstract: Users of the norelgestromin-containing patch reported more breast discomfort, dysmenorrhea, nausea, and vomiting.
-In the levonorgestrel-containing patch trial, patch users reported less vomiting, headaches, and fatigue…[53]
+Scientific Abstract: Users of the norelgestromin‐containing patch reported more breast discomfort, dysmenorrhea, nausea, and vomiting.
+In the levonorgestrel‐containing patch trial, patch users reported less vomiting, headaches, and fatigue…[55]
 Generated Question: The patch is a small, thin, what kind of square applied to the skin?
@@
 [54]
-Scientific Abstract: To assess the effects of environmental interventions (excluding taxation) on the consumption of sugar-sweetened beverages and sugar-sweetened milk, diet-related anthropometric measures and health outcomes, and on any reported unintended consequences or adverse outcomes…[52]
+Scientific Abstract: To assess the effects of environmental interventions (excluding taxation) on the consumption of sugar‐sweetened beverages and sugar‐sweetened milk, diet‐related anthropometric measures and health outcomes, and on any reported unintended consequences or adverse outcomes…[54]
 Generated Question: What officials, business people and health professionals implementing such measures should work together with researchers to find out more about their effects in the short and long term?
-Retrieved Knowledge: Implementation should be accompanied by high-quality evaluations using appropriate study designs, with a particular focus on the long-term effects of approaches suitable for large-scale implementation.
-Final Answer: long-term effects
+Retrieved Knowledge: Implementation should be accompanied by high‐quality evaluations using appropriate study designs, with a particular focus on the long‐term effects of approaches suitable for large‐scale implementation.
+Final Answer: long‐term effects
 Table 5: Error analysis of PlainQAFact with intermediate metric outputs.
@@
 Color legend: extracted answer, answer origin, correct answer, incorrect answer.
-5.4 Error Analysis
+5.4 Error Analysis on PlainFact
 To analyze cases where PlainQAFact fails on the PlainFact benchmark, we categorize one correct example and three types of errors in Table 5.
@@
 Addressing these challenges requires further improvements in classification, retrieval, and domain-specific question generation.
+5.5 Error Analysis on FareBio
+Type Target PLS Sentence QA Extraction Analysis Rationale
+NF This performance was identical in the comparative analysis subgroup and was within the range of physicians at different levels of expertise: 0.86-0.97 and 0. Ref: physicians
+Q: This performance was identical in the comparative analysis subgroup and was within the range of what at different levels of expertise?
+Pred: physicians
+The non-factual error is a value mistake.
+The target sentence incorrectly truncates the specificity range to “0,” which does not appear in the source and therefore being labeled as non-factual.
+However, our answer extractor fails to identify the numerical values as gold answer to verify its factual consistency.
+NF Imaging also showed chronic compression of the coeliac axis with compensatory hypertrophy of the gastroduodenal artery. Ref: gastroduodenal artery
+Q: Imaging also showed chronic compression of the coeliac axis with compensatory hypertrophy of what artery?
+Pred: gastroduodenal
+The target sentence correctly restates information that is directly present in the source, but it is labeled as non-factual in FareBio.
+F In summary, the article discusses the potential impact of COVID-19 on the musculoskeletal system… and emphasizes the need for further research to understand the underlying mechanisms and develop appropriate interventions.
+Ref: interventions
+Q: In summary, the article discusses the potential impact of COVID-19 on the musculoskeletal system, particularly in the context of aging and Long-COVID, and emphasizes the need for further research to understand the underlying mechanisms and develop appropriate what?
+Pred: multidisciplinary approach
+The predicted answer comes from the original source. This target PLS sentence is classified as a simplification; however, the term “interventions” does not appear in the source.
+F Some patients experience persistent symptoms after recovering from the initial infection, a condition known as “Long-COVID.” Ref: infection
+Q: Some patients experience persistent symptoms after recovering from what condition known as Long-COVID?
+Pred: SARS-CoV-2
+In the source, SARS CoV 2 always appears together with the word “infection.” Our QA model captures the relevant answer, but our answer overlap evaluation approach does not align this prediction with the gold answer, which results in a lower PlainQAFact.
+Table 6: Error analysis of PlainQAFact on the FareBio dataset. NF/F = non-factual or factual; Ref = gold/extracted answer; Q = generated question; Pred = predicted answer.
+As shown in Table 6, we provide several failure cases on the FareBio dataset based on the results in Table 2.
+We rank the PlainQAFact scores for all summaries and select two examples with the lowest scores among those labeled NF and two examples with the lowest scores among those labeled F.
+The first NF example in Table 6 shows a true factual error.
+The target plain language sentence changes the numerical range and introduces the value zero, which does not appear in the source.
+However, our answer extractor does not reliably extract the exact numbers, so PlainQAFact does not capture the fact errors.
+The second NF example shows a label noise in FareBio.
+The target sentence restates information that is clearly present in the source, but the sentence is labeled as NF.
+Third, we find pipeline errors even when the FareBio label is factual.
+In the third example, the target sentence correctly summarizes the current knowledge and is classified as simplification by our classifier, but the answer extractor extracts the word “interventions,” which does not appear in the source and thus leads to a low score.
+In the fourth example, the QA module predicts “SARS CoV 2” as the answer, which is consistent with the “infection” described in the plain language sentence, but the final answer overlap evaluation step cannot match this prediction to the gold answer “infection.” Overall, we conclude that most errors of our PlainQAFact come from the answer extraction and answer overlap evaluation stages in the FareBio dataset, while the dataset contains noisy factuality labels.
+To mitigate these errors (§5.4, §5.5), we provide several directions for future work.
+First, misclassification of plain language sentences by the factuality type classifier can be reduced by training on more annotated examples or use advanced low-resource model tuning strategies for text classification [56].
+Second, retrieval errors can be mitigated by conducting precise filtering method to select the most relevant snippets, such as limiting the length of retrieved snippets and chunk size, using a combination of dense and sparse retrieval, and expanding the domain corpus (e.g., PubMed) to cover more specialized topics.
+Third, to handle unanswerable or vague questions, future work can improve the question filtering stage that removes questions that do not have clear grounding in both source retrieved context, or that are too general to support reliable factual checks using heuristic approaches.
+Finally, for the answer overlap evaluation module, more advanced entailment-based models can be applied to replace the BERTScore semantic matching to check whether the predicted answer is consistent with the gold answer after simple normalization of entities and synonyms.
+We report the quantitative error analysis of FareBio in  J.
 6 Discussion and Conclusion
-Our study advances the assessment of hallucinations in the field of PLS by supplying both the first domain expert-annotated biomedical PLS benchmark PlainFact and a novel retrieval-augmented QA-based factual consistency evaluation metric PlainQAFact.
-Unlike existing biomedical domain corpora that either lack fine-grained labels for elaborative content [54] or focus solely on text simplification (e.g., FactPICO[9]), PlainFact captures sentence-level distinctions, including simplification versus explanation, functional roles, and explicit alignment to source sentences.
-This granularity not only enables precise error analysis of hallucinations introduced for clarity, but also provides a reusable framework for researchers in other disciplines (e.g., legal, technical) to replicate our annotation protocol and build high-quality factual consistency datasets tailored to their specialized texts.
-By offering these data and detailed annotation protocol (A), we anticipate community-driven extensions, such as multilingual adaptations or integration with domain-specific ontologies that will democratize development of trustworthy summarization models.
-Building on this resource, PlainQAFact differentiates from prior QA-based metrics (e.g., QAFactEval[21], QuestEval[50]) by first classifying whether a sentence requires external verification given elaborative explanations are often ignored by metrics relying solely on source text, and improve the evaluation efficiency without retrieving for every plain language summary instance.
-Then, retrieving targeted domain knowledge for each elaborative explanation before posing and answering fact-checking questions.
-The ablations demonstrate that selective retrieval yields substantial gains in Kendall’s $\tau$, Pearson correlation, and AUC-ROC over both end-to-end LLM judges[41] and alignment-based scorers[19], highlighting that judicious incorporation of external evidence is essential for robust factual consistency assessment.
-Beyond its empirical strengths, PlainQAFact serves as a plug-and-play evaluation metric for future summarization systems.
+Our study advances the assessment of hallucinations in the field of PLS by supplying both the first domain expert-annotated biomedical PLS benchmark PlainFact and a novel retrieval‐augmented QA‐based factual consistency evaluation metric PlainQAFact.
+Unlike existing biomedical domain corpora that either lack fine‐grained labels for elaborative content [57] or focus solely on text simplification (e.g., FactPICO[9]), PlainFact captures sentence‐level distinctions, including simplification versus explanation, functional roles, and explicit alignment to source sentences.
+This granularity not only enables precise error analysis of hallucinations introduced for clarity, but also provides a reusable framework for researchers in other disciplines (e.g., legal, technical) to replicate our annotation protocol and build high‐quality factual consistency datasets tailored to their specialized texts.
+By offering these data and detailed annotation protocol (A), we anticipate community‐driven extensions, such as multilingual adaptations or integration with domain‐specific ontologies that will democratize development of trustworthy summarization models.
+Building on this resource, PlainQAFact differentiates from prior QA‐based metrics (e.g., QAFactEval[21], QuestEval[50]) by first classifying whether a sentence requires external verification given elaborative explanations are often ignored by metrics relying solely on source text, and improve the evaluation accuracy without retrieving for every plain language summary instance.
+Then, retrieving targeted domain knowledge for each elaborative explanation before posing and answering fact‐checking questions.
+The ablations demonstrate that selective retrieval yields substantial gains in Kendall’s $\tau$, Pearson correlation, and AUC‐ROC over both end‑to‑end LLM judges[41] and alignment‐based scorers[19], highlighting that judicious incorporation of external evidence is essential for robust factual consistency assessment.
+Beyond its empirical strengths, PlainQAFact serves as a plug‑and‑play evaluation metric for future summarization systems.
 Developers can leverage our classifier to classify sentences, apply retrieval only where necessary, thus containing computational costs and obtain interpretable question–answer pairs that manifest non-factual hallucinations.
-More importantly, PlainQAFact is developed with the open-source backbone model (i.e., Llama 3.1), ensuring broad use with transparency, reproducibility, and fewer computational costs.
-Looking forward, extending our proposed evaluation method to interactive human–AI workflows, where clinicians or subject-matter experts verify QA outputs on-the-fly, could further enhance trust and adoption in high-stakes domains, including medicine, healthcare, or law.
+More importantly, PlainQAFact is developed with the open-source backbone model (i.e., Llama 3.1), ensuring broad use with transparency, reproducibility, and lower computational costs.
+Looking forward, extending our proposed evaluation method to interactive human–AI workflows, where clinicians or subject‐matter experts verify QA outputs on‐the‑fly, could further enhance trust and adoption in high‐stakes domains, including medicine, healthcare, or law.
 Despite these promising results, we consider the following limitations of this study.
@@
 We suggest to balance wisely based on the trade-offs we report in this study regarding the evaluation time and accuracy.
+PlainQAFact is a research-oriented metric that prioritizes interpretability and transparency over speed.
+(3) Given the fair IRA scores, especially for the factuality type field, we acknowledge that our classification method is not fully reliable.
+The annotated dataset reflects subjective judgments, particularly when distinguishing source simplification from elaborative explanation, and the training set is limited in size.
+We advise users to apply our pretrained classifier with caution when using it for later factual consistency evaluation steps.
 Acknowledgment
@@
 For every sentence-abstract pair, the annotators are required to label three features: External, Category, and Relation, with the appropriate labels.
-Step 1: Sentence Annotation: Compared to the scientific abstract, analyze each sentence of the plain language summary across three dimensions: external information, category, and relation.
+Step 1: Sentence Annotation Instructions Compared to the scientific abstract, analyze each sentence of the plain language summary across three dimensions: external information, category, and relation.
 1.
@@
 In Step 2, we provide explanations of the existing sentence-level indexes for plain language sentences and scientific abstracts.
-Step 2: Using the Annotation Spreadsheet: You will work within a structured spreadsheet containing the segmented sentences from both the summaries and the scientific abstracts.
+Step 2: Using the Annotation Spreadsheet You will work within a structured spreadsheet containing the segmented sentences from both the summaries and the scientific abstracts.
 1. Target_Summary_ID: A unique identifier for each plain language summary. Sentences from the same plain language summary share the same ID.
@@
 For each row of the spreadsheet, they need to annotate three columns: “External,” “Category,” and “Relation.”
-<<TO BE ANNOTATED>> External: Mark “Yes” or “No” to indicate if the sentence in the ‘Target_Sentence’ column contains external information.
+Annotation Fields to Be Completed <<TO BE ANNOTATED>> External: Mark “Yes” or “No” to indicate if the sentence in the ‘Target_Sentence’ column contains external information.
 <<TO BE ANNOTATED>> Category: Choose the most fitting category of the sentence from the list (Definition, Background, Example, Method/Result, Other).
@@
 We select candidates from freelancers with a medical education background, and only those who pass the screening test are finalized as annotators.
+To ensure annotators fully understand the context and task requirements, we provide comprehensive annotation training and a screening test prior to the annotation process.
+We select candidates from freelancers with a medical education background, and only those who pass the screening test are finalized as annotators.
+Appendix B Dataset Examples
 Feature Annotation
@@
 Relation s15_16,s15_17,s15_20
-Corresponding Abstract <<s15_16>>For treatment failure, the difference between doxycycline and tetracycline is uncertain (very low-certainty evidence).
-<<s15_17>>Doxycycline compared to tetracycline may make little or no difference in resolution of fever within 48 hours (risk ratio (RR) 1.14, 95% confidence interval (CI) 0.90 to 1.44, 55 participants; one trial; low-certainty evidence) and in time to defervescence (116 participants; one trial; low-certainty evidence).
-<<s15_20>>For most outcomes, including treatment failure, resolution of fever within 48 hours, time to defervescence, and serious adverse events, we are uncertain whether study results show a difference between doxycycline and macrolides (very low-certainty evidence).
+Corresponding Abstract <<s15_16>>For treatment failure, the difference between doxycycline and tetracycline is uncertain (very low‐certainty evidence).
+<<s15_17>>Doxycycline compared to tetracycline may make little or no difference in resolution of fever within 48 hours (risk ratio (RR) 1.14, 95% confidence interval (CI) 0.90 to 1.44, 55 participants; one trial; low‐certainty evidence) and in time to defervescence (116 participants; one trial; low‐certainty evidence).
+<<s15_20>>For most outcomes, including treatment failure, resolution of fever within 48 hours, time to defervescence, and serious adverse events, we are uncertain whether study results show a difference between doxycycline and macrolides (very low‐certainty evidence).
 Table 7: Examples of our curated dataset.
@@
 Conversely, a label of “no” shows that the sentence is a simplification that can be validated using only the source abstract.
-Appendix B Dataset Examples
 In Table 7, we presents three representative examples from PlainFact.
@@
 Additionally, indexes have been created for both the plain language sentences and the abstract sentences to facilitate annotation.
-Appendix C LLM Prompts
-We utilize two types of prompts to guide LLMs through two stages of FC evaluation: classification and answer extraction.
+Appendix C Qualitative Analysis of PlainFact
+ID Target PLS Sentence A1 A2 Final Rationale
+Ex1 Obesity is associated with many health problems and a higher risk of death.
+“Health problems” and “higher risk of death” are not in the original abstract.
+No association between obesity and health/death issues exists in the source.
+Ex2 It can be hard for the person to gain control of the eye, which can cause distress in social situations.
+Most relevant source (from A1): “Spontaneously manifest DVD can be difficult to control and often causes psychosocial concerns.” However, no direct correlation between eye control and DVD exists; social situations are not equal to psychosocial concerns.
+Ex3 Most participants were middle-aged.
+A2 provides supporting source: “Applicability of evidence was limited to middle-aged participants who were relatively free of co-morbidity and were undergoing elective abdominal surgery.” This sentence is supported by the source.
+Ex4 In some people, winter blues becomes depression, which seriously affects their daily lives.
+A2 highlights relevant source: “SAD is a seasonal pattern of recurrent major depressive episodes that most commonly occurs during autumn or winter and remits in spring.” However, no causal relation between winter blues and depression in source.
+Ex5 RDS is the most common cause of disease and death in babies born before 34 weeks’ gestation.
+Most relevant source: “RDS is the single most important cause of morbidity and mortality in preterm infants.” However, “babies born before 34 weeks’ gestation” is not mentioned in the abstract.
+Table 8: Qualitative analysis of PlainFact annotation examples. $A1/A2$ = Annotator 1 or 2; ✓/✗ indicate yes or no judgments for the “Need External Information?” field.
+Given the fair inter-rater agreement for the “Need External Information?” field in PlainFact, we conduct a targeted qualitative error analysis.
+We randomly sample one representative plain language sentence from each of five annotation batches where the two annotators disagree.
+Two annotators independently judge whether external information is required, and a third judge provides a brief rationale.
+As shown in Table 8, most disagreements fall into two categories: (i) differing interpretations of whether an elaborative explanation introduces new information beyond the source, and (ii) failures to locate or recognize supporting evidence in the source abstract.
+These examples highlight that the “Need External Information?” decision often depends on subtle judgments about sufficiency and specificity of source coverage, suggesting that clearer guidelines and additional training examples may further improve annotation consistency.
+Appendix D Data and Indexing Statement
+In this study, we follow the same protocol as used in MedRAG [37] to conduct sentence-level retrieval of the PLS sentences.
+Specifically, we use top-K as 3 across all of our experiments on both StatPearls [^9] https://huggingface.co/datasets/MedRAG/statpearls and medical Textbooks [^10] https://huggingface.co/datasets/MedRAG/textbooks [38] provided in the MedRAG repository [^11] https://github.com/Teddy-XiongGZ/MedRAG?tab=readme-ov-file#corpus.
+Specifically, the medical Textbooks corpus comprises 18 widely used medical textbooks (e.g., for USMLE preparation) originally collected by Jin et al.
+[38], while StatPearls corpus consists of point-of-care clinical decision support articles.
+The authors of MedRAG collect 9,330 publicly available articles from the NCBI Bookshelf [^12] https://www.ncbi.nlm.nih.gov/books/NBK430685/, which permits redistribution for non-commercial research purposes.
+The StatPearls corpus applies a hierarchical approach for chunking, treating each paragraph in an article as a snippet and splice all relevant hierarchical headings to form the corresponding title.
+The medical Textbooks is processed into chunks of no more than 1,000 characters.
+These corpus are made publicly available for research replication via the MedRAG HuggingFace repository [^13] https://huggingface.co/MedRAG under the original authors’ distribution terms.
+We do not redistribute raw StatPearls and Textbooks content beyond the MedRAG release and rely exclusively on their officially published corpus to ensure replicability.
+Researchers wishing to replicate the corpus index should (i) access the same StatPearls [^14] https://huggingface.co/datasets/MedRAG/statpearls and Textbooks [^15] https://huggingface.co/datasets/MedRAG/textbooks datasets, (ii) ensure their use complies with the copyright and permissions of the source license of MedRAG [37] and MedQA [38].
+Following MedRAG, we use the same RecursiveCharacterTextSplitter from the LangChain [^16] https://www.langchain.com/ library to perform chunking.
+We directly retrieve top 3 most relevant snippets from both corpus.
+For the retriever, we use MedCPT [36] (specifically the ncbi/MedCPT-Query-Encoder [^17] https://huggingface.co/ncbi/MedCPT-Query-Encoder version, commit hash: d83a36c) for dense retrieval with CLS token pooling.
+Documents are indexed using FAISS [58] with an IndexFlatIP configuration (exact inner product search, , no quantization or clustering) for 768-dimensional embeddings.
+The retrieval system computes the inner product between query and document vectors to identify the most relevant snippets.
+This matches the default MedRAG implementation for MedCPT.
+Appendix E Reproducibility
+E.1 Fine-tuning QG Model
+We fine-tune the QG model (i.e., BART-large) with the same datasets used in QAFactEval [39], including SQuAD [42] and QA2D [43].
+We use a batch size of 16, a learning rate of 3e-5, maximum sequence length of 512 tokens, and two training epochs for QG model fine-tuning.
+E.2 Fine-tuning Factuality Type Classifier
+As mentioned in Section 3.2.1, we fine-tune a classifier using PubMedBERT-base model through our curated PlainFact dataset.
+Specifically, we set the random seed to 42 to split PlainFact into training, validation, and test sets.
+We tune the PubMedBERT-base model for three epochs with early stopping under the validation set.
+The final fine-tuning accuracy of the classifier on the test set is 0.77.
+The same seed (42) is also used to sample 200 summary-abstract pairs from the CELLS test set for the comparison in Table 2.
+The classifier used in PlainQAFact is fine-tuned on PlainFact with a batch size of 32 for 10 epochs.
+We apply early stopping during training based on validation loss.
+E.3 General Details
+For other open-source models used in this study, we apply the default version.
+For the QA model, we apply the original Electra-large model [^18] Can be downloaded here: https://github.com/salesforce/QAFactEval/blob/master/download_models.sh used in QAFactEval [39].
+All experiments we conduct are under one NVIDIA A100 GPU with 40 GB GPU memory.
+We employ the Natural Language Toolkit (NLTK) [^19] https://www.nltk.org/ to split the plain language summaries into sentences using the punkt package.
+For question filtering strategy, we follow the same protocol in QAFactEval.
+The temperature is set to 0 for GPT-4o and 0.01 for Llama 3.1.
+The Llama 3.1 model used in AE also uses a temperature of 0.01.
+The maximum input length is set to 512 for all models in PlainQAFact.The QA model (i.e., Electra-large) used in our PlainQAFact is directly downloaded from QAFactEval [39].
+However, we set its maximum input length to 512 (from 364) tokens to incorporate more source context.
+Similarly, to ensure that the answers extracted during the AE stage are valid (i.e., within the 512-token limit to maintain consistency with the subsequent QA model), we configure the LLMs’ input lengths to 512 tokens.
+In MedCPT retrieval, we set k (the number of retrieved snippets) to 3, ensuring the retrieved information remains within a short context window for the subsequent QA process (E).
+In the QF stage, we follow QAFactEval [39] and remove the unanswerable questions using the pre-trained Electra-large QA model [^20] Download here: https://github.com/salesforce/QAFactEval/blob/master/download_models.sh.
+We apply the default setting (i.e., RoBERTa-base) for AlignScore [19] and SummaC-Conv for SummaC [16].
+For GPT-4o experiments as a factual consistency evaluator, we set the temperature as 1.0 in this paper.
+For the FareBio dataset, each plain language sentence is annotated with two binary labels: faithfulness and factual hallucination.
+In the released data, there are no sentences that are faithful but factually wrong.
+For the dataset pre-processing, we treat a sentence as not factual only when both labels are “no.” If either faithfulness or factual hallucination is “yes,” we treat the sentence as factual.
+To construct summary level labels, we first apply this rule to every sentence and then group sentences by their original FareBio summary.
+If a summary contains at least one sentence that is not factual, we label the whole summary as not factual, which produces 33 summaries.
+If all sentences in a summary are factual, we label the summary as factual, which produces 174 summaries.
+For the subset of 53 valid summaries used in our explanation focused evaluation, we further keep only those sentences that meet our filtering criteria for factual elaborative explanations and then aggregate them back into summaries.
+For the computational costs of GPT-4o, it costs about \$0.01 per plain language summary during the factual consistency evaluation on the CELLS dataset experiment, combining input and output tokens, and takes about 2 seconds per summary.
+In contrast, PlainQAFact takes about 30 seconds per summary on our hardware.
+However, PlainQAFact is able to provide intermediate outputs for each step in the QA based evaluation, which makes our metric more transparent and easier to inspect than the black box GPT-4o evaluation, even in most cases where GPT-4o achieves slightly higher performance than PlainQAFact.
+E.4 LLM Prompts
+We utilize two types of prompts to guide LLMs through two stages of factual consistency evaluation: classification and answer extraction.
 In accordance with the benchmark annotation protocol (A), we employ GPT-4o as a classifier to determine whether a given sentence or summary requires elaborative explanations for factual verification.
-For both stages, we set the max_tokens parameter to 512 and configure the temperature to 0 for GPT-4o and 0.01 for the Llama 3.1 8B Instruc model.
+For both stages, we set the max_tokens parameter to 512 and configure the temperature to 0 for GPT-4o and 0.01 for the Llama 3.1 8B Instruct model.
 Developer
@@
 Example 1: Prompt of GPT-4o as the Classifier.
-For the AE stage, we explore both GPT-4o and Llama 3.1 8B Instruct as backbone models to assess the FC evaluation performance of PlainQAFact.
+For the AE stage, we explore both GPT-4o and Llama 3.1 8B Instruct as backbone models to assess the factual consistency evaluation performance of PlainQAFact.
 Following the task description outlined in QAFactEval [39], we instruct both LLMs to extract potential answer entities from the input PLS.
@@
 Example 3: Prompt of Llama 3.1 8B Instruct as the Answer Extractor.
-Additionally, we report the performance of using Llama 3.1 8B Instruct as a judge to evaluate the FC of a given PLS based on its source scientific abstract.
+Additionally, we report the performance of using Llama 3.1 8B Instruct as a judge to evaluate the factual consistency of a given PLS based on its source scientific abstract.
 System
@@
 Factuality score (only output a numeric score): <score>
-Example 4: Prompt of Llama 3.1 8B Instruct as a FC judge.
+Example 4: Prompt of Llama 3.1 8B Instruct as a factual consistency judge.
 For the LLM-based perturbation of PlainFact and CELLS datasets, we follow the protocol of APPLS [12] on faithfulness criteria.
@@
 Example 5: Prompt of GPT-4o for faithfulness perturbation
-Appendix D Detailed Experiment Settings
-All experiments we conduct are under one NVIDIA A100 GPU with 40 GB GPU memory.
-We employ the Natural Language Toolkit (NLTK) [^5] https://www.nltk.org/ to split the plain language summaries into sentences using the punkt package.
-For the classifier fine-tuning, we set the random seed to 42 to split PlainFact into training, validation, and test sets.
-We tune the PubMedBERT-base model for three epochs with early stopping under the validation set.
-The final fine-tuning accuracy of the classifier on the test set is 0.77.
-The same seed (42) is also used to sample 200 summary-abstract pairs from the CELLS test set for the comparison in Table 2.
-The QA model used in our PlainQAFact is downloaded from QAFactEval [39].
-However, we set its maximum input length to 512 (from 364) tokens to incorporate more source context.
-Similarly, to ensure that the answers extracted during the AE stage are valid (i.e., within the 512-token limit to maintain consistency with the subsequent QA model), we configure the LLMs’ input lengths to 512 tokens.
-In MedCPT retrieval, we set k (the number of retrieved snippets) to 3, ensuring the retrieved information remains within a short context window for the subsequent QA process (D).
-We apply the default setting (i.e., RoBERTa-base) for AlignScore [19] and SummaC-Conv for SummaC [16].
-For GPT-4o experiments as an factual consistency evaluator, we set the temperature as 1.0 in this paper.
-Appendix E Statistical Testing
-Following SummaC [16] and QAFactEval [39], we test whether our proposed metric achieves statistically significant improvements over other methods.
-Given Kendall’s $\tau$ and Pearson correlations do not show the discriminative feature of continuous scores in various evaluation metric, we run a systematic evaluation on four datasets (CELLS, FareBio, FactPICO, and PlainFact) through bootstrap resampling [56] on the AUC-ROC results.
-We compare our best metric to other methods using the confidence intervals as significance level of 0.05 and apply Bonferroni correction [57] following SummaC [16].
-Our results show that PlainQAFact achieves statistically significant improvements over QuestEval [50], SummaC [16], and DAE [48] on the CELLS and PlainFact datasets.
-As shown in Figure 2, PlainQAFact also outperforms other automatic metrics in evaluating elaborative explanation sentences, except for QAFactEval [39].
-This highlights the effectiveness of PlainQAFact for biomedical PLS tasks that include external explanations.
-On the other hand, when evaluating elaborative content in the FactPICO and FareBIO datasets (Figures 3 and 4), we do not find statistically significant differences across metrics.
-While some existing metrics perform similarly in certain settings, our metric consistently shows better performance across all three evaluation criteria.
-Figure 3: Overall performance on summaries containing added information (i.e., elaborative explanations) from FactPICO [9].
-The std.
-of PlainQAFact, Llama 3.1, and GPT-4o are 0.2, 0.2, and 3.3, respectively, based on five runs of each metric (details in  E).
-Figure 4: Overall performance on summaries with elaborative explanations from FareBio [47].
-The std.
-for PlainQAFact, Llama 3.1, and GPT-4o are 0.03, 0.3, and 5.7, respectively, computed over five runs (details in  E.)
-Appendix F Summary-Level Explanation-Only Evaluation
-Similar to our sentence-level explanation annotations in PlainFact, FactPICO [9] and FareBio [47] datasets also provide human-annotated explanation information in plain language summaries.
-According to the results shown in Section 5.2, we also assess the performance of five factual consistency evaluation metrics under FactPICO and FareBio datasets.
-In the FactPICO, the elaborative explanation is defined as “added information.” We first select summaries that include added information, and then we label a summary as “non-factual” if it contains any non-factual added information as determined by annotators.
-We generate a summary-level explanation-only FactPICO dataset consisting of 231 summary-abstract pairs.
-As shown in Figure 3, AlignScore achieves the best performance across all three evaluation criteria.
-However, since the summaries in the FactPICO dataset are generated using LLMs, the factual consistency between the generated summaries and the original source abstracts is not guaranteed.
-For instance, we treat a summary as “factual” indicates that all the added information is factual, but it does not necessarily reflect the factual consistency of other content.
-Therefore, the results presented in Figure 3 may become biased.
-For FareBio [47], we collect explanation sentences based on both “faithfulness” and “factual hallucination” labels in the original dataset.
-As shown in Figure 4, PlainQAFact achieves the best performance compared with all other metrics, showing a clear advantage over GPT-4o.
-Since the FareBio dataset provides sentence-level annotations of factual consistency for each sentence in the plain language summaries, the results are more reliable than those from FactPICO.
-Appendix G Pilot Study on FactPICO
+Appendix F Pilot Study on FactPICO
 To investigate the performance of existing automatic factual consistency evaluation metrics on plain language generation tasks, we employ the recently introduced FactPICO dataset [9].
@@
 To validate this assumption, we conduct pilot studies using four factual consistency evaluation metrics from FactPICO alongside one NLI-based metric: DAE [48], AlignScore [19], SummaC [16], QAFactEval [39], and QuestEval [50].
-G.1 Dataset Pre-processing
+F.1 Dataset Pre-processing
 FactPICO provides span-level annotations for LLM-generated summaries, assessing whether the additional information is present and determining its factuality, labeled as either “yes” or “no.” We first remove all special identification tags from the abstracts, such as “ABSTRACT.BACKGROUND,” “ABSTRACT.RESULTS,” and “ABSTRACT.CONCLUSIONS.” We then deduplicate the generated summaries, while retaining duplicates within abstracts, as each abstract has three summaries generated by different LLMs, and each abstract may have varying numbers of generated summaries, from none to multiple.
 This pilot study focuses on two key research questions: evaluating the effectiveness of existing factuality metrics in assessing added information and non-factual added information.
-G.2 Experiments and Results
+F.2 Experiments and Results
 We conduct experiments using the processed FactPICO dataset to conduct pilot studies that assess the performance of existing factual consistency evaluation metrics in detecting added information in plain language summarization tasks.
@@
 Nevertheless, based on the results of RQ 1 and 2, our findings suggest that existing factual consistency evaluation metrics have limited capacity to accurately distinguish between factual and non-factual added information in plain language summarization tasks.
+Appendix G Statistical Testing
+Following SummaC [16] and QAFactEval [39], we test whether our proposed metric achieves statistically significant improvements over other methods.
+Given Kendall’s $\tau$ and Pearson correlations do not show the discriminative feature of continuous scores in various evaluation metric, we run a systematic evaluation on four datasets (CELLS, FareBio, FactPICO, and PlainFact) through bootstrap resampling [59] on the AUC-ROC results.
+We compare our best metric to other methods using the confidence intervals as significance level of 0.01 and apply Bonferroni correction [60] following SummaC [16].
+Our results show that PlainQAFact achieves statistically significant improvements over QuestEval [50], SummaC [16], and DAE [48] on the CELLS and PlainFact datasets.
+As shown in Figure 2, PlainQAFact also outperforms other automatic metrics in evaluating elaborative explanation sentences, except for QAFactEval [39].
+This highlights the effectiveness of PlainQAFact for biomedical PLS tasks that include external explanations.
+For hypothesis testing, we treat AUC-ROC as the primary metric.
+For each baseline $b$ we conduct a paired bootstrap test on the difference in AUC-ROC, $\Delta^{(k)}=\mathrm{AUC}^{(k)}_{\text{{PlainQAFact}}}-\mathrm{AUC}^{(k)}_{b}$ for bootstrap replicate $k$.
+The one-sided $p$-value for the alternative “PlainQAFact >> baseline” is estimated as $p=\Pr(\Delta\leq 0)$ over the 10,000 replicates.
+Across the seven baselines we apply Holm-Bonferroni [52] correction to control the family-wise error rate at $\alpha=0.01$.
+Appendix H Entity Swap Perturbation
+Dataset Metrics Kendall’s $\tau$ Pearson AUC (99% CI)
+CELLS Llama 3.1 24.6 29.9 65.7* [57.8, 73.3]
+GPT-4o 69.3 74.3 98.4 [96.4, 99.7]
+QAFactEval 55.9 64.6 89.4 [85.5, 93.2]
+QuestEval 22.6 27.0 66.0* [58.7, 72.8]
+SummaC 16.4 24.5 61.7* [54.2, 68.7]
+AlignScore 36.7 44.6 75.9* [69.5, 81.8]
+DAE 20.6 20.6 56.8* [52.6, 60.9]
+PlainQAFact 56.9 68.2 90.2 [85.9, 93.9]
+Table 9: Evaluation results of automatic metrics on the CELLS dataset [28] with entity swap non-factual plain language summary perturbation.
+Results are evaluated using Kendall’s $\tau$, Pearson correlation, and AUC (AUC-ROC).
+* indicates a statistically significant difference compared to PlainQAFact ($p<0.01$).
+The standard deviations (std.) over five runs are: 0.1 (PlainQAFact), 0.2 (Llama 3.1), and 2.5 (GPT-4o).
+Dataset: CELLS AE Kendall’s $\tau$ Pearson AUC (99% CI) std.
+w/ Heuristic Entity Swap Llama 3.1 56.9 68.2 90.2 [85.9, 93.9] 0.1
+PromptRank 49.6 58.8 85.0* [79.7, 89.6] -
+w/ LLM Perturbation Llama 3.1 62.1 75.2 93.8 [90.3, 96.7] 0.1
+PromptRank 56.3 66.6 89.6* [85.3, 93.3] -
+Table 10: Ablation study of answer extraction approach in the PlainQAFact on the CELLS dataset.
+w/ Heuristic Entity Swap uses APPLS entity swap [12] approach to create non-factual plain language summaries, while w/ LLM Perturbation applies GPT-4o to perturb the plain language summaries.
+AE = answer extraction.
+AUC = AUC-ROC.
+* indicates a statistically significant difference compared to Llama 3.1 ($p<0.01$).
+We run each setting for five times and report the standard deviation (std.).
+Note that only the settings using LLMs may produce fluctuated scores.
+As we discussed in Section 5.1, in addition to evaluating the performance of automatic factual consistency evaluation metrics on LLM-perturbed datasets, we also use the non-LLM entity swap approach introduced in Guo et al.
+[12] to assess the robustness of our proposed metric.
+Specifically, we perturb plain language summaries using the CELLS dataset following the same entity swap protocol as APPLS [12].
+The entity swap system identifies specific “entity spans” in the text and substitutes them with existing terms from the UMLS knowledge base.
+As shown in Table 9, PlainQAFact achieves the best performance among all automatic evaluation metrics on the CELLS dataset.
+PlainQAFact achieves the highest correlation with human judgement (AUC=90.2, 99% CI [85.9, 93.9]).
+As comparative effect sizes, we report pairwise differences between PlainQAFact and the strongest baseline QAFactEval.
+The $\Delta\mathrm{AUC}=0.7$, where $\Delta$ is PlainQAFact minus the baseline.
+After correction, PlainQAFact shows significantly higher AUC than Llama 3.1, QuestEval, SummaC, AlignScore, and DAE ($p<0.01$, Holm-Bonferroni corrected for all), but is not significantly different from QAFactEval ($p=0.64$) or GPT-4o ($p=1.0$).
+We also report an ablation study using a non-LLM answer extractor to show the effectiveness of our LLM-powered answer extraction approach.
+In Table 10, our answer extraction method using Llama 3.1 consistently achieves higher correlation (AUC=90.2, 99% CI [85.9, 93.9] in w/ heuristic entity swap and AUC=93.8, 99% CI [90.3, 96.7] in w/ LLM perturbation) compared with the non-LLM answer extractor (i.e., PromptRank).
+Using Llama 3.1 shows significantly higher AUC than PromptRank in both settings ($p<0.01$, Holm-Bonferroni corrected for all), highlighting the robustness and effectiveness of LLM-powered answer extraction in factual consistency evaluation.
+Appendix I Summary-Level Explanation-Only Evaluation
+Figure 5: Overall performance on summaries containing added information (i.e., elaborative explanations) from FactPICO [9].
+The std.
+of PlainQAFact, Llama 3.1, and GPT-4o are 0.2, 0.2, and 3.3, respectively, based on five runs of each metric.
+Figure 6: Overall performance on summaries with elaborative explanations from FareBio [47].
+The std.
+for PlainQAFact, Llama 3.1, and GPT-4o are 0.03, 0.3, and 5.7, respectively, computed over five runs.
+Similar to our sentence-level explanation annotations in PlainFact, FactPICO [9] and FareBio [47] datasets also provide human-annotated explanation information in plain language summaries.
+According to the results shown in Section 5.2, we also assess the performance of five factual consistency evaluation metrics under FactPICO and FareBio datasets.
+In the FactPICO, the elaborative explanation is defined as “added information.” We first select summaries that include added information, and then we label a summary as “non-factual” if it contains any non-factual added information as determined by annotators.
+We generate a summary-level explanation-only FactPICO dataset consisting of 231 summary-abstract pairs.
+As shown in Figure 5, AlignScore achieves the best performance across all three evaluation criteria.
+However, since the summaries in the FactPICO dataset are generated using LLMs, the factual consistency between the generated summaries and the original source abstracts is not guaranteed.
+For instance, we treat a summary as “factual” indicates that all the added information is factual, but it does not necessarily reflect the factual consistency of other content.
+Therefore, the results presented in Figure 5 may become biased.
+For FareBio [47], we collect explanation sentences based on both “faithfulness” and “factual hallucination” labels in the original dataset.
+As shown in Figure 6, PlainQAFact achieves the best performance compared with all other metrics, showing a clear advantage over GPT-4o.
+Since the FareBio dataset provides sentence-level annotations of factual consistency for each sentence in the plain language summaries, the results are more reliable than those from FactPICO.
+On both datasets, we do not find statistically significant differences in AUC-ROC metric by $p<0.01$.
+While some existing metrics perform similarly in certain settings, our metric consistently shows better performance in correlation.
+Appendix J Quantitative Error Analysis of FareBio
+Given the Table 6, we further conduct quantitative analysis regarding failure cases in FareBio using our proposed metric.
+We randomly sample 30 plain language summaries (both factual and non-factual) from FareBio and evaluate all sentences within them, resulting in 324 split sentences.
+We categorize errors into three main types:
+1. QA Pipeline Errors (34.0% of errors): misclassification of sentence types, retrieval of irrelevant context, and generation of unanswerable questions.
+2. Answer Extraction and Evaluation Issues (18.4% of errors): failure to extract numerical values and answer mismatches when using BERTScore.
+3. Dataset Labeling Issues (47.7% of errors): inconsistencies in FaithBio’s annotation protocol.
+Table 11 presents the complete quantitative breakdown.
+We identify 109 errors across the 324 sentences (33.6% error rate).
+Notably, 47.7% of errors occurs in sentences mislabeled as factual and non-factual, with most being false negatives.
+FareBio uses a different annotation protocol than PlainFact.
+All source-grounded simplification sentences are implicitly treated as factual.
+Annotators do not evaluate factuality once a sentence is labeled as faithful.
+This creates label noise when evaluating simplification sentences, as the ground truth may not reflect actual factual consistency.
+Our analysis reveals that PlainQAFact performs better on PlainFact (where annotation protocols align with our metric’s design) than on FareBio.
+The performance gap reflects both limitations in our approach, particularly for numerical verification, and incompatibilities between FareBio’s annotation protocol and our QA components.
+Error Type Count % of Errors % of Total In F In NF
+QA Pipeline Errors
+Misclassification 13 11.9% 4.0% 10 3
+Retrieval quality 9 8.3% 2.8% 8 1
+Unanswerable generated question 15 13.8% 4.6% 11 4
+Answer Extraction and Evaluation Issues
+Numerical value mis-extraction 16 14.7% 4.9% 11 5
+Answer match misalignment 4 3.7% 1.2% 3 1
+Dataset Issues
+False positive (F labeled as NF) 10 9.2% 3.1% 0 10
+False negative (NF labeled as F) 42 38.5% 13.0% 42 0
+Total Errors 109 100.0% 33.6% 85 24
+Correct Classifications 215 66.4% 114 101
+Total Sentences Analyzed 324 100.0% 198 126
+Table 11: Quantitative error analysis of PlainQAFact on 60 randomly sampled summaries from FareBio dataset (30 factual, 30 non-factual).
+Summaries are split into sentences when evaluating the factual consistency.
+F = factual sentences; NF = non-factual sentences.

### v4

@@
 We further analyze PlainQAFact’s effectiveness across external knowledge sources, answer extraction strategies, answer overlap measures, and document granularity levels, refining its overall factual consistency assessment.
-Taken together, our work presents a sentence-aware, retrieval-augmented metric targeted at elaborative explanations in biomedical PLS tasks, providing the community with both a robust benchmark and a practical tool to advance reliable and safe plain language communication in the medical domain.
+Taken together, our work presents a sentence-aware, retrieval-augmented metric targeted at elaborative explanations in biomedical PLS tasks, providing the community with both a new benchmark and a practical evaluation tool to advance reliable and safe plain language communication in the medical domain.
 PlainQAFact and PlainFact are available at: [Uncaptioned image] https://github.com/zhiwenyou103/PlainQAFact.
@@
 While human evaluation is reliable [11], it is costly and difficult to scale, particularly in biomedical contexts where domain expertise is required.
-Commonly used factuality metrics can effectively verify content supported by the source but fail to assess factual consistency of added information [12].
+Commonly used factual consistency evaluation metrics can effectively verify content supported by the source but fail to assess factual consistency of added information [12].
 However, these metrics depend heavily on high-quality reference summaries, which are often unavailable in plain language summaries.
 Recent prompt-based evaluation techniques show promise [13, 14], but their sensitivity to factual perturbations in elaborative content remains limited [12].
-The lack of suitable benchmark datasets further hinders progress.
+The lack of suitable benchmark datasets further hinders the progress of factual consistency evaluation.
 Many existing datasets are constructed from LLM-generated summaries or apply rule-based perturbations to simulate non-factual content.
@@
 PlainQAFact assigns a high score, reflecting strong alignment between the extracted and gold answers.
-To address these challenges, we propose a targeted retrieval-based metric for factual consistency evaluation in PLS (Figure 1).
-We introduce a new expert-annotated dataset, PlainFact, consisting of human-written plain language summaries aligned with scientific abstracts.
+To address these challenges, we first introduce a new expert-annotated dataset, PlainFact, consisting of human-written plain language summaries aligned with scientific abstracts.
 Each sentence is labeled with its type (elaborative explanation vs.
 source simplification), functional role, and alignment to the source (§3.1).
-Building on this and prior studies for the QA-based evaluation metrics in the general domain [20, 21], we present PlainQAFact, a dual-stage QA-based evaluation metric that selectively applies retrieval only to elaborative explanations (§3.2).
-Experiment results on several PLS datasets demonstrate the effectiveness of our approach, particularly in assessing complex, elaborative content (§4, §5).
-Taken together, our work presents the first metric specifically designed to evaluate the factual consistency in elaborative explanations in PLS, providing the community with both a robust benchmark and a practical tool to advance reliable and safe plain language communication in the medical domain.
+Building on this and prior studies for the QA-based evaluation metrics in the general domain [20, 21], we propose a targeted retrieval-based metric, PlainQAFact, for factual consistency evaluation in PLS (Figure 1).
+It is a dual-stage QA-based evaluation metric that selectively applies retrieval only to elaborative explanations (§3.2).
+Experiment results on several PLS datasets demonstrate the effectiveness of our metric, particularly in assessing complex, elaborative content (§4, §5).
+Taken together, our work introduces an automatic evaluation metric specifically designed to assess factual consistency for PLS.
+In addition, we provide the community with a new benchmark that supports nuanced evaluation of simplified and elaborative content.
+These contributions aim to promote more reliable and safe plain language communication in the medical domain.
 2 Related Work
@@
 Evaluating factual consistency in biomedical PLS is particularly challenging, as it requires domain expertise.
-Entailment- [15, 16], similarity- [17, 18], model- [19] and QA-based [20, 21] metrics are commonly used for factual consistency assessment but rely heavily on high-quality reference summaries, which are often unavailable or difficult to obtain for PLS.
+Entailment- [15, 16], similarity- [17, 18], model- [19], and QA-based [20, 21] metrics are commonly used for factual consistency assessment but rely heavily on high-quality reference summaries, which are often unavailable or difficult to obtain for PLS.
 Recent advancements in prompt-based evaluation show promise [13]; however, their sensitivity to factuality perturbations in PLS remains limited [12].
-To address these limitations, we propose a reference-free solution for factual consistency evaluation of PLS that effectively assesses factual consistency with external information retrieval to augment the reference summary.
+To address these limitations, we propose a reference-free solution for evaluating factual consistency in PLS, which effectively assesses factual consistency by leveraging external information retrieval to augment the reference summary.
 Retrieval-Augmented Generation
@@
 However, retrieval-augmented approaches have not been extensively explored for factual consistency evaluation in PLS, despite their potential for addressing elaborative explanations.
-In this work, we investigate retrieval-augmented QA to enhance PLS factual consistency assessment while also examining its limitations.
+In this work, we investigate retrieval-augmented QA scoring to enhance factual consistency assessment in PLS while also examining its limitations.
 3 Methods
 We first introduce the protocol of PlainFact dataset curation (§3.1).
-Based on this dataset, we propose PlainQAFact, a two-stage retrieval-augmented QA framework for factual consistency evaluation in PLS tasks (§3.2).
+Based on this dataset, we propose PlainQAFact, a two-stage retrieval-augmented QA scoring framework for factual consistency evaluation in PLS tasks (§3.2).
 3.1 PlainFact Benchmark
@@
 3.2 PlainQAFact Framework
-PlainQAFact conducts fine-grained factual consistency evaluation for plain language summaries by first segmenting each summary into sentences, classifying as either simplification or explanation, and retrieving tailored external knowledge within a retrieval-augmented QA framework.
+PlainQAFact conducts fine-grained factual consistency evaluation for plain language summaries by first segmenting each summary into sentences, classifying as either simplification or explanation, and retrieving tailored external knowledge within a retrieval-augmented QA scoring framework.
 Figure 1 provides an overview of its three key components: sentence-level classification (§3.2.1), domain knowledge retrieval (§3.2.2), and dual-stage QA-based factual consistency evaluation (§3.2.3).
@@
 3.2.3 QA Evaluation Components
-QA-based metrics have proved to be effective than other factual consistency evaluation metrics [12] in general summarization tasks and align more closely with human annotations [20, 39].
-In this study, we adopt a QA-based approach as a backbone of our evaluation metric to verify factual consistency, incorporating answer extraction, question generation, question filtering, question answering, and answer overlap calculation modules.
+QA-based scoring metrics have proved to be effective than other factual consistency evaluation metrics [12] in general summarization tasks and align more closely with human annotations [20, 39].
+In this study, we adopt a QA-based scoring approach as a backbone of our evaluation metric to verify factual consistency, incorporating answer extraction, question generation, question filtering, question answering, and answer overlap calculation modules.
 Gold Answer Extraction (AE)
@@
 To compare answer extraction strategies, we use PromptRank as the baseline, and evaluate LLM-based extractors, including an open-source Llama 3.1 8B Instruct model (Llama 3.1) [41] and close-source GPT-4o (GPT-4o) [34].
-Prompts for Llama 3.1 and GPT-4o are in E.4.
+Prompts for Llama 3.1 and GPT-4o are presented in E.4.
 Question Generation (QG)
@@
 Questions generated by the QG model (§3.2.3) may not always be answerable.
-For example, the QG model generates “<What can occur over several days?>,” which is vague and hard to predict the correct answer.
+For example, the QG model generates “What can occur over several days?,” which is vague and hard to predict the correct answer.
 To prevent these unanswerable questions from impacting the final evaluation performance, we follow QAFactEval [39] and remove the unanswerable questions using a pre-trained Electra-large model [44].
@@
 (4) QAFactEval [39] is a QA-based factuality evaluation metric that assesses summary consistency by generating questions from the source document and comparing the model-generated answers with expected answers.
-(5) QuestEval [50] is a reference‐less evaluation metric employs a T5 model to generate questions from the source document and verifies whether the summary can correctly answer them.
+(5) QuestEval [50] is a reference‐less evaluation metric employs a T5 [51] model to generate questions from the source document and verifies whether the summary can correctly answer them.
 5 Results and Analysis
@@
 Detailed experiment results are reported in  H.
-In summary, PlainQAFact is a more consistent and effective metric for factual consistency compared to existing automatic approaches.
+In summary, PlainQAFact is a more consistent and effective metric for factual consistency evaluation compared to existing automatic metrics.
 It clearly improves over prior QA-based metrics on CELLS and matches or exceeds them on PlainFact.
@@
 PlainQAFact detects added elaborative explanations and checks them with domain knowledge retrieval, which helps when factual consistency depends on facts beyond simple restatement from the source.
-Since 44% of sentences in our curated benchmark are labeled as elaborative by human annotators, we believe that PlainQAFact is the more suitable and robust factual consistency evaluation metric in this scenario.
+Since 44% of sentences in our curated benchmark are labeled as elaborative explanations by human annotators, we believe that PlainQAFact is the more suitable and robust factual consistency evaluation metric in this scenario.
 Component Method Choice $\tau$ Pearson AUC (99% CI) std.
@@
 Table 5: Error analysis of PlainQAFact with intermediate metric outputs.
-“Retrieved Knowledge” refers to the source texts retrieved for each plain language sentence.
+Retrieved Knowledge refers to the source texts retrieved for each plain language sentence.
 We present correct and failure cases sampled only from elaborative explanation examples (i.e., sentences classified as explanations during evaluation).
@@
 6 Discussion and Conclusion
-Our study advances the assessment of hallucinations in the field of PLS by supplying both the first domain expert-annotated biomedical PLS benchmark PlainFact and a novel retrieval‐augmented QA‐based factual consistency evaluation metric PlainQAFact.
+Our study advances the assessment of hallucinations in the field of PLS by supplying both a domain expert-annotated biomedical PLS benchmark PlainFact and a novel retrieval‐augmented QA‐based factual consistency evaluation metric PlainQAFact.
 Unlike existing biomedical domain corpora that either lack fine‐grained labels for elaborative content [58] or focus solely on text simplification (e.g., FactPICO[9]), PlainFact captures sentence‐level distinctions, including simplification versus explanation, functional roles, and explicit alignment to source sentences.
@@
 Then, retrieving targeted domain knowledge for each elaborative explanation before posing and answering fact‐checking questions.
-The ablations demonstrate that selective retrieval yields substantial gains in Kendall’s $\tau$, Pearson correlation, and AUC‐ROC over both end‑to‑end LLM judges[41] and alignment‐based scorers[19], highlighting that judicious incorporation of external evidence is essential for robust factual consistency assessment.
+The ablations demonstrate that selective retrieval yields substantial gains in Kendall’s $\tau$, Pearson correlation, and AUC‐ROC over both end‑to‑end LLM judges[41] and alignment‐based scorers[19], highlighting that the incorporation of external evidence is essential for robust factual consistency assessment.
 Beyond its empirical strengths, PlainQAFact serves as a plug‑and‑play evaluation metric for future summarization systems.
@@
 More importantly, PlainQAFact is developed with the open-source backbone model (i.e., Llama 3.1), ensuring broad use with transparency, reproducibility, and lower computational costs.
-Looking forward, extending our proposed evaluation method to interactive human–AI workflows, where clinicians or subject‐matter experts verify QA outputs on‐the‑fly, could further enhance trust and adoption in high‐stakes domains, including medicine, healthcare, or law.
-Despite these promising results, we consider the following limitations of this study.
-(1) Our error analysis reveals that misclassifications, particularly between source simplifications and elaborative explanations, can lead to retrieval mismatches and unanswerable QA prompts.
+Looking forward, our proposed evaluation metric could serve as a diagnostic tool for researchers and developers to identify and analyze specific factual inconsistencies in PLS systems, guiding future improvements in both model development and evaluation methodology.
+Despite these promising results, we consider the following limitations of this study: (1) Our error analysis reveals that misclassifications, particularly between source simplifications and elaborative explanations, can lead to retrieval mismatches and unanswerable QA prompts.
 These issues underscore the need for more nuanced classification and question generation modules.
-Although we provide substitute solutions for classifying input summaries or sentences, our fine-tuned classifier is limited to the biomedical domain and specifically designed for PLS tasks, fine-tuned through PlainFact.
-We acknowledge that the dataset used for model fine-tuning is insufficient, resulting in only moderate classification accuracy.
-Given the scarcity of high-quality human-annotated data for classifying source simplification and elaborative explanation information, future efforts are needed on developing domain-specific (e.g., law, healthcare, social science, etc.) classifiers [56] for factual consistency evaluation.
+Given the fair IRA scores, especially for the factuality type field, we acknowledge that our classification method is not fully reliable.
+The annotated dataset reflects subjective judgments, particularly when distinguishing source simplification from elaborative explanation, and the training set is limited in size.
+We advise users to apply our pretrained classifier with caution when using it for later factual consistency evaluation steps.
+Future efforts are needed on developing domain-specific (e.g., law, healthcare, social science, etc.) classifiers [57] for factual consistency evaluation.
 (2) While our experiments indicate that our retrieval-augmented evaluation metric can improve factual consistency assessment in most of the settings, especially in elaborative explanation evaluation, the computational time increases compared to NLI-based evaluation metrics and LLM-based evaluators.
@@
 PlainQAFact is a research-oriented metric that prioritizes interpretability and transparency over speed.
-(3) Given the fair IRA scores, especially for the factuality type field, we acknowledge that our classification method is not fully reliable.
-The annotated dataset reflects subjective judgments, particularly when distinguishing source simplification from elaborative explanation, and the training set is limited in size.
-We advise users to apply our pretrained classifier with caution when using it for later factual consistency evaluation steps.
+(3) The quantitative error analysis on FareBio (Table 11) shows a notable limitation of PlainQAFact in handling numerical content.
+The answer extraction step may overlook numerical values, preventing them from being verified in subsequent question generation and answering stages.
+Consequently, PlainQAFact may underestimate factual inconsistency in summaries with incorrect numerical information.
+(4) PlainFact and PlainQAFact are built upon the CELLS corpus, whose plain language summaries are written by the original article authors.
+While this ensures domain expertise, it does not guarantee complete factual consistency, as expert-authored summaries may contain unintentional inaccuracies or oversimplifications.
 Acknowledgment
@@
 Appendix A Dataset Annotation Protocol
-We developed a comprehensive annotation procedure for freelancers on Upwork to conduct fact-checking annotations.
+We develop a comprehensive annotation procedure for freelancers on Upwork to conduct fact-checking annotations.
 Our annotation procedure involves two stages, starting with thorough training using detailed examples to ensure consistent understanding of the task between annotators.
 Annotators receive a spreadsheet where each row contains a pair of data: a sentence extracted from the plain language summary and its corresponding scientific abstract.
-For every sentence-abstract pair, the annotators are required to label three features: External, Category, and Relation, with the appropriate labels.
+For every sentence-abstract pair, the annotators are required to label three features: Factuality type, Functional role, and Sentence alignment, with the appropriate labels.
 Step 1: Sentence Annotation Instructions Compared to the scientific abstract, analyze each sentence of the plain language summary across three dimensions: external information, category, and relation.
 1.
-External: Determine whether the sentence includes information does not present in the scientific abstract.
+Factuality type: Determine whether the sentence includes information does not present in the scientific abstract.
 1) Yes: The sentence contains external information that is not explicitly mentioned, paraphrased, or implied in the scientific abstract.
@@
 2.
-Category: Classify the sentence of the plain language summary into one of the following categories (you can only choose one category per annotation) 1) Definition: Provides a fundamental explanation of a term.
+Functional role: Classify the sentence of the plain language summary into one of the following categories (you can only choose one category per annotation) 1) Definition: Provides a fundamental explanation of a term.
 2) Background: Information that helps understand the term within the context of the abstract, such as relevance, significance, or motivation.
@@
 5) Other: For sentences that do not fit into the categories above, please indicate the category 3.
-Relation: Identify the sentence(s) in the scientific abstract that the sentence of the plain language summary is related to.
+Sentence alignment: Identify the sentence(s) in the scientific abstract that the sentence of the plain language summary is related to.
 Use indices of sentences from the scientific abstract to link the sentence of the plain language summary.
@@
 Each annotator will annotate 40 summary-abstract pairs to ensure each sentence of the plain language summary has two sets of annotations from different people.
-For each row of the spreadsheet, they need to annotate three columns: “External,” “Category,” and “Relation.”
-Annotation Fields to Be Completed <<TO BE ANNOTATED>> External: Mark “Yes” or “No” to indicate if the sentence in the ‘Target_Sentence’ column contains external information.
-<<TO BE ANNOTATED>> Category: Choose the most fitting category of the sentence from the list (Definition, Background, Example, Method/Result, Other).
-<<TO BE ANNOTATED>> Relation: List the relevant sentence indices from the abstract in the ‘Original_Abstract’ column that relate to the plain language sentence (e.g., s10_1,s10_5).
+For each row of the spreadsheet, they need to annotate three columns: “Factuality type,” “Functional role,” and “Sentence alignment.”
+Annotation Fields to Be Completed <<TO BE ANNOTATED>> Factuality type: Mark “Yes” or “No” to indicate if the sentence in the ‘Target_Sentence’ column contains external information.
+<<TO BE ANNOTATED>> Functional role: Choose the most fitting category of the sentence from the list (Definition, Background, Example, Method/Result, Other).
+<<TO BE ANNOTATED>> Sentence alignment: List the relevant sentence indices from the abstract in the ‘Original_Abstract’ column that relate to the plain language sentence (e.g., s10_1,s10_5).
 For example, filling in ‘s10_1,s10_5,s10_9’ if you think these three sentences from the abstract are relevant to sentence t1_3.
@@
 Need External Information? yes
-Category Background
-Relation external
+Functional role Background
+Sentence alignment external
 Corresponding Abstract None
@@
 Need External Information? yes
-Category Background
-Relation s10_1,s10_2
+Functional role Background
+Sentence alignment s10_1,s10_2
 Corresponding Abstract <<s10_1>>Acute cardiovascular events represent a therapeutic challenge.
@@
 Need External Information? no
-Category Method/Result
-Relation s15_16,s15_17,s15_20
+Functional role Method/Result
+Sentence alignment s15_16,s15_17,s15_20
 Corresponding Abstract <<s15_16>>For treatment failure, the difference between doxycycline and tetracycline is uncertain (very low‐certainty evidence).
@@
 A “Yes” label indicates that the sentence includes information not explicitly mentioned in the abstract, and vice versa.
-The “Relation” feature identifies the sentence(s) in the scientific abstract most relevant to the plain language summaries; if no corresponding content exists in the abstract, the relation is marked as “external.” We randomly sample three examples from the dataset to to illustrate the dataset’s structure.
+The “Sentence alignment” feature identifies the sentence(s) in the scientific abstract most relevant to the plain language summaries; if no corresponding content exists in the abstract, the relation is marked as “external.” We randomly sample three examples from the dataset to to illustrate the dataset’s structure.
 Additionally, indexes have been created for both the plain language sentences and the abstract sentences to facilitate annotation.
@@
 In accordance with the benchmark annotation protocol (A), we employ GPT-4o as a classifier to determine whether a given sentence or summary requires elaborative explanations for factual verification.
-For both stages, we set the max_tokens parameter to 512 and configure the temperature to 0 for GPT-4o and 0.01 for the Llama 3.1 8B Instruct model.
+For both stages, we set the max_tokens parameter to 512 and configure the temperature to 0 for GPT-4o and 0.01 for the Llama 3.1 model.
 Developer
@@
 Example 1: Prompt of GPT-4o as the Classifier.
-For the AE stage, we explore both GPT-4o and Llama 3.1 8B Instruct as backbone models to assess the factual consistency evaluation performance of PlainQAFact.
+For the AE stage, we explore both GPT-4o and Llama 3.1 as backbone models to assess the factual consistency evaluation performance of PlainQAFact.
 Following the task description outlined in QAFactEval [39], we instruct both LLMs to extract potential answer entities from the input PLS.
@@
 Extract a comma-separated list of the most important keywords from the following text: <input>
-Example 3: Prompt of Llama 3.1 8B Instruct as the Answer Extractor.
-Additionally, we report the performance of using Llama 3.1 8B Instruct as a judge to evaluate the factual consistency of a given PLS based on its source scientific abstract.
+Example 3: Prompt of Llama 3.1 as the Answer Extractor.
+Additionally, we report the performance of using Llama 3.1 as a judge to evaluate the factual consistency of a given PLS based on its source scientific abstract.
 System
@@
 Factuality score (only output a numeric score): <score>
-Example 4: Prompt of Llama 3.1 8B Instruct as a factual consistency judge.
+Example 4: Prompt of Llama 3.1 as a factual consistency judge.
 For the LLM-based perturbation of PlainFact and CELLS datasets, we follow the protocol of APPLS [12] on faithfulness criteria.